### v1

# A direct proof of the existence of MME for finite horizon Sinai billiards

Abstract.
The Sinai billiard map $T$ on the two-torus, i.e., the periodic Lorentz gaz, is a discontinuous map. Assuming finite horizon and bounded complexity, we prove that the Kolmogorov–Sinai entropy map associated with the billiard map $T$ is upper semi-continuous, as well as the compactness of the set of $T$-invariant measures. In particular, for the potentials $g\equiv 0$ and $g=-h_{\rm top}(\Phi_{1})\tau$, we recover the recent results of the existence of measures of maximal entropy (MME) for both the billiard map and flow $\Phi_{t}$, due to Baladi and Demers for $T$, jointly with Carrand for $\Phi_{t}$. For general finite horizon Sinai billiards, we provide bounds on the defect of upper semi-continuity of the entropy map and on the topological tail entropy. The bounded complexity condition is expected to hold generically.
The author is grateful to Viviane Baladi, Mark Demers and David Burguet for useful discussions and comments. Research supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 787304).
1. Introduction
1.1. Equilibrium States in Dynamical Systems
Taking inspiration from physics, more specifically thermodynamic, where “nature tends to maximize free energy”, it is tempting to investigate the existence of maximizing measures (or states in physics terminology) for
(1.1) $\mu\mapsto h_{\mu}(T)+\int_{M}g\,\mathrm{d}\mu\,,$
where $\mu$ ranges over the set $\mathcal{M}(M,T)$ of Borel probability measures on $M$ invariant under some transformation $T:M\to M$, $h_{\mu}(T)$ is the Kolmogorov-Sinai entropy of $(T,\mu)$, and $g:M\to\mathbbm{R}$ is some fixed function called potential – in the physical analogy, $\int_{M}-g\,\mathrm{d}\mu$ corresponds to the internal energy of the state $\mu$. Maximal measures are called equilibrium states.
This problem has been introduced by Ruelle in the case of finite rank lattices [16, 17], where a particle lies at each vertex. The rank one case, assuming translation symmetry of states, corresponds to a subshift of finite type. There, for any smooth potential $g$ (e.g. Hölder), there exists a unique equilibrium state $\mu_{g}$, and it enjoys strong statistical properties (exponential mixing, CLT) [7]. Furthermore, $\mu_{g}$ can be constructed by joining the left and right eigenvectors with maximal eigenvalue of a twisted Ruelle transfer operator.
More general transformations can be studied: through the use of coding with Markov partitions, the case of Anosov or Axiom A maps $T$ can be carried out [7]. In this case, for any Hölder potential, the uniquely associated equilibrium state is Bernoulli. Note that the process of coding can be bypassed by studying a twisted Ruelle transfer operator acting this time on anisotropic Banach spaces [1]. Although this strategy is usually very technical, it can lead to strong statistical properties (exponential mixing, CLT, ASIP) – provided that the operator has, for example, good spectral properties – and has the advantage to be applicable to more general situations.
Finally, note that if $T$ is uniformly hyperbolic on a compact manifold $M$, then it is expansive [14, Corollary 6.4.10], which in turn implies that $\mu\mapsto h_{\mu}(T)$ is upper semi-continuous [19, Theorem 8.2], giving existence for $\mu_{g}$ for any continuous potential $g$ since $\mathcal{M}(M,T)$ is compact. For continuous maps $T$, the defect of upper semicontinuity can be bounded by the topological tail entropy $h^{*}(T)$ [15]. Uniqueness can be deduced from expansiveness and the so called specification property [6].
1.2. The case of Sinai Billiards
In this paper we are interested in the case where $T:M\to M$ is the collision map associated to a finite horizon Sinai billiard (see definition below).
A Sinai billiard on the torus is (the quotient modulo $\mathbbm{Z}^{2}$, for position, of) the periodic planar Lorentz gas (1905) model for the motion of a single dilute electron in a metal. The scatterers (corresponding to atoms of the metal) are assumed to be strictly convex (but not necessarily discs). Such billiards have become fundamental models in mathematical physics.
Many equilibrium states for the billiard map (resp. for the billiard flow) have already been studied, such as the SRB measure [10] which enjoys exponential mixing [20, 13] (resp. [5]), the measure of maximal entropy, constructed in [3] (resp. [9, 2]) enjoys polynomial mixing [12], equilibrium states associated to piecewise-Hölder potentials [9, 2] which are Bernoulli, and finally equilibrium states with geometric potentials (which are not Hölder) [4] which are exponentially mixing. Most of the constructions – those of [3, 9, 2, 4] – or the study of the equilibrium states – as for [13, 5] – relie on the use of a transfer operator acting on specific anisotropic Banach spaces. In [2, 3, 4, 9], the studied equilibrium states are proved to be unique. In the construction performed in [3, 9, 2], an additional assumption on the billiard table is made, namely sparse recurrence to singularities, that is
(1.2) $P_{\rm top}(T,g)-\sup g>s_{0}\log 2,$
where $P_{\rm top}(T,g)$ is the topological pressure of $g$ under the transformation $T$, and $s_{0}\in(0,1)$ is defined in the next subsection.
Bounded complexity is a more restrictive assumption than sparse recurrence (used in [3, 2]). Indeed, for $g\equiv 0$ (respectively, for $g=-h_{\rm top}(\Phi_{1})\tau$), the left hand side of (1.2) reduces to $P_{\rm top}(T,g)-\sup g=h_{\rm top}(T)>0$ (resp. $P_{\rm top}(T,g)-\sup g\geqslant h_{\rm top}(\Phi_{1})\min\tau>0$), but Demers and Korepanov proved in [12] that if the complexity is bounded, then $\inf_{n_{0},\varphi_{0}}s_{0}(n_{0},\varphi_{0})=0$. According to [18], bounded complexity is a generic property among finite horizon Sinai billiards.
In this paper, we prove in Section 2 that for any finite horizon Sinai billiard with bounded complexity and any piecewise continuous potential, the map (1.1) is upper semi-continuous. In Section 3, we prove that for any finite horizon Sinai billiard, $\mathcal{M}(M,T)$ is compact. In particular for $g\equiv 0$ and $g=-h_{\rm top}(\Phi_{1})\tau$ (see [9]), we recover the existence of the MME for the collision map and the billiard flow respectively. Additionally, in Section 4, we give bounds on the defect of upper semi-continuity and on the topological tail entropy for general finite horizon Sinai billiards.
The problem of uniqueness is not covered by this paper.
1.3. Precise statement of results
Formally, a Sinai billiard table $Q$ on the two-torus $\mathbbm{T}^{2}$ is a set $Q=\mathbbm{T}^{2}\smallsetminus B$, with $B=\bigcup_{i=1}^{D}B_{i}$ for some finite number $D\geqslant 1$ of pairwise disjoint closed domains $B_{i}$, called scatterers, with $\mathcal{C}^{3}$ boundaries having strictly positive curvature – in particular, the scatterers are strictly convex. The billiard flow $\Phi_{t}:\Omega\to\Omega$ is the motion of a point particle travelling at unit speed in $Q$ with specular reflections off the boundary of the scatterers. Identifying outgoing collisions with incoming ones in the phase space, that is $\Omega=Q\times\mathbbm{S}^{1}/\sim$, the billiard flow is continuous. However, the grazing collisions – those tangential to scatterers – give rise to singularities in the derivative [10]. The Sinai billiard map $T:M\to M$ – also called collision map – is the first return map of the single point particle to the scatterers, that is $M=\partial Q\times[-\pi/2,\pi/2]$. Because of the grazing collisions, the Sinai billiard map is a discontinuous map.
As in [3, 20], we say that the billiard table $Q$ has finite horizon if the billiard flow does not have any trajectories making only tangential collisions – in particular, this implies that the return time function $\tau$ to the scatterers is bounded. All Sinai billiards in this paper are assumed to have finite horizon. In this case, $\Phi_{t}$ is a suspension flow over $T$, under $\tau$.
Define $\mathcal{S}_{0}=\{(r,\varphi)\in M\mid|\varphi|=\pi/2\}$ the set of grazing collisions, and $\mathcal{S}_{\pm n}=\bigcup_{i=0}^{n}T^{\mp i}\mathcal{S}_{0}$ the singular set of $T^{\pm n}$. Notice that $\mathcal{S}_{n}$ (resp. $\mathcal{S}_{-n}$) is made of a finite number of decreasing (resp. increasing) curves in the $(r,\varphi)$ coordinates, whereas $M$ is a two dimensional manifold with boundaries. Let $\mathcal{M}_{-k}^{n}$ be the partition of $M\smallsetminus(\mathcal{S}_{-k}\cup\mathcal{S}_{n})$ into maximal connected components (which are all open sets of $M$). Furthermore, define $\mathcal{P}$ the partition of $M$ into maximal connected components on which both $T$ and $T^{-1}$ are continuous, and $\mathcal{P}_{-k}^{n}=\bigvee_{i=-k}^{n}T^{-i}\mathcal{P}$. As in [3], the topological entropy of $T$ is defined as $h_{*}\coloneqq\lim_{n}\tfrac{1}{n}\log\#\mathcal{P}_{0}^{n}$. We say that a potential $g$ is $\mathcal{M}_{-k}^{n}$-continuous if $g$ is bounded and $g|_{A}$ is continuous for each $A\in\mathcal{M}_{-k}^{n}$. We simply say that a potential is piecewise continuous if it is $\mathcal{M}_{-k}^{n}$-continuous for some $k,n\geqslant 0$.
One crucial object in the rest of this paper is sequence $K_{n}$, $n\geqslant 1$, called the complexity: each $K_{n}$ is defined as the maximal number of curves of $S_{n}$ intersecting at some common point. Chernov [11] proved that for any finite horizon Sinai billiard, there exists a constant $K$ such that $K_{n}\leqslant K\,n$. A linear growth of the complexity can be achieved if there exists a periodic orbit with a grazing collision. In a work in preparation [18], Tòth proved that for generic billiard table, the complexity is bounded.
Sections 2 and 3 are devoted to the proof of the following result.
Theorem 1.1.
For any finite horizon Sinai billiard with bounded complexity, and any piecewise-continuous potential $g:M\to\mathbbm{R}$, the map
$\mu\in\mathcal{M}(M,T)\mapsto h_{\mu}(T)+\int g\,\mathrm{d}\mu$
is upper semi-continuous.
In particular, $g\equiv 0$ and $g=-h_{\rm top}(\phi_{1})\tau$ admit equilibrium measures. In other words, $T$ and $\Phi_{t}$ have MME.
The upper semi-continuity statement in the above theorem follows from Proposition 2.2. The particular cases follow from the compactness of $\mathcal{M}(M,T)$ proved in Proposition 3.1. The relation between $g=-h_{\rm top}(\phi_{1})\tau$ and the MME of the Sinai billiard flow is proved in [9, Corollary 2.6].
In order to state our second result, we need to quantify the frequency of return of an orbit in a neighbourhood of $\mathcal{S}_{0}$. More precisely, for any $n_{0}\geqslant 1$ and $20<\varphi_{0}<\pi/$ 2, define
(1.3) $s_{0}(n_{0},\varphi_{0})\coloneqq\frac{1}{n_{0}}\sup_{M}\sum_{i=0}^{n_{0}-1}\mathbbm{1}_{(|\varphi|>\varphi_{0})}\circ T^{i},$
that is, the maximal frequency of $\varphi_{0}$-grazing collisions among $n_{0}$ consecutive collisions.
Finally, we recall the formal definition of the topological tail entropy (as in [8]). Let $d(x,y)$ be the Euclidean metric between $x$ and $y$ if they belong to the same connected component $M_{i}$ of $M=\bigsqcup_{i=1}^{D}M_{i}$, and otherwise fix $d(x,y)=10\max_{1\leqslant i\leqslant D}\mbox{diam}(M_{i})$. Denote $d_{n}$ the $n$-th Bowen distance, that is
$d_{n}(x,y)\coloneqq\max_{0\leqslant i<n}d(T^{i}x,T^{i}y),$
and $B(x,\varepsilon,n)$ the open ball centred at $x$ of radius $\varepsilon$ for the $d_{n}$ metric. A set $X\subset Y$ ($Y\subset M$) is said to be a $(n,\delta)$-spanning subset of $Y$ if $Y\subset\bigcup_{x\in X}B(x,\delta,n)$ and a $(n,\delta)$-separated subset of $Y$ if for all distinct $x,y\in X$, $d_{n}(x,y)\geqslant\delta$. We use the notation $r(n,\delta,Y)$ (respectively, $s(n,\delta,Y)$) for the minimal cardinality of a $(n,\delta)$-spanning subset of a set $Y\subset M$ (respectively, the maximal cardinality of a $(n,\delta)$-separated subset of $Y$). Now, define the topological tail entropy to be
(1.4) $h^{*}(T)\coloneqq\lim_{\varepsilon\to 0}\lim_{\delta\to 0}\limsup_{n\to+\infty}\frac{1}{n}\sup_{x\in M}r(n,\delta,B(x,\varepsilon,n)).$
The Section 4 is devoted to the proof of the following result.
Theorem 1.2.
For any finite horizon Sinai billiard table:
i) $\limsup_{\nu\rightharpoonup\mu}h_{\nu}(T)-h_{\mu}(T)\leqslant\left(2\mu\left(\bigcup_{n\geqslant 0}\mathcal{S}_{n}\right)+\mu(\mathcal{S}_{0})\right)h_{*}$,
ii) $h^{*}(T)\leqslant\left(3+2\left\lfloor\frac{\max\tau}{\min\tau}\right\rfloor\right)s_{0}\log 2K$, where $K$ is such that $K_{n}\leqslant Kn$ for all $n\geqslant 1$.
In particular, if the complexity is bounded, i) implies the upper semi-continuity of the entropy map, and ii) implies that $h^{*}(T)=0$.
The above theorem will follows from Corollary 4.3 for i) and from Corollary 4.7 for ii). Concerning the bounded complexity case, i) can be deduced from the proof of Proposition 2.1, [12, Proposition 5.6] and the invariance of the measure, and ii) follows from [12, Proposition 5.6].
2. Bounded complexity case
In this section we prove the upper semi-continuity of (1.1) under the bounded complexity assumption. We begin with an intermediate result. Since $\mathcal{P}$ is a generating partition, we begin by proving the continuity of each $\mu\mapsto\frac{1}{n}H_{\mu}(\mathcal{P}_{l}^{n})$ in Proposition 2.1, where
$H_{\mu}(\mathcal{A})\coloneqq-\sum_{A\in\mathcal{A}}\mu(A)\log\mu(A)$
denotes the static entropy of a partition $\mathcal{A}$. For this, we will use that when the complexity is bounded, $\mu(\partial A)=0$ for all $A\in\mathcal{P}_{l}^{n}$, all $n,l\geqslant 0$ and all $\mu\in\mathcal{M}(M,T)$. We deduce the upper semi-continuity of (1.1) in Proposition 2.2.
Proposition 2.1.
For any finite horizon Sinai billiard with bounded complexity and for any $n,\ell\geqslant 0$, the map $\mu\mapsto\frac{1}{n}H_{\mu}(\mathcal{P}_{-\ell}^{n})$ is continuous on $\mathcal{M}(M,T)$.
Proof.
Fix some $n,\ell\geqslant 0$. We start by proving that for all $A\in\mathcal{P}_{-\ell}^{n}$, $\partial A$ has measure zero. Notice that for all $A\in\mathcal{P}_{-\ell}^{n}$, $\partial A\subset\mathcal{S}_{-\ell-1}\cup\mathcal{S}_{n+1}$ (see [3, Lemma 3.3]). Now, for all $T$-invariant measure $\mu$, we have
$\mu(\mathcal{S}_{0})\leqslant\mu(\{|\varphi|>\varphi_{0}\})=\frac{1}{n_{0}}\int\sum_{i=0}^{n_{0}-1}\mathbbm{1}_{\{|\varphi|>\varphi_{0}\}}\circ T^{i}\,\mathrm{d}\mu\leqslant\frac{1}{n_{0}}\sup_{M}\sum_{i=0}^{n_{0}-1}\mathbbm{1}_{\{|\varphi|>\varphi_{0}\}}\circ T^{i}=s_{0}(n_{0},\varphi_{0}).$
Assuming bounded complexity, we can use [12, Proposition 5.6]: for any $\varepsilon_{0}>0$, there exists $n_{0},\varphi_{0}$ such that $s_{0}(n_{0},\varphi_{0})<\varepsilon_{0}$. Thus $\mu(\mathcal{S}_{0})=0$. Since $\mathcal{S}_{\pm k}\coloneqq\bigcup_{i=0}^{n}T^{\mp i}\mathcal{S}_{0}$, we get that $\mu(\mathcal{S}_{\pm k})=0$, for all $k\geqslant 0$ and all $\mu\in\mathcal{M}(M,T)$.
Let $\mu_{k}\rightharpoonup\mu$ in $\mathcal{M}(M,T)$ as $k\to\infty$. Let $A\in\mathcal{P}_{-\ell}^{n}$. By the above, $\mu(\partial A)=0$. Applying Portmanteau theorem, we thus get $\lim_{k\to\infty}\mu_{k}(A)=\mu(A)$. Since $\mathcal{P}_{-\ell}^{n}$ is a finite partition, and by continuity of $x\mapsto-x\log x$, we get the desired result. ∎
We can now prove the upper semi-continuity of (1.1).
Proposition 2.2.
Assuming bounded complexity, then for any piecewise-continuous potential $g:M\to\mathbbm{R}$ the map $\mu\mapsto h_{\mu}(T)+\int g\,\mathrm{d}\mu$ is upper semi-continuous on $\mathcal{M}(M,T)$.
Proof.
First, we prove that $\mu\mapsto h_{\mu}(T)$ is upper semi-continuous. Let $\mu\in\mathcal{M}(M,T)$, $\varepsilon>0$. By the continuity from Proposition 2.1, we get that for all $n\geqslant 1$, there is a neighbourhood $U_{n,\varepsilon}$ of $\mu$ such that for all $\nu\in U_{n,\varepsilon}$ $\frac{1}{n}H_{\mu}(\mathcal{P}_{0}^{n})\geqslant\frac{1}{n}H_{\nu}(\mathcal{P}_{0}^{n})-\varepsilon$. Therefore, since $\mathcal{P}$ is a generating partition,
$h_{\mu}(T)$ $=h_{\mu}(T,\mathcal{P})=\lim_{n\to\infty}\frac{1}{n}H_{\mu}(\mathcal{P}_{0}^{n})\geqslant\lim_{n\to\infty}\sup_{\nu\in U_{n,\varepsilon}}\frac{1}{n}H_{\nu}(\mathcal{P}_{0}^{n})-\varepsilon\geqslant\lim_{n\to\infty}\sup_{\nu\in U_{n,\varepsilon}}h_{\nu}(T,\mathcal{P})-\varepsilon$
$\geqslant\lim_{n\to\infty}\sup_{\nu\in U_{n,\varepsilon}}h_{\nu}(T)-\varepsilon=\limsup_{\nu\rightharpoonup\mu}h_{\nu}(T)-\varepsilon.$
Since this holds for any $\varepsilon>0$, we get the upper semi-continuity of the entropy.
We now prove that $\mu\mapsto\int g\,\mathrm{d}\mu$ is continuous. Let $g$ be a $\mathcal{M}_{-\ell}^{n}$-continuous potential and $\mu\in\mathcal{M}(M,T)$. Since
$\int g\,\mathrm{d}\mu=\sum_{A\in\mathcal{M}_{-\ell}^{n}}\int g\mathbbm{1}_{A}\,\mathrm{d}\mu,$
we only need to prove the continuity of $\mu\mapsto\int g\mathbbm{1}_{A}\,\mathrm{d}\mu$ for any fixed $A\in\mathcal{M}_{-\ell}^{n}$. Notice that, since $\partial A$ is compact, we have $\partial A=\bigcap_{\delta>0}\mathcal{N}_{\delta}(\partial A)$, where $\mathcal{N}_{\delta}(\partial A)$ is a $\delta$-neighbourhood of $\partial A$. Let $\varepsilon>0$. Then, since $\mu(\partial A)=0$, there exists some $\delta>0$ such that $\mu(\mathcal{N}_{\delta}(\partial A))<\varepsilon$. Up to decreasing the value of $\delta$, we can also assume that $\mu(\partial\mathcal{N}_{\delta}(\partial A))=0$. Let $\mu_{k}\rightharpoonup\mu$ as $k\to\infty$. Define
(2.1) $\varphi_{A,\delta}(x)=\begin{cases}1&\text{if }x\in A\smallsetminus\mathcal{N}_{\delta}(\partial A),\\ \delta^{-1}d(x,\partial A)&\text{if }x\in A\cap\mathcal{N}_{\delta}(\partial A),\\ 0&\text{if }x\notin A,\end{cases}$
and notice that $\varphi_{A}$ is continuous. Thus, we need to estimate the three terms in
(2.2) $\begin{split}\left|\int g\mathbbm{1}_{A}\,\mathrm{d}\mu_{k}-\int g\mathbbm{1}_{A}\,\mathrm{d}\mu\right|&\leqslant\left|\int g\mathbbm{1}_{A}\,\mathrm{d}\mu_{k}-\int g\varphi_{A,\delta}\,\mathrm{d}\mu_{k}\right|\\ &\,\,+\left|\int g\varphi_{A,\delta}\,\mathrm{d}\mu_{k}-\int g\varphi_{A,\delta}\,\mathrm{d}\mu\right|+\left|\int g\varphi_{A,\delta}\,\mathrm{d}\mu-\int g\mathbbm{1}_{A}\,\mathrm{d}\mu\right|.\end{split}$
The third term is bounded by $\mu(\mathcal{N}_{\delta}(\partial A))\sup_{A}|g|\leqslant\varepsilon\sup_{A}|g|$. For the first term, we use that since $\mu(\partial\mathcal{N}_{\delta}(\partial A))=0$, there is some $k_{A}$ such that for all $k\geqslant k_{A}$, $\mu_{k}(\mathcal{N}_{\delta}(\partial A))\leqslant\mu(\mathcal{N}_{\delta}(\partial A))+\varepsilon\leqslant 2\varepsilon$. Thus, the first term in (2.2) is bounded by $2\varepsilon\sup_{A}|g|$. Now, for the second term, we use that $g\varphi_{A,\delta}$ is continuous on $M$. Thus, from the $-*{}^{*}$ convergence, for all $k\geqslant k_{A}$ (up to increasing the value of $k_{A}$), the second term in (2.2) is bounded by $\varepsilon$. This prove the continuity of $\mu\mapsto\int g\mathbbm{1}_{A}\,\mathrm{d}\mu$, and hence the continuity of $\mu\mapsto\int g\,\mathrm{d}\mu$. ∎
3. $\mathcal{M}(M,T)$ is closed
Since $T$ is not continuous, the transfer operator $T_{*}$ has no reason, a priori, to be continuous on $\mathcal{M}(M)$ with respect to the $-*{}^{*}$ topology. It is therefore unclear whether the set of fixed points $\mathcal{M}(M,T)$ of $T_{*}$ is a closed set or not. Fortunately, the billiard flow is continuous and since it is a suspension flow over $T$ (because of the finite horizon), invariant measures with respect to $T$ are strongly related to the ones of $\Phi_{t}$. This is the strategy we use.
Proposition 3.1.
For any finite horizon Sinai billiard, the set of $T$-invariant probability measures $\mathcal{M}(M,T)$ is a closed subset of $\mathcal{M}(M)$ for the $-*{}^{*}$ topology.
Proof.
Let $\mu_{n}\in\mathcal{M}(M,T)$ be a sequence converging to some $\mu_{\infty}\in\mathcal{M}(M)$. Then, in the coordinates of the suspension, $\bar{\mu}_{n}\coloneqq(\int_{M}\tau\,\mathrm{d}\mu_{n})^{-1}\mu_{n}\otimes\lambda$ is a sequence of $\Phi_{t}$-invariant probability measures, where $\lambda$ is the Lebesgue measure in the direction of the flow. Up to taking a subsequence, we can assume that $\bar{\mu}_{n}$ converges to some $\bar{\mu}$. Furthermore, since the billiard flow is continuous, $\bar{\mu}$ is a flow invariant measure.
In particular, there exists a $T$-invariant measure $\mu$ such that $\bar{\mu}=(\int_{M}\tau\,\mathrm{d}\mu)^{-1}\mu\otimes\lambda$. We now prove that $\mu_{\infty}=\mu$.
Consider $\psi\in C^{0}(M)$ and a smooth function $\rho:\mathbbm{R}\to\mathbbm{R}_{+}$ supported in $[-\min\tau/2,\min\tau/2]$ such that $\int_{\mathbbm{R}}\rho\,\mathrm{d}\lambda>0$. Consider a fundamental domain containing $M\times[-\min\tau/2,\min\tau/2]$ of $\Omega\cong(M\times\mathbbm{R})/\sim$ where $\sim$ is such that
$(x,t)\sim(y,u)\Leftrightarrow\exists n\geqslant 0,(y,u)=(T^{n}x,t-\sum_{i=0}^{n-1}\tau(T^{i}x))\text{ or }(x,t)=(T^{n}y,u-\sum_{i=0}^{n-1}\tau(T^{i}y)).$
On this domain, $\varphi(x,t)=\psi(x)\rho(t)$ is well defined and descends into a continuous function on $\Omega$, also called $\varphi$. Thus
(3.1) $\frac{\int_{M}\psi\,\mathrm{d}\mu_{n}}{\int_{M}\tau\,\mathrm{d}\mu_{n}}\int_{\mathbbm{R}}\rho\,\mathrm{d}\lambda=\bar{\mu}_{n}(\varphi)\xrightarrow[n\to\infty]{}\bar{\mu}(\varphi)=\frac{\int_{M}\psi\,\mathrm{d}\mu}{\int_{M}\tau\,\mathrm{d}\mu}\int_{\mathbbm{R}}\rho\,\mathrm{d}\lambda.$
Now, since $\mu_{n}(\varphi)$ converges to $\mu_{\infty}(\varphi)$, and $\int_{\mathbbm{R}}\rho\,\mathrm{d}\lambda>0$, we get from (3.1) that for all $\psi\in C^{0}(M)$, if $\mu_{\infty}(\psi)\neq 0$ then
$\frac{\mu(\tau)}{\mu_{n}(\tau)}\xrightarrow[n\to\infty]{}\frac{\mu(\psi)}{\mu_{\infty}(\psi)},$
and if $\mu_{\infty}(\psi)=0$, then $\mu(\psi)=0$. Therefore, there is a constant $c\geqslant\tau_{\min}/\tau_{\max}>0$ such that for all $\psi\in C^{0}(M)$, $\mu(\psi)=c\,\mu_{\infty}(\psi)$. Since both $\mu$ and $\mu_{\infty}$ are probability measures, it must be that $c=1$, and so $\mu_{\infty}=\mu$ is invariant by $T$. Hence, $\mathcal{M}(M,T)$ is a closed set. ∎
4. General estimates
In this section we consider general finite horizon Sinai billiards (with potentially unbounded complexity). We prove the bounds on the defect of upper semi-continuity of the entropy map in Subsection 4.1 and on the topological tail entropy in Subsection 4.2.
4.1. Bound on the defect of upper semi-continuity
Before investigating the smoothness of the entropy map, we give a preliminary result estimating the size of eventual discontinuities of each $\mu\mapsto\frac{1}{n}H_{\mu}(\mathcal{P}_{0}^{n})$.
Proposition 4.1.
For all $\mu\in\mathcal{M}(M,T)$, all small enough $\varepsilon>0$ and all large enough $n$, there is a neighbourhood $V_{\varepsilon,n}\subset\mathcal{M}(M,T)$ of $\mu$ such that for all $\nu\in V_{\varepsilon,n}$,
$\left|\frac{1}{n}H_{\nu}(\mathcal{P}_{0}^{n})-\frac{1}{n}H_{\mu}(\mathcal{P}_{0}^{n})\right|\leqslant\frac{\log\#\mathcal{P}_{0}^{n}}{n}(2(1+\varepsilon)\mu(S_{n+1})+\mu(\mathcal{S}_{0})+3\varepsilon)+\frac{2\log 2+e^{-1}}{n}$
To prove the above proposition, we split the contributions to $H_{\mu}(\mathcal{P}_{0}^{n})$ of a neighbourhood of $\partial\mathcal{P}_{0}^{n}$ and its complement. To do so, we rely on the use of conditional entropy. Recall that for any two finite partitions $\mathcal{A}$ and $\mathcal{C}$, the conditional entropy of $\mathcal{A}$ with respect to $\mathcal{C}$ and $\mu$ is defined by
$H_{\mu}(\mathcal{A}\mid\mathcal{C})\coloneqq\sum_{C\in\mathcal{C}}\mu(C)H_{\mu}(\mathcal{A}\mid C)\quad\text{with }H_{\mu}(\mathcal{A}\mid C)\coloneqq H_{\mu_{C}}(\mathcal{A}),$
where $\mu_{C}$ is the conditional probability of $\mu$ on $C$. By convention, if $\mu(C)=0$, we set $H_{\mu}(\mathcal{A}\mid C)=0$.
In order to estimate the contribution away from singularities, we need the following technical lemma.
Lemma 4.2.
For all $x,y\in[0,1]$ if $|x-y|\leqslant 1/2$, then $|x\log x-y\log y|\leqslant-|x-y|\log|x-y|$.
Proof of Lemma 4.2.
Let $f_{\delta}(x)=-x\log x+(x+\delta)\log(x+\delta)$, for $20\leqslant\delta\leqslant 1/$ 2 and $x\in[0,1-\delta]$. Differentiating, we get $f^{\prime}_{\delta}(x)=\log((x+\delta)/x)>0$ on $(0,1-\delta)$. Thus
$\delta\log\delta=f_{\delta}(0)\leqslant f_{\delta}\leqslant f_{\delta}(1-\delta)=-(1-\delta)\log(1-\delta).$
Let $g(\delta)=-\delta\log\delta+(1-\delta)\log(1-\delta)$, for $20\leqslant\delta\leqslant 1/$ 2. Differentiating twice, we get $g^{\prime\prime}(\delta)=\frac{2\delta-1}{\delta(1-\delta)}<0$ on $(0,1/2)$. Thus $g$ is concave on $[0,1/2]$ and $g(\delta)\geqslant 2\delta g(1/2)+(1-2\delta)g(0)=0$. In other words, $-(1-\delta)\log(1-\delta)\leqslant-\delta\log\delta$. Taking $x\leqslant y$ and $\delta=y-x$ ends the proof. ∎
Proof of Proposition 4.1.
Let $\varepsilon>0$. Remark that for any partition $\alpha$ of $M$ into two elements, we have on the one hand, for all $\nu\in\mathcal{M}(M,T)$ and $n\geqslant 1$,
$H_{\nu}(\mathcal{P}_{0}^{n}\vee\alpha)=H_{\nu}(\alpha)+H_{\nu}(\mathcal{P}_{0}^{n}\mid\alpha),$
where $0\leqslant H_{\nu}(\alpha)\leqslant\log 2$ by [19, Corollary 4.2.1], and on the other hand,
$H_{\nu}(\mathcal{P}_{0}^{n}\vee\alpha)=H_{\nu}(\mathcal{P}_{0}^{n})+H_{\nu}(\alpha\mid\mathcal{P}_{0}^{n}),$
where, $0\leqslant H_{\nu}(\alpha\mid\mathcal{P}_{0}^{n})\leqslant H_{\nu}(\alpha)\leqslant\log 2$ by [19, Theorem 4.3(vi)]. Thus, for any partition $\alpha$ of $M$ with #$\#\alpha=2$ and any $T$-invariant measure $\nu$,
$\left|H_{\nu}(\mathcal{P}_{0}^{n})-H_{\nu}(\mathcal{P}_{0}^{n}\mid\alpha)\right|\leqslant\log 2.$
Fix $n$ such that $2\log 2<n\varepsilon$. Let $\mu\in\mathcal{M}(M,T)$. Let $\delta$ be small enough so that $|\mu(\mathcal{S}_{n+1})-\mu(\mathcal{N}_{\delta}(\mathcal{S}_{n+1}))|\leqslant\varepsilon$. Notice that, up to decreasing the value of $\delta$, we can assume that $\mu(\partial\mathcal{N}_{\delta}(\mathcal{S}_{n+1})\smallsetminus\mathcal{S}_{0})=0$. Now, denote $U=\mathcal{N}_{\delta}(\mathcal{S}_{n+1})$ and consider the partition $\alpha_{\varepsilon,n}=\{U,M\smallsetminus U\}$ into two elements. Therefore, for any $\nu\in\mathcal{M}(M,T)$,
(4.1) $\left|\frac{1}{n}H_{\mu}(\mathcal{P}_{0}^{n})-\frac{1}{n}H_{\nu}(\mathcal{P}_{0}^{n})\right|\leqslant\frac{2\log 2}{n}+\left|\frac{1}{n}H_{\mu}(\mathcal{P}_{0}^{n}\mid\alpha_{\varepsilon,n})-\frac{1}{n}H_{\nu}(\mathcal{P}_{0}^{n}\mid\alpha_{\varepsilon,n})\right|.$
We want to estimate the right hand side. First, notice that since for all $A\in\mathcal{P}_{0}^{n}$, $\partial(A\smallsetminus U)\subset\partial U\smallsetminus\mathcal{S}_{0}$ has $\mu$-measure zero, there is a neighbourhood $V_{A}$ of $\mu$ such that for any $\nu\in V_{A}$,
$\left|\nu(A\smallsetminus U)-\mu(A\smallsetminus U)\right|\leqslant\frac{\varepsilon}{\#\mathcal{P}_{0}^{n}}.$
Also, since $\mu(\partial U\smallsetminus\mathcal{S}_{0})=0$, there is a neighbourhood $\tilde{V}_{1}$ of $\mu$ such that for any $\nu\in\tilde{V}_{1}$,
$\left|\nu(U)-\mu(U)\right|\leqslant(1+\varepsilon)\mu(\mathcal{S}_{0}).$
Finally, since $\mu(\partial(M\smallsetminus U))=0$, there is a neighbourhood $\tilde{V}_{2}$ of $\mu$ such that for any $\nu\in\tilde{V}_{2}$,
$\left|\nu(M\smallsetminus U)-\mu(M\smallsetminus U)\right|\leqslant\varepsilon.$
Define $V_{\varepsilon,n}\coloneqq\tilde{V}_{1}\cap\tilde{V}_{2}\cap\bigcap_{A\in\mathcal{P}_{0}^{n}}V_{A}$. Note that $V_{\varepsilon,n}$ is a neighbourhood of $\mu$.
In the case where $\mu(M\smallsetminus U)>0$, we will need to consider an even smaller neighbourhood. More precisely, notice that if the value of $\varepsilon$ decreases (as well as $\delta$, accordingly), then $\mu(M\smallsetminus U)$ increases. Thus, if $\mu(M\smallsetminus U_{\varepsilon})>0$ for some $\varepsilon$, there is an $\tilde{\varepsilon}>0$ such that for all $0<\varepsilon<\tilde{\varepsilon}$, $22\varepsilon/\mu(M\smallsetminus U)<1/$ 2. Furthermore, there is a neighbourhood $\tilde{V}_{3}$ of $\mu$ such that for all $\nu\in\tilde{V}_{3}$
$\left|\frac{\mu(M\smallsetminus U)}{\nu(M\smallsetminus U)}-1\right|\leqslant\varepsilon.$
If $\mu(M\smallsetminus U)>0$, replace $V_{\varepsilon,n}$ by $V_{\varepsilon,n}\cap\tilde{V}_{3}$ and take $\varepsilon<\tilde{\varepsilon}$.
Expanding the second term of the right-hand-side of (4.1) and assuming $\nu\in V_{\varepsilon,n}$, we get
(4.2) $\begin{split}&\left|H_{\mu}(\mathcal{P}_{0}^{n}\mid\alpha_{\varepsilon,n})-H_{\nu}(\mathcal{P}_{0}^{n}\mid\alpha_{\varepsilon,n})\right|\leqslant\left|\mu(U)H_{\mu}(\mathcal{P}_{0}^{n}\mid U)\right|+\left|\nu(U)H_{\nu}(\mathcal{P}_{0}^{n}\mid U)\right|\\ &\qquad\qquad\qquad\qquad\qquad+\left|\mu(M\smallsetminus U)H_{\mu}(\mathcal{P}_{0}^{n}\mid M\smallsetminus U)-\nu(M\smallsetminus U)H_{\nu}(\mathcal{P}_{0}^{n}\mid M\smallsetminus U)\right|.\end{split}$
We begin with the contributions of $U$. The first term in the right-hand-side of (4.2) can be bounded as follows
$\left|\mu(U)H_{\mu}(\mathcal{P}_{0}^{n}\mid U)\right|\leqslant\mu(U)\log\#\mathcal{P}_{0}^{n}\leqslant(1+\varepsilon)\mu(\mathcal{S}_{n+1})\log\#\mathcal{P}_{0}^{n}.$
Similarly, the second term in (4.2) is estimated as follows
$\left|\nu(U)H_{\nu}(\mathcal{P}_{0}^{n}\mid U)\right|\leqslant((1+\varepsilon)\mu(\mathcal{S}_{n+1})+\mu(\mathcal{S}_{0}))\log\#\mathcal{P}_{0}^{n}$
We now consider the contributions of $M\smallsetminus U$, that is, the third term in (4.2). If $\mu$ is such that $\mu(M\smallsetminus U)=0$, then
$|\mu(M\smallsetminus U)H_{\mu}(\mathcal{P}_{0}^{n}\mid M\smallsetminus U)-$ $\nu(M\smallsetminus U)H_{\nu}(\mathcal{P}_{0}^{n}\mid M\smallsetminus U)|$
$\quad\leqslant\mu(M\smallsetminus U)\left|H_{\mu}(\mathcal{P}_{0}^{n}\mid M\smallsetminus U)-H_{\nu}(\mathcal{P}_{0}^{n}\mid M\smallsetminus U)\right|$
$\qquad+\left|\left(\mu(M\smallsetminus U)-\nu(M\smallsetminus U)\right)H_{\nu}(\mathcal{P}_{0}^{n}\mid M\smallsetminus U)\right|$
$\quad\leqslant\varepsilon\log\#\mathcal{P}_{0}^{n}$
Otherwise, if $\mu(M\smallsetminus U)>0$, then
$\left|\mu(M\smallsetminus U)H_{\mu}(\mathcal{P}_{0}^{n}\mid M\smallsetminus U)-\nu(M\smallsetminus U)H_{\nu}(\mathcal{P}_{0}^{n}\mid M\smallsetminus U)\right|$
$\quad\leqslant\mu(M\smallsetminus U)\left|H_{\mu}(\mathcal{P}_{0}^{n}\mid M\smallsetminus U)-H_{\nu}(\mathcal{P}_{0}^{n}\mid M\smallsetminus U)\right|$
$\qquad+\left|\left(\mu(M\smallsetminus U)-\nu(M\smallsetminus U)\right)H_{\nu}(\mathcal{P}_{0}^{n}\mid M\smallsetminus U)\right|$
$\quad\leqslant\varepsilon\log\#\mathcal{P}_{0}^{n}+\mu(M\smallsetminus U)\sum_{A\in\mathcal{P}_{0}^{n}}\left|-\frac{\mu(A\smallsetminus U)}{\mu(M\smallsetminus U)}\log\frac{\mu(A\smallsetminus U)}{\mu(M\smallsetminus U)}+\frac{\nu(A\smallsetminus U)}{\nu(M\smallsetminus U)}\log\frac{\nu(A\smallsetminus U)}{\nu(M\smallsetminus U)}\right|$
$\quad\leqslant\varepsilon\log\#\mathcal{P}_{0}^{n}+\mu(M\smallsetminus U)\sum_{A\in\mathcal{P}_{0}^{n}}-\left|\frac{\mu(A\smallsetminus U)}{\mu(M\smallsetminus U)}-\frac{\nu(A\smallsetminus U)}{\nu(M\smallsetminus U)}\right|\log\left|\frac{\mu(A\smallsetminus U)}{\mu(M\smallsetminus U)}-\frac{\nu(A\smallsetminus U)}{\nu(M\smallsetminus U)}\right|$
$\quad\leqslant\varepsilon\log\#\mathcal{P}_{0}^{n}+\mu(M\smallsetminus U)S\left(\log\#\mathcal{P}_{0}^{n}-\log S\right)\text{ where }S\coloneqq\sum_{A\in\mathcal{P}_{0}^{n}}\left|\frac{\mu(A\smallsetminus U)}{\mu(M\smallsetminus U)}-\frac{\nu(A\smallsetminus U)}{\nu(M\smallsetminus U)}\right|,$
where we used Lemma 4.2 for the third inequality (the use of Lemma 4.2 is justified by (4.3)), and the Jensen inequality for the last line. Now,
(4.3) $\begin{split}&\sum_{A\in\mathcal{P}_{0}^{n}}\left|\frac{\mu(A\smallsetminus U)}{\mu(M\smallsetminus U)}-\frac{\nu(A\smallsetminus U)}{\nu(M\smallsetminus U)}\right|\\ &\qquad\qquad\quad\leqslant\frac{1}{\mu(M\smallsetminus U)}\sum_{A\in\mathcal{P}_{0}^{n}}\left|\left(\frac{\mu(M\smallsetminus U)}{\nu(M\smallsetminus U)}-1\right)\nu(A\smallsetminus U)\right|+\left|\mu(A\smallsetminus U)-\nu(A\smallsetminus U)\right|\\ &\qquad\qquad\quad\leqslant\frac{1}{\mu(M\smallsetminus U)}\sum_{A\in\mathcal{P}_{0}^{n}}\left(\varepsilon\nu(A\smallsetminus U)+\frac{\varepsilon}{\#\mathcal{P}_{0}^{n}}\right)\leqslant\frac{2\varepsilon}{\mu(M\smallsetminus U)}<\frac{1}{2}.\end{split}$
Thus, in both cases
$\big{|}\mu(M\smallsetminus U)H_{\mu}(\mathcal{P}_{0}^{n}\mid M\smallsetminus U)-\nu(M\smallsetminus U)H_{\nu}(\mathcal{P}_{0}^{n}\mid M\smallsetminus U)\big{|}\leqslant 3\varepsilon\log\#\mathcal{P}_{0}^{n}+e^{-1}.$
Combining the estimates of the contributions of $U$ and $M\smallsetminus U$ inside (4.2), we get
$\big{|}H_{\mu}(\mathcal{P}_{0}^{n}\mid\alpha_{\varepsilon,n})-H_{\nu}(\mathcal{P}_{0}^{n}\mid\alpha_{\varepsilon,n})\big{|}\leqslant\left(2(1+\varepsilon)\mu(\mathcal{S}_{n+1})+\mu(\mathcal{S}_{0})+3\varepsilon\right)\frac{\log\#\mathcal{P}_{0}^{n}}{n}+\frac{e^{-1}}{n}.$
Then, inserting this last estimate inside (4.1), we obtain the desired claim. ∎
We are now able to prove the main bound of this section.
Corollary 4.3.
For all finite horizon Sinai billiard and all $T$-invariant probability measure $\mu$,
$\limsup_{\nu\rightharpoonup\mu}h_{\nu}(T)\leqslant h_{\mu}(T)+\left(2\mu\left(\bigcup_{n\geqslant 0}\mathcal{S}_{n}\right)+\mu(\mathcal{S}_{0})\right)h_{*},$
[…]
Proof.
Let $\mu\in\mathcal{M}(M,T)$ and $\varepsilon>0$ small enough as in Proposition 4.1. Since $\mathcal{P}$ is a generating partition, we can choose $n=n(\varepsilon)$ large enough so that
$h_{\mu}(T)\geqslant\frac{1}{n}H_{\mu}(\mathcal{P}_{0}^{n})-\varepsilon.$
Let $V_{\varepsilon,n}$ be the neighbourhood of $\mu$ given by Proposition 4.1 (up to increasing the value of $n$). Therefore
$h_{\mu}(T)$ $\geqslant\sup_{\nu\in V_{\varepsilon,n}}\frac{1}{n}H_{\nu}(\mathcal{P}_{0}^{n})-\left(\frac{\log\#\mathcal{P}_{0}^{n}}{n}(2(1+\varepsilon)\mu(S_{n+1})+\mu(\mathcal{S}_{0})+3\varepsilon)+\frac{2\log 2}{n}\right)-\varepsilon$
$\geqslant\sup_{\nu\in V_{\varepsilon,n}}h_{\nu}(T)-\left(\frac{\log\#\mathcal{P}_{0}^{n}}{n}(2(1+\varepsilon)\mu(S_{n+1})+\mu(\mathcal{S}_{0})+3\varepsilon)+\frac{2\log 2}{n}\right)-\varepsilon.$
We can choose $n=n(\varepsilon)$ such that, taking the limit $\varepsilon\to 0$, we get $n(\varepsilon)\to+\infty$. Therefore, taking the limit as $\varepsilon$ goes to zero of the previous expression yields
$h_{\mu}(T)\geqslant\limsup_{\nu\rightharpoonup\mu}h_{\nu}(T)-\left(2\mu\left(\bigcup\limits_{n\geqslant 0}\mathcal{S}_{n}\right)+\mu(\mathcal{S}_{0})\right)h_{*}\,,$
as claimed. ∎
4.2. Bound on the topological tail entropy
In this section, we give an estimate on the topological tail entropy, as defined in (1.4), in Corollary 4.7. To do so, we split the contributions of points in $M\smallsetminus\mathcal{S}_{n}$ and $\mathcal{S}_{n}$ in Propositions 4.4 and 4.5 respectively.
First, recall that in the proof of [3, Lemma 3.4], Baladi and Demers proved that [^1] The definition of the Bowen metric $d_{n}$ in [3] corresponds here to $d_{n+1}$. there exists $\varepsilon_{0}>0$ with the property that for any $n\geqslant 0$, if $x,y$ lie in different elements of $\mathcal{M}_{0}^{n}$, then $d_{n+1}(x,y)\geqslant\varepsilon_{0}$.
Proposition 4.4.
There exists a constants $C>0$ such that for all $0<\varepsilon<\varepsilon_{0}$, all $\delta>0$, all $n\geqslant 2$,
$\log r(n,\delta,B(x,\varepsilon,n))\leqslant 2\frac{\log\#\mathcal{P}}{\log\Lambda}\log\delta^{-1}+C,\quad\forall x\in M\smallsetminus\mathcal{S}_{n-1}.$
In particular, taking the appropriate limits in $n$, $\delta$ and $\varepsilon$, notice that points of $M\smallsetminus\bigcup_{n\geqslant 0}\mathcal{S}_{n}$ don’t contribute to $h^{*}(T)$.
Proof.
Let $x\in M\smallsetminus\mathcal{S}_{n-1}$, $\varepsilon<\varepsilon_{0}$ and $\delta>0$. Let $\mathcal{M}_{0}^{n-1}(x)$ denote the unique element of $\mathcal{M}_{0}^{n-1}$ containing $x$. Then, $B(x,\varepsilon,n)\subset\mathcal{M}_{0}^{n-1}(x)$.
Note that if $F^{\prime}\subset B(x,\varepsilon,n)$ is $(\delta/2,n)$-separated of maximal cardinality, then $F^{\prime}$ is $(\delta,n)$-spanning $B(x,\varepsilon,n)$. In particular, $r(\delta,n,B(x,\varepsilon,n))\leqslant\#F^{\prime}$. Furthermore, if $F\subset\mathcal{M}_{0}^{n-1}(x)$ is $(\delta/2,n)$-separated with maximal cardinality, then $F$ is $(\delta,n)$-spanning and #$\#F^{\prime}\leqslant\#F$. Let $F$ be such a set.
Since $F$ is $(\delta/2,n)$-separated, according to the proof of [3, Lemma 3.4], each element of $\mathring{\mathcal{P}}_{-k_{\delta}}^{n-1+k_{\delta}}$ contains at most one point of $F$, where $k_{\delta}$ is the smallest integer such that $C_{1}\Lambda^{-k_{\delta}}<c_{1}\delta/2$, where $c_{1}$ and $C_{1}$ are constants coming respectively from [3, Lemma 3.4] and [3, eq. (3.1)]. Since $\mathcal{M}_{0}^{n-1}(x)$ contains at most #$\#\mathcal{M}_{-1}^{0}$ elements of $\mathcal{M}_{-1}^{n-1}$ and, from [3, Lemma 3.3], $\mathcal{M}_{-1}^{n-1}=\mathring{\mathcal{P}}_{0}^{n-2}$, we get that $\mathcal{M}_{0}^{n-1}(x)$ contains at most #$\#\mathcal{M}_{-1}^{0}\#\mathcal{P}_{-k_{\delta}}^{0}\#\mathcal{P}_{0}^{k_{\delta}+1}$ elements of $\mathring{\mathcal{P}}_{-k_{\delta}}^{n-1+k_{\delta}}$. Therefore
$r(n,\delta,B(x,\varepsilon,n))\leqslant\#F\leqslant\#\mathcal{M}_{-1}^{0}(\#\mathcal{P})^{2k_{\delta}+1}\leqslant\#\mathcal{M}_{-1}^{0}\#\mathcal{P}\exp\left(-2\log\#\mathcal{P}\left\lceil\frac{\log c_{1}C_{1}^{-1}\delta}{\log\Lambda}\right\rceil\right).$
Taking the $\log$ finishes the proof. ∎
Proposition 4.5.
There exists a constant $C>0$ such that for all $0<\varepsilon<\varepsilon_{0}$ small enough, all $\delta>0$ and all $n\geqslant n_{0}$,
$\log r(n,\delta,B(x,\varepsilon,n))\leqslant 2\frac{\log\#\mathcal{P}}{\log\Lambda}\log\delta^{-1}+C+\left(3+2\left\lfloor\frac{\max\tau}{\min\tau}\right\rfloor\right)(n+n_{0})s_{0}\log 2K,\quad\forall x\in\mathcal{S}_{n-1},$
where $K$ is such that $K_{n}\leqslant Kn$ for all $n$.
Actually, the estimate in Proposition 4.5 holds for all $x\in M$, but in lights of Proposition 4.4 we only state it for points of $\mathcal{S}_{n-1}$.
To prove the above, we rely on a following result which allows to count the number of times a flow orbit passes near a scatterer without making a collision.
Lemma 4.6.
For all $n_{0}$ fixed, there exists $\tilde{\varepsilon}=\tilde{\varepsilon}(n_{0})$ such that for all $\varepsilon\in(0,\tilde{\varepsilon})$ and all $x\in M$, the curve $\{\Phi_{t}(x)\mid t\in[0,\sum_{i=0}^{n_{0}-1}\tau(T^{i}x)]\}\subset\Omega$ passes at distance less than $\varepsilon$ to the scatterers without making a collision at most $\lfloor\max\tau/\min\tau\rfloor s_{0}n_{0}$ times.
Proof of Lemma 4.6.
Denote $t_{0}\coloneqq\lfloor n_{0}\max\tau/\min\tau\rfloor$. First, note that $(x,t)\mapsto\Phi_{t}(x)$ is continuous on the compact set $\Omega\times[-t_{0}\max\tau,t_{0}\max\tau]$ and, therefore, is uniformly continuous on this same set. The idea of the proof is to find for all $x$ a point $y$ in a neighbourhood such that the flow orbit of $y$ makes a grazing collision on the scatterers getting close to the trajectory of $x$. We start by considering the points making multiple grazing collisions. Let
$R_{2}\coloneqq\{x\in M\mid(T^{i}x)_{0\leqslant i\leqslant t_{0}}\text{ comprises at least two grazing collisions}\}.$
Remark that each point of $R_{2}$ must be a corner of $\bar{A}$, for some $A\in\mathcal{M}_{0}^{t_{0}}$. In particular $R_{2}$ is a finite set. Informally, define $\varepsilon_{2}$ to be the minimal distance (in $\Omega$) between the scatterers and the segment $[T^{i}x,T^{i+1}x]\subset\Omega$, $x\in R_{2}$ and $0\leqslant i<t_{0}$. More precisely, define $\pi:\Omega\to Q$ by $\pi(x)=\pi(\bar{x},v)=\bar{x}$ where $x=(\bar{x},v)\in\Omega=Q\times\mathbbm{S}^{1}/\sim$, and
$\tilde{d}([x,Tx],\partial Q)\coloneqq\inf\{\delta>0\mid\partial Q\cap(\pi([x,Tx])\times[-\delta,\delta])\text{ has $3$ or more connected components}\}$
Therefore, let
$\varepsilon_{2}\coloneqq\min_{x\in R_{2}}\min_{0\leqslant i<t_{0}}\tilde{d}([T^{i}x,T^{i+1}x],\partial Q)>0,$
and let $\delta_{2}$ be the delta associate to $\varepsilon_{2}/2$ by the uniform continuity of the flow.
We now consider the points making a single grazing collision: let
$R_{1}\coloneqq\{x\in M\mid(T^{i}x)_{0\leqslant i\leqslant t_{0}}\text{ comprises exactly one grazing collisions}\}=\mathcal{S}_{t_{0}}\smallsetminus R_{2}.$
Then $R_{1}^{\prime}\coloneqq\mathcal{S}_{t_{0}}\smallsetminus\mathcal{N}_{\delta_{2}/2}(R_{2})$ is compact. Let
$\varepsilon_{1}\coloneqq\min_{x\in R_{1}^{\prime}}\min_{0\leqslant i<t_{0}}\tilde{d}([T^{i}x,T^{i+1}x],\partial Q).$
By the uniform continuity of the flow, it follows that $\varepsilon_{1}>0$. Let $\delta_{1}>0$ be the delta associated with $\min(\varepsilon_{2}/2,\varepsilon_{1}/2,\delta_{2}/2,\mbox{diam}R_{2}/4)$ by the uniform continuity of the flow. Up to decreasing the value of $\delta_{1}$, we can assume that $\delta_{1}<\min(\varepsilon_{1}/2,\varepsilon_{2}/2)$. Finally, let $\delta_{0}$ be the delta associated to $\delta_{1}$ by the uniform continuity of the flow. We can assume that $\delta_{0}<\delta_{1}$.
We claim that any $\varepsilon>0$ such that $\varepsilon<\delta_{0}$ has the desired property. Indeed, assume that for some fixed $x\in M$ there exist two or more distinct $0\leqslant i_{j}<n_{0}$ such that $\tilde{d}([T^{i_{j}}x,T^{i_{j}+1}x],\partial Q)<\varepsilon$. Therefore, to each $i_{j}$ is associated a point $x_{j}\in M$, in a $\delta_{1}$-neighbourhood of $x$, such that the flow trajectories of $x_{j}$ and $x$ stay at distance at most $\min(\varepsilon_{2}/2,\varepsilon_{1}/2,\delta_{2}/2,\mbox{diam}R_{2}/4)$ for all times $t\in[0,n_{0}\max\tau]$. By this shadowing, we get that the orbit of $x_{j}$ passes at distance less than $\delta_{0}+\varepsilon_{1}/2<\varepsilon_{1}$ to $\partial Q$ (corresponding to the in-between collisions associated with $i_{k}$, $k\neq j$). In particular, by the definition of $\varepsilon_{1}$, we get that $x_{j}\notin R_{1}^{\prime}$. Thus $d_{M}(x_{j},R_{2})<\delta_{2}/2$ for all $j$. Since $d_{M}(x,x_{j})<\delta_{1}\leqslant\mbox{diam}R_{2}/4$, there exists a unique $y\in R_{2}$ such that $d_{M}(x_{j},y)<\delta_{2}/2$. Therefore $d(x,y)<\delta_{2}$. Thus, the flow orbits of $x$ and $y$ stay at distance less than $\varepsilon_{2}/2$ to each other for times $t\in[0,\sum_{i=0}^{n_{0}-1}\tau(T^{i}x)]$. Thus, the flow orbit of $y$ passes at distance less than $\varepsilon+\varepsilon_{2}/2<\varepsilon_{2}$ to $\partial Q$ when the orbit of $x$ passes at distance less than $\varepsilon$. Therefore, by the definition of $\varepsilon_{2}$, the flow orbit of $y$ makes a grazing collision at this instants. However, the flow orbit of $y$, for $t\in[0,n_{0}\max\tau]$ makes at most $\lfloor\max\tau/\min\tau\rfloor n_{0}$ collisions, and thus at most $\lfloor\max\tau/\min\tau\rfloor s_{0}n_{0}$ grazing collisions. Taking $\tilde{\varepsilon}=\delta_{0}$ finishes the proof. ∎
We have now all the tools in order to prove Proposition 4.5.
Proof of Proposition 4.5.
Let $0<\varepsilon<\varepsilon_{0}$ be small enough so that $\sin\varepsilon\leqslant\tilde{\varepsilon}/\max\tau$ (where $\tilde{\varepsilon}$ comes from Lemma 4.6),
$A\cap\mathcal{N}_{\varepsilon}(A\cap(\mathcal{S}_{-1}\cup\mathcal{S}_{1}))\subset\bigcup_{-1\leqslant i\leqslant 1}T^{i}\{|\varphi|>\varphi_{0}\},\quad\forall A\in\mathcal{P},$
and $B(x,\varepsilon)$ intersects at most $2K$ elements of $\mathcal{M}_{-1}^{1}$ for all $x\in M$.
Let $x\in M$. Then $B(x,\varepsilon,n)$ might intersects several elements of $\mathcal{M}_{-1}^{n}=\mathring{\mathcal{P}}_{0}^{n-1}$. We start by estimate the number of those sets.
Assume first that $n=n_{0}$. For any $A\in\mathcal{P}_{0}^{n-1}$, there exist $A_{i}\in\mathcal{P}$ such that $A=\bigcap_{i=0}^{n-1}T^{-i}A_{i}$. Thus $A\cap B(x,\varepsilon,n)\neq\emptyset$ if and only if $A_{i}\cap B(T^{-i}x,\varepsilon)\neq\emptyset$ for all $0\leqslant i<n$. In order to count the number of admissible $A_{i}$, we distinguish three mutually exclusive cases:
i) $B(y,\varepsilon)\cap(\mathcal{S}_{-1}\cup\mathcal{S}_{1})=\emptyset$. Therefore, $B(y,\varepsilon)$ is contained is a single element of $\mathcal{P}$.
ii) $y\in A\cap\mathcal{N}_{\varepsilon}(A\cap(\mathcal{S}_{-1}\cup\mathcal{S}_{1}))$ for some $A\in\mathcal{P}$. Therefore at least one of $y$, $Ty$ or $T^{-1}y$ is a $\varphi_{0}$-grazing collision, and $B(y,\varepsilon)$ intersects at most $2K$ elements of $\mathcal{P}$. This case occurs at most $3s_{0}n_{0}$ times.
iii) Otherwise, at least one of $[y,Ty]$ or $[T^{-1}y,y]$ passes $\varepsilon$ close to the scatterers without making a collision. Furthermore, $B(y,\varepsilon)$ intersects at most $2K$ elements of $\mathcal{P}$. By the Sublemma 4.6, this case occurs at most $2\lfloor\max\tau/\min\tau\rfloor s_{0}n_{0}$ times.
Combining the three above situations, we obtain
#$\#\{A\in\mathcal{P}_{0}^{n_{0}-1}\mid A\cap B(x,\varepsilon,n_{0})\neq\emptyset\}\leqslant(2K)^{\left(3+2\left\lfloor\frac{\max\tau}{\min\tau}\right\rfloor\right)s_{0}n_{0}}.$
Now, for general $n\geqslant n_{0}$, write $n=kn_{0}+\ell$, $0\leqslant\ell<n_{0}$. Then, any $A\in\mathcal{P}_{0}^{n-1}$ can be written $A=\bigcap_{i=0}^{k-1}T^{in_{0}}A_{i}\cap B$, where $A_{i}\in\mathcal{P}_{0}^{n_{0}-1}$ and $B\in\mathcal{P}_{0}^{\ell-1}$. Thus, $A\cap B(x,\varepsilon,n)\neq\emptyset$ if and only if $A_{i}\cap B(T^{in_{0}}x,\varepsilon,n_{0})\neq\emptyset$ for all $i$ and $B\cap B(T^{kn_{0}}x,\varepsilon,\ell)\neq\emptyset$. Therefore, by the computation done in the case $n=n_{0}$, and since $\mathcal{P}_{0}^{n-1}=\mathcal{M}_{-1}^{n}$, we obtain
#$\#\{A\in\mathcal{M}_{0}^{n}\mid A\cap B(x,\varepsilon,n)\neq\emptyset\}\leqslant\#\mathcal{M}_{-1}^{0}\left((2K)^{\left(3+2\left\lfloor\frac{\max\tau}{\min\tau}\right\rfloor\right)s_{0}n_{0}}\right)^{\left\lceil\frac{n}{n_{0}}\right\rceil}.$
For each $A\in\mathcal{M}_{0}^{n-1}$, let $F_{A}\subset A$ be $(\delta/2,n)$-separated with maximal cardinality if $A\cap B(x,\varepsilon,n)\neq\emptyset$, and $F_{A}=\emptyset$ otherwise. Thus, the union of the $F_{A}$ is $(\delta,n)$-spanning $B(x,\varepsilon,n)$. Therefore, using that
#$\#\{A\in\mathcal{M}_{0}^{n-1}\mid A\cap B(x,\varepsilon,n)\neq\emptyset\}\leqslant\#\{A\in\mathcal{M}_{0}^{n}\mid A\cap B(x,\varepsilon,n)\neq\emptyset\},$
and estimates on each #$\#F_{A}$ obtained in the proof of Proposition 4.4 yield
$r(n,\delta,B(x,$ $\varepsilon,n))\leqslant\sum_{A\in\mathcal{M}_{0}^{n-1}}\#F_{A}$
(#$\leqslant(\#\mathcal{M}_{-1}^{0})^{2}\#\mathcal{P}\exp\left(-2\log\#\mathcal{P}\left\lceil\frac{\log c_{1}C_{1}^{-1}\delta}{\log\Lambda}\right\rceil\right)\left((2K)^{\left(3+2\left\lfloor\frac{\max\tau}{\min\tau}\right\rfloor\right)s_{0}n_{0}}\right)^{\left\lceil\frac{n}{n_{0}}\right\rceil}.$
Taking the $\log$ of the above estimate ends the proof. ∎
Corollary 4.7.
For any finite horizon billiard table, $h^{*}(T)\leqslant\left(3+2\left\lfloor\frac{\max\tau}{\min\tau}\right\rfloor\right)s_{0}\log 2K$.
Proof.
By Propositions 4.4 and 4.5, we get that
$\limsup_{n\to\infty}\frac{1}{n}\log\sup_{x\in M}r(n,\delta,B(x,\varepsilon,n))\leqslant\left(3+2\left\lfloor\frac{\max\tau}{\min\tau}\right\rfloor\right)s_{0}\log 2K.$
Taking the appropriate limits in $\delta$ and in $\varepsilon$ yields the result. ∎