### v2

# A direct proof of the existence of MME for finite horizon Sinai billiards

Abstract.
The Sinai billiard map $T$ on the two-torus, i.e., the periodic Lorentz gaz, is a discontinuous map. Assuming finite horizon and another condition we introduce – namely negligible singularities – we prove that the metric pressure map associated with the billiard map $T$ is upper semi-continuous, as well as the compactness of the set of $T$-invariant measures. In particular, for the potentials $g\equiv 0$ and $g=-h_{\rm top}(\Phi^{1})\tau$, we recover the recent results of the existence of measures of maximal entropy (MME) for both the billiard map and flow $\Phi^{t}$, due to Baladi and Demers for $T$, jointly with Carrand for $\Phi^{t}$. We prove that the negligible singularities condition is generic among the billiard table with $C^{3+\alpha}$ boundary, with respect to the $C^{3+\alpha}$ topology. For finite horizon Sinai billiards, we provide bounds on the defect of upper semi-continuity of the metric pressure map and on the topological tail entropy. Assuming that singularities are negligible and sparse recurrence hold, we deduce the equidistribution of the periodic orbits with respect to the unique MME. We provide examples of billiard tables satisfying both conditions.
The author is grateful to Viviane Baladi, Mark Demers, David Burguet and Yuri Lima for useful discussions and comments. Research supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 787304).
1. Introduction
1.1. Equilibrium States in Dynamical Systems
Taking inspiration from physics, more specifically thermodynamic, where “nature tends to maximize free energy”, it is tempting to investigate the existence of maximizing measures (or states in physics terminology) for
(1.1) $\mu\mapsto P_{\mu}(T,g)\coloneqq h_{\mu}(T)+\int_{M}g\,{\rm d}\mu\,,$
where $\mu$ ranges over the set $\mathcal{M}(M,T)$ of Borel probability measures on $M$ invariant under some transformation $T:M\to M$, $h_{\mu}(T)$ is the Kolmogorov-Sinai entropy of $(T,\mu)$, and $g:M\to\mathbbm{R}$ is some fixed function called potential – in the physical analogy, $\int_{M}-g\,\mathrm{d}\mu$ corresponds to the internal energy of the state $\mu$. Maximal measures are called equilibrium states.
This problem has been introduced by Ruelle in the case of finite rank lattices [28, 29], where a particle lies at each vertex. The rank one case, assuming translation symmetry of states, corresponds to a subshift of finite type. There, for any smooth potential $g$ (e.g. Hölder), there exists a unique equilibrium state $\mu_{g}$, and it enjoys strong statistical properties (exponential mixing, CLT) [7]. Furthermore, $\mu_{g}$ can be constructed by joining the left and right eigenvectors with maximal eigenvalue of a twisted Ruelle transfer operator.
More general transformations can be studied: through the use of coding with Markov partitions, the case of Anosov or Axiom A maps $T$ can be carried out [7]. In this case, for any Hölder potential, the uniquely associated equilibrium state is Bernoulli. Note that the process of coding can be bypassed by studying a twisted Ruelle transfer operator acting this time on anisotropic Banach spaces [1]. Although this strategy is usually very technical, it can lead to strong statistical properties (exponential mixing, CLT, ASIP) – provided that the operator has, for example, good spectral properties – and has the advantage to be applicable to more general situations.
Finally, note that if $T$ is uniformly hyperbolic on a compact manifold $M$, then it is expansive [23, Corollary 6.4.10], which in turn implies that $\mu\mapsto h_{\mu}(T)$ is upper semi-continuous [33, Theorem 8.2], giving existence for $\mu_{g}$ for any continuous potential $g$ since $\mathcal{M}(M,T)$ is compact. For continuous maps $T$, the defect of upper semicontinuity can be bounded by the topological tail entropy $h^{*}(T)$ [26]. Uniqueness can be deduced from expansiveness and the so called specification property [6].
1.2. The case of Sinai Billiards
In this paper we are interested in the case where $T:M\to M$ is the collision map associated to a finite horizon Sinai billiard (see definition below).
A Sinai billiard on the torus is (the quotient modulo $\mathbbm{Z}^{2}$, for position, of) the periodic planar Lorentz gas (1905) model for the motion of a single dilute electron in a metal. The scatterers (corresponding to atoms of the metal) are assumed to be strictly convex (but not necessarily discs). Such billiards have become fundamental models in mathematical physics.
Assuming finite horizon (see definition below), many equilibrium states for the billiard map (resp. for the billiard flow) have already been studied, such as the SRB measure [12] which enjoys exponential mixing [34, 16] (resp. [5]), the measure of maximal entropy, constructed in [3] (resp. [11, 2]) enjoys polynomial mixing [15], equilibrium states associated to piecewise-Hölder potentials [11, 2] which are Bernoulli, and finally equilibrium states with geometric potentials (which are not Hölder) [4] which are exponentially mixing. Most of the constructions – those of [3, 11, 2, 4] – or the study of the equilibrium states – as for [16, 5] – relie on the use of a transfer operator acting on specific anisotropic Banach spaces. In [2, 3, 4, 11, 25], the studied equilibrium states are proved to be unique. In the construction performed in [3, 11, 2], an additional assumption on the billiard table is made, namely sparse recurrence to singularities, that is
(1.2) $P_{\rm top}(T,g)-\sup g>s_{0}\log 2,$
where $P_{\rm top}(T,g)$ is the topological pressure of $g$ under the transformation $T$, and $s_{0}\in(0,1)$ is defined in the next subsection.
Bounded complexity is a more restrictive assumption than sparse recurrence (used in [3, 2]). Indeed, for $g\equiv 0$ (respectively, for $g=-h_{\rm top}(\Phi^{1})\tau$, where $\tau$ is the time of first return map), the left hand side of (1.2) reduces to $P_{\rm top}(T,g)-\sup g=h_{\rm top}(T)>0$ (resp. $P_{\rm top}(T,g)-\sup g=h_{\rm top}(\Phi^{1})\min\tau>0$), but Demers and Korepanov proved in [15] that if the complexity is bounded, then $\inf_{n_{0},\varphi_{0}}s_{0}(n_{0},\varphi_{0})=0$. According to [31], bounded complexity is a generic property among finite horizon Sinai billiards.
We say that a billiard table has negligible singularities when its singular set is a null set with respect to every invariant measure. Using again the result due to Demers and Korepanov, this condition is less restrictive than bounded complexity. It is unclear how it relates to sparse recurrence.
The first half of this paper is dedicated to prove that for for generic billiard tables with finite horizon, its pressure map is upper semi-continuous for any piecewise continuous potential. In Section 2, we show that for any finite horizon Sinai billiard with negligeable singularities and any piecewise continuous potential, the map (1.1) is upper semi-continuous. In Section 3, we prove that for any finite horizon Sinai billiard, $\mathcal{M}(M,T)$ is compact. In particular for $g\equiv 0$ and $g=-h_{\rm top}(\Phi^{1})\tau$ (see [11]), we recover the existence of the MME for the collision map and the billiard flow respectively. Furthermore, we prove that the variational principle is satisfied. In Section 4 we prove that the negligible singularities condition is generic with respect to the $C^{3+\alpha}$ topology.
In the second half of the paper, we study the defect of upper semi-continuity of the pressure map for general finite horizon billiard tables as well as the topological tail entropy. In Section 5, we provide a bound the defect of upper semi-continuity, similar to the ones obtained in [21, 17, 22, 9, 32] in other contexts, as well as its consequences, such that the equidistribution of the periodic orbits with respect to the MME (assuming both sparse recurrence and negligible singularities). In Section 6 we provide an upper bound on the topological tail entropy of the billiard map. Finally, in Section 7, we give some examples of billiard tables satisfying both sparse recurrence and negligible singularities.
The problem of uniqueness is not covered by this paper. In [25], it is shown – without requiring sparse recurrence – that for any fixed piecewise Hölder potential, among all associated equilibrium states, at most one is adapted.
1.3. Precise statement of results
Formally, a Sinai billiard table $Q$ on the two-torus $\mathbbm{T}^{2}$ is a set $Q=\mathbbm{T}^{2}\smallsetminus B$, with $B=\bigcup_{i=1}^{D}B_{i}$ for some finite number $D\geqslant 1$ of pairwise disjoint closed domains $B_{i}$, called scatterers, with $\mathcal{C}^{3}$ boundaries having strictly positive curvature – in particular, the scatterers are strictly convex. The billiard flow $\Phi^{t}:\Omega\to\Omega$ is the motion of a point particle travelling at unit speed in $Q$ with specular reflections off the boundary of the scatterers. Identifying outgoing collisions with incoming ones in the phase space, that is $\Omega=Q\times\mathbbm{S}^{1}/\sim$, the billiard flow is continuous. However, the grazing collisions – those tangential to scatterers – give rise to singularities in the derivative [12]. The Sinai billiard map $T:M\to M$ – also called collision map – is the first return map of the single point particle to the scatterers, that is $M=\partial Q\times[-\pi/2,\pi/2]$. Because of the grazing collisions, the Sinai billiard map is a discontinuous map.
As in [3, 34], we say that the billiard table $Q$ has finite horizon if the billiard flow does not have any trajectories making only tangential collisions – in particular, this implies that the return time function $\tau$ to the scatterers is bounded. All Sinai billiards in this paper are assumed to have finite horizon. In this case, $\Phi^{t}$ is a suspension flow over $T$, under $\tau$.
Define $\mathcal{S}_{0}=\{(r,\varphi)\in M\mid|\varphi|=\pi/2\}$ the set of grazing collisions, $\mathcal{S}_{\pm n}=\bigcup_{i=0}^{n}T^{\mp i}\mathcal{S}_{0}$ the singular set of $T^{\pm n}$, and $\mathcal{S}=\bigcup_{i=-\infty}^{+\infty}$ the set of all singularities. Notice that $\mathcal{S}_{n}$ (resp. $\mathcal{S}_{-n}$) is made of a finite number of decreasing (resp. increasing) curves in the $(r,\varphi)$ coordinates, whereas $M$ is a two dimensional manifold with boundary $\mathcal{S}_{0}$. Let $\mathcal{M}_{-k}^{n}$ be the partition of $M\smallsetminus(\mathcal{S}_{-k}\cup\mathcal{S}_{n})$ into maximal connected components (which are all open sets of $M$). Furthermore, define $\mathcal{P}$ the partition of $M$ into maximal connected components on which both $T$ and $T^{-1}$ are continuous, and $\mathcal{P}_{-k}^{n}=\bigvee_{i=-k}^{n}T^{-i}\mathcal{P}$. As in [3], the topological entropy of $T$ is defined as $h_{*}\coloneqq\lim_{n}\tfrac{1}{n}\log\#\mathcal{P}_{0}^{n}$. We say that a potential $g$ is $\mathcal{M}_{-k}^{n}$-continuous if $g$ is bounded and $g|_{A}$ is continuous for each $A\in\mathcal{M}_{-k}^{n}$. We simply say that a potential is piecewise continuous if it is $\mathcal{M}_{-k}^{n}$-continuous for some $k,n\geqslant 0$.
The complexity of a billiard is a sequence $(K_{n})_{n\geqslant 1}$, where each $K_{n}$ is defined as the maximal number of curves of $S_{n}$ intersecting at some common point. Chernov [13] proved that for any finite horizon Sinai billiard, there exists a constant $K$ such that $K_{n}\leqslant K\,n$. A linear growth of the complexity can be achieved if there exists a periodic orbit with a grazing collision. In a recent seminar [31], Tòth announced that for generic billiard table, the complexity is bounded.
In this paper, we introduce a weaker condition, namely Negligible Singularities. A billiard table satisfies this condition whenever $\mu(\mathcal{S})=0$ for all $T$-invariant measure $\mu$.
The first half of the present paper – that is, Sections 23 and 4 – is devoted to the following result.
Theorem 1.1.
For any finite horizon Sinai billiard with negligible singularities, and any piecewise-continuous potential $g:M\to\mathbbm{R}$, the map
$\mu\in\mathcal{M}(M,T)\mapsto h_{\mu}(T)+\int g\,{\rm d}\mu$
is upper semi-continuous, and $\mathcal{M}(M,T)$ is compact with respect to the weak- topology. Thus equilibrium states exist and they all satisfy the variational principle.
In particular, $g\equiv 0$ and $g=-h_{\rm top}(\phi_{1})\tau$ admit equilibrium measures. In other words, $T$ and $\Phi^{t}$ have MME.
Furthermore, finite horizon Sinai billiard table with negligible singularities form a dense $G_{\delta}$ set within the set of finite horizon tables with respect to the $C^{3}$ topology.
The upper semi-continuity statement in the above theorem follows from Proposition 2.2. The compactness of $\mathcal{M}(M,T)$ is proved in Proposition 3.1. The relation between $g=-h_{\rm top}(\phi_{1})\tau$ and the MME of the Sinai billiard flow is proved in [11, Corollary 2.6]. The genericness is proved in Proposition 4.6.
In order to state our second result, we need to quantify the frequency of return of an orbit in a neighbourhood of $\mathcal{S}_{0}$. More precisely, for any $n_{0}\geqslant 1$ and $20<\varphi_{0}<\pi/$ 2, define
(1.3) $s_{0}(n_{0},\varphi_{0})\coloneqq\frac{1}{n_{0}}\sup_{M}\sum_{i=0}^{n_{0}-1}\mathbbm{1}_{(|\varphi|>\varphi_{0})}\circ T^{i},$
that is, the maximal frequency of $\varphi_{0}$-grazing collisions among $n_{0}$ consecutive collisions.
Finally, we recall the formal definition of the topological tail entropy (as in [8]). Let $d(x,y)$ be the Euclidean metric between $x$ and $y$ if they belong to the same connected component $M_{i}$ of $M=\bigsqcup_{i=1}^{D}M_{i}$, and otherwise fix $d(x,y)=10\max_{1\leqslant i\leqslant D}\mbox{diam}(M_{i})$. Denote $d_{n}$ the $n$-th Bowen distance, that is
$d_{n}(x,y)\coloneqq\max_{0\leqslant i<n}d(T^{i}x,T^{i}y),$
and $B(x,\varepsilon,n)$ the open ball centred at $x$ of radius $\varepsilon$ for the $d_{n}$ metric. A set $X\subset Y$ ($Y\subset M$) is said to be a $(n,\delta)$-spanning subset of $Y$ if $Y\subset\bigcup_{x\in X}B(x,\delta,n)$ and a $(n,\delta)$-separated subset of $Y$ if for all distinct $x,y\in X$, $d_{n}(x,y)\geqslant\delta$. We use the notation $r(n,\delta,Y)$ (respectively, $s(n,\delta,Y)$) for the minimal cardinality of a $(n,\delta)$-spanning subset of a set $Y\subset M$ (respectively, the maximal cardinality of a $(n,\delta)$-separated subset of $Y$). Now, define the topological tail entropy to be
(1.4) $h^{*}(T)\coloneqq\lim_{\varepsilon\to 0}\lim_{\delta\to 0}\limsup_{n\to+\infty}\frac{1}{n}\sup_{x\in M}r(n,\delta,B(x,\varepsilon,n)).$
The second half of the paper, namely Sections 5 and 6, is devoted to the proof of the following result.
Theorem 1.2.
For any finite horizon Sinai billiard table:
i) $\limsup_{\nu\rightharpoonup\mu}P_{\nu}(T,g)\leqslant P_{\mu}(T,g)+\mu(\mathcal{S})\left(P_{\rm top}(T,g)-P_{\mu_{\mathcal{S}}}(T,g)\right)$, for all $\mu\in\mathcal{M}(M,T)$ and all $\mathcal{M}_{0}^{1}$-piecewise continuous potential $g$,
ii) $h^{*}(T)\leqslant\left(3+2\left\lfloor\frac{\max\tau}{\min\tau}\right\rfloor\right)s_{0}\log 2K$, where $K$ is such that $K_{n}\leqslant Kn$ for all $n\geqslant 1$.
In particular, if the singularities are negligible then i) implies the upper semi-continuity of the pressure map (and of the entropy map in the case $g\equiv 0$). And if the complexity is bounded then ii) implies that $h^{*}(T)=0$.
The above theorem will follows from Corollary 5.3 for i) and from Corollary 6.4 for ii). Concerning the negligible singularities case, i) can be deduced directly. For the bounded complexity case, ii) follows from [15, Proposition 5.6].
Corollary 1.3.
For any finite horizon Sinai billiard table:
i) If there is a maximizing sequence of measures for the entropy, then there exists measures of maximal entropy;
ii) The Kolmogorov–Sinai entropy is upper semi-continuous at every ergodic measure;
iii) If sparse recurrence and negligible recurrence holds, the periodic orbits equidistribute with respect to the unique measure of maximal entropy.
The corollary follows from Corollary 5.5 (in the case $g\equiv 0$) for i), from Corollary 5.6 for ii) and from Corollary 5.8 for iii). Existence of billiard tables satisfying both sparse recurrence and negligible singularities is carried out in Section 7.
2. Negligible Singularities case
In this section we prove the upper semi-continuity of (1.1) under negligible singularities assumption. We begin with an intermediate result. Since $\mathcal{P}$ is a generating partition, we begin by proving the continuity of each $\mu\mapsto\frac{1}{n}H_{\mu}(\mathcal{P}_{l}^{n})$ in Lemma 2.1, where
$H_{\mu}(\mathcal{A})\coloneqq-\sum_{A\in\mathcal{A}}\mu(A)\log\mu(A)$
denotes the static entropy of a partition $\mathcal{A}$. For this, we will use that when the singularities are negligible, $\mu(\partial A)=0$ for all $A\in\mathcal{P}_{l}^{n}$, all $n,l\geqslant 0$ and all $\mu\in\mathcal{M}(M,T)$. We deduce the upper semi-continuity of (1.1) in Proposition 2.2.
Lemma 2.1.
For any finite horizon Sinai billiard with negligible singularities and for any $n,\ell\geqslant 0$, the map $\mu\mapsto\frac{1}{n}H_{\mu}(\mathcal{P}_{-\ell}^{n})$ is continuous on $\mathcal{M}(M,T)$.
Proof.
Fix some $n,\ell\geqslant 0$. We start by proving that for all $A\in\mathcal{P}_{-\ell}^{n}$, $\partial A$ has measure zero. Notice that for all $A\in\mathcal{P}_{-\ell}^{n}$, $\partial A\subset\mathcal{S}_{-\ell-1}\cup\mathcal{S}_{n+1}\subset\mathcal{S}$ (see [3, Lemma 3.3]).
Let $\mu_{k}\rightharpoonup\mu$ in $\mathcal{M}(M,T)$ as $k\to\infty$. Let $A\in\mathcal{P}_{-\ell}^{n}$. By the above, $\mu(\partial A)=0$. Applying Portmanteau theorem, we thus get $\lim_{k\to\infty}\mu_{k}(A)=\mu(A)$. Since $\mathcal{P}_{-\ell}^{n}$ is a finite partition, and by continuity of $x\mapsto-x\log x$, we get the desired result. ∎
We can now prove the upper semi-continuity of (1.1).
Proposition 2.2.
Assuming negligible singularities, then for any $\mathcal{M}_{-\ell}^{n}$-piecewise continuous potential $g:M\to\mathbbm{R}$ the map $\mu\mapsto h_{\mu}(T)+\int g\,\mathrm{d}\mu$ is upper semi-continuous on $\mathcal{M}(M,T)$.
Proof.
First, we prove that $\mu\mapsto h_{\mu}(T)$ is upper semi-continuous. Let $\mu\in\mathcal{M}(M,T)$, $\varepsilon>0$. By the continuity from Lemma 2.1, we get that for all $n\geqslant 1$, there is a neighbourhood $U_{n,\varepsilon}$ of $\mu$ such that for all $\nu\in U_{n,\varepsilon}$ $\frac{1}{n}H_{\mu}(\mathcal{P}_{0}^{n})\geqslant\frac{1}{n}H_{\nu}(\mathcal{P}_{0}^{n})-\varepsilon$. Therefore, since $\mathcal{P}$ is a generating partition,
$h_{\mu}(T)$ $=h_{\mu}(T,\mathcal{P})=\lim_{n\to\infty}\frac{1}{n}H_{\mu}(\mathcal{P}_{0}^{n})\geqslant\lim_{n\to\infty}\sup_{\nu\in U_{n,\varepsilon}}\frac{1}{n}H_{\nu}(\mathcal{P}_{0}^{n})-\varepsilon\geqslant\lim_{n\to\infty}\sup_{\nu\in U_{n,\varepsilon}}h_{\nu}(T,\mathcal{P})-\varepsilon$
$\geqslant\lim_{n\to\infty}\sup_{\nu\in U_{n,\varepsilon}}h_{\nu}(T)-\varepsilon=\limsup_{\nu\rightharpoonup\mu}h_{\nu}(T)-\varepsilon.$
Since this holds for any $\varepsilon>0$, we get the upper semi-continuity of the entropy.
We now prove that $\mu\mapsto\int g\,\mathrm{d}\mu$ is continuous. Let $g$ be a $\mathcal{M}_{-\ell}^{n}$-continuous potential and $\mu\in\mathcal{M}(M,T)$. Since
$\int g\,\mathrm{d}\mu=\sum_{A\in\mathcal{M}_{-\ell}^{n}}\int g\mathbbm{1}_{A}\,{\rm d}\mu,$
we only need to prove the continuity of $\mu\mapsto\int g\mathbbm{1}_{A}\,\mathrm{d}\mu$ for any fixed $A\in\mathcal{M}_{-\ell}^{n}$. Notice that, since $\partial A$ is compact, we have $\partial A=\bigcap_{\delta>0}\mathcal{N}_{\delta}(\partial A)$, where $\mathcal{N}_{\delta}(\partial A)$ is a $\delta$-neighbourhood of $\partial A$. Let $\varepsilon>0$. Then, since $\mu(\partial A)=0$, there exists some $\delta>0$ such that $\mu(\mathcal{N}_{\delta}(\partial A))<\varepsilon$. Up to decreasing the value of $\delta$, we can also assume that $\mu(\partial\mathcal{N}_{\delta}(\partial A))=0$. Let $\mu_{k}\rightharpoonup\mu$ as $k\to\infty$. Define
(2.1) $\varphi_{A,\delta}(x)=\begin{cases}1&\text{if }x\in A\smallsetminus\mathcal{N}_{\delta}(\partial A),\\ \delta^{-1}d(x,\partial A)&\text{if }x\in A\cap\mathcal{N}_{\delta}(\partial A),\\ 0&\text{if }x\notin A,\end{cases}$
and notice that $\varphi_{A}$ is continuous. Thus, we need to estimate the three terms in
(2.2) $\begin{split}\left|\int g\mathbbm{1}_{A}\,{\rm d}\mu_{k}-\int g\mathbbm{1}_{A}\,{\rm d}\mu\right|&\leqslant\left|\int g\mathbbm{1}_{A}\,{\rm d}\mu_{k}-\int g\varphi_{A,\delta}\,{\rm d}\mu_{k}\right|\\ &\,\,+\left|\int g\varphi_{A,\delta}\,{\rm d}\mu_{k}-\int g\varphi_{A,\delta}\,{\rm d}\mu\right|+\left|\int g\varphi_{A,\delta}\,{\rm d}\mu-\int g\mathbbm{1}_{A}\,{\rm d}\mu\right|.\end{split}$
The third term is bounded by $\mu(\mathcal{N}_{\delta}(\partial A))\sup_{A}|g|\leqslant\varepsilon\sup_{A}|g|$. For the first term, we use that since $\mu(\partial\mathcal{N}_{\delta}(\partial A))=0$, there is some $k_{A}$ such that for all $k\geqslant k_{A}$, $\mu_{k}(\mathcal{N}_{\delta}(\partial A))\leqslant\mu(\mathcal{N}_{\delta}(\partial A))+\varepsilon\leqslant 2\varepsilon$. Thus, the first term in (2.2) is bounded by $2\varepsilon\sup_{A}|g|$. Now, for the second term, we use that $g\varphi_{A,\delta}$ is continuous on $M$. Thus, from the weak- convergence, for all $k\geqslant k_{A}$ (up to increasing the value of $k_{A}$), the second term in (2.2) is bounded by $\varepsilon$. This prove the continuity of $\mu\mapsto\int g\mathbbm{1}_{A}\,\mathrm{d}\mu$, and hence the continuity of $\mu\mapsto\int g\,\mathrm{d}\mu$. ∎
From the upper semi-continuity of $\mu\mapsto P_{\mu}(T,g)$ and the compactness of $\mathcal{M}(M,T)$ (see Proposition 3.1), we deduce the existence of equilibrium states. Furthermore, using Proposition 3.2, we can identify the maximal value of the metric pressure.
Proposition 2.3.
For any finite horizon Sinai billiard with negligible singularities and for any piecewise-continuous potential $g:M\to\mathbbm{R}$, the Variational Principle is satisfied, that is $\sup_{\mu\in\mathcal{M}(M,T)}P_{\mu}(T,g)=P_{\rm top}(T,g)$.
The proof follows the same strategies as the one exposed in [23, Theorem 20.2.4]. We write it for completeness. Note that in the process, we construct an equilibrium state.
Proof.
Let $g$ be a piecewise continuous potential. By [11], for any invariant measure $\mu$, $P_{\mu}(T,g)\leqslant P_{\rm top}(T,g)$. We now prove the reverse inequality.
Let $10<\chi_{n}<$ 1 be a sequence converging to $1$. For each $n$, let $E_{n}$ be a finite set containing exactly one point $x_{A}$ of each $A\in\mathcal{P}_{0}^{n}$ such that $e^{S_{n}g(x_{A})}\geqslant\chi_{n}\sup_{A}e^{S_{n}g}$. Let
$\nu_{n}\coloneqq\left(\sum_{x\in E_{n}}e^{S_{n}g(x)}\right)^{-1}\sum_{x\in E_{n}}e^{S_{n}g(x)}\delta_{x},\quad\mu_{n}=\frac{1}{n}\sum_{i=0}^{n-1}T^{i}_{*}\nu_{n}.$
Since $\mathcal{P}_{0}^{n}$ is a partition, each $x\in E_{n}$ belongs to a single $A\in\mathcal{P}_{0}^{n}$, we get
$H_{\nu_{n}}(\mathcal{P}_{0}^{n})+n\int g\,{\rm d}\mu_{n}$ $=H_{\nu_{n}}(\mathcal{P}_{0}^{n})+\int S_{n}g\,{\rm d}\nu_{n}=\sum_{x\in E_{n}}\nu_{n}(\{x\})(S_{n}g(x)-\log\nu_{n}(\{x\}))$
$=\log\sum_{x\in E_{n}}e^{S_{n}g(x)}\geqslant\log\chi_{n}+\log\sum_{A\in\mathcal{P}_{n}}\sup_{A}e^{S_{n}g}$
where the last equality is due to the equality case of [23, Lemma 20.2.2]. Let $q\leqslant n$ and $0\leqslant r<q$. Let $i_{r}$ be the smallest integer such that $n-(r+qi_{r})<q$. Since
$\mathcal{P}_{0}^{n}=\mathcal{P}_{0}^{r}\vee\left(\bigvee_{i=0}^{i_{r}-1}T^{-qi-r}\mathcal{P}_{0}^{q}\right)\vee T^{r+qi_{r}}\mathcal{P}_{0}^{n-(r+qi_{r})},$
we obtain, by sub-additivity,
$H_{\nu_{n}}(\mathcal{P}_{0}^{n})$ $\leqslant H_{\nu_{n}}(\mathcal{P}_{0}^{r})+\sum_{i=0}^{i_{r}-1}H_{\nu_{n}}(T^{-qi-r}\mathcal{P}_{0}^{q})+H_{\nu_{n}}(T^{r+qi_{r}}\mathcal{P}_{0}^{n-(r+qi_{r})})$
$\leqslant 2q\log\#\mathcal{P}+\sum_{i=0}^{i_{r}-1}H_{\nu_{n}}(T^{-qi-r}\mathcal{P}_{0}^{q}).$
Summing over $r$, we get
$qH_{\nu_{n}}(\mathcal{P}_{0}^{n})$ $\leqslant 2q^{2}\log\#\mathcal{P}+\sum_{i=0}^{n}H_{\nu_{n}}(T^{-i}\mathcal{P}_{0}^{q}).$
On the other hand, by concavity of $\phi:t\mapsto-t\log t$, we get
$H_{\mu_{n}}(\mathcal{P}_{0}^{q})$ $=\sum_{A\in\mathcal{P}_{0}^{q}}\phi(\mu_{n}(A))=\sum_{A\in\mathcal{P}_{0}^{q}}\phi\left(\frac{1}{n}\sum_{i=0}^{n-1}T^{i}_{*}\nu_{n}(A)\right)$
$\geqslant\frac{1}{n}\sum_{i=0}^{n-1}\sum_{A\in\mathcal{P}_{0}^{q}}\phi(T^{i}_{*}\nu_{n}(A))=\frac{1}{n}\sum_{i=0}^{n-1}H_{\nu_{n}}(T^{-i}\mathcal{P}_{0}^{q}).$
Putting all of the above estimates together yields
(2.3) $\frac{\log\chi_{n}}{n}+\frac{1}{n}\log\sum_{A\in\mathcal{P}_{0}^{n}}\sup_{A}e^{S_{n}g}\leqslant\frac{2q}{n}\log\#\mathcal{P}+\frac{1}{q}H_{\mu_{n}}(\mathcal{P}_{0}^{q}).$
By Proposition 3.2, there exist $n_{k}\to\infty$ and $\mu_{g}\in\mathcal{M}(M,T)$ such that $\mu_{n_{k}}\rightharpoonup\mu_{g}$. Replacing $n$ by $n_{k}$ in (2.3) and taking the limit, we obtain using Lemma 2.1
$P_{\rm top}(T,g)\leqslant\frac{1}{q}H_{\mu_{g}}(\mathcal{P}_{0}^{q})+\int g\,{\rm d}\mu_{g}.$
Finally, taking the limit $q\to\infty$ yields $P_{\rm top}(T,g)\leqslant P_{\mu_{g}}(T,g)$. ∎
3. $\mathcal{M}(M,T)$ is closed
Since $T$ is not continuous, the transfer operator $T_{*}$ has no reason, a priori, to be continuous on $\mathcal{M}(M)$ with respect to the weak- topology. It is therefore unclear whether the set of fixed points $\mathcal{M}(M,T)$ of $T_{*}$ is a closed set or not, nor if the Krylov–Bogolyubov construction of invariant measures holds. Fortunately, the billiard flow is continuous and since it is a suspension flow over $T$ (because of the finite horizon), invariant measures with respect to $T$ are strongly related to the ones of $\Phi^{t}$. This is the strategy we use.
Proposition 3.1.
For any finite horizon Sinai billiard, the set of $T$-invariant probability measures $\mathcal{M}(M,T)$ is a closed subset of $\mathcal{M}(M)$ for the weak- topology.
Proof.
Let $\mu_{n}\in\mathcal{M}(M,T)$ be a sequence converging to some $\mu_{\infty}\in\mathcal{M}(M)$. Then, in the coordinates of the suspension, $\bar{\mu}_{n}\coloneqq(\int_{M}\tau\,\mathrm{d}\mu_{n})^{-1}\mu_{n}\otimes\lambda$ is a sequence of $\Phi^{t}$-invariant probability measures, where $\lambda$ is the Lebesgue measure in the direction of the flow. Up to taking a subsequence, we can assume that $\bar{\mu}_{n}$ converges to some $\bar{\mu}$. Furthermore, since the billiard flow is continuous, $\bar{\mu}$ is a flow invariant measure.
In particular, there exists a $T$-invariant measure $\mu$ such that $\bar{\mu}=(\int_{M}\tau\,\mathrm{d}\mu)^{-1}\mu\otimes\lambda$. We now prove that $\mu_{\infty}=\mu$.
Consider $\psi\in C^{0}(M)$ and a smooth function $\rho:\mathbbm{R}\to\mathbbm{R}_{+}$ supported in $[-\min\tau/2,\min\tau/2]$ such that $\int_{\mathbbm{R}}\rho\,\mathrm{d}\lambda>0$. Consider a fundamental domain containing $M\times[-\min\tau/2,\min\tau/2]$ of $\Omega\cong(M\times\mathbbm{R})/\sim$ where $\sim$ is such that
$(x,t)\sim(y,u)\Leftrightarrow\exists n\geqslant 0,(y,u)=(T^{n}x,t-\sum_{i=0}^{n-1}\tau(T^{i}x))\text{ or }(x,t)=(T^{n}y,u-\sum_{i=0}^{n-1}\tau(T^{i}y)).$
On this domain, $\varphi(x,t)=\psi(x)\rho(t)$ is well defined and descends into a continuous function on $\Omega$, also called $\varphi$. Thus
(3.1) $\frac{\int_{M}\psi\,{\rm d}\mu_{n}}{\int_{M}\tau\,{\rm d}\mu_{n}}\int_{\mathbbm{R}}\rho\,{\rm d}\lambda=\bar{\mu}_{n}(\varphi)\xrightarrow[n\to\infty]{}\bar{\mu}(\varphi)=\frac{\int_{M}\psi\,{\rm d}\mu}{\int_{M}\tau\,{\rm d}\mu}\int_{\mathbbm{R}}\rho\,{\rm d}\lambda.$
Now, since $\mu_{n}(\varphi)$ converges to $\mu_{\infty}(\varphi)$, and $\int_{\mathbbm{R}}\rho\,\mathrm{d}\lambda>0$, we get from (3.1) that for all $\psi\in C^{0}(M)$, if $\mu_{\infty}(\psi)\neq 0$ then
$\frac{\mu(\tau)}{\mu_{n}(\tau)}\xrightarrow[n\to\infty]{}\frac{\mu(\psi)}{\mu_{\infty}(\psi)},$
and if $\mu_{\infty}(\psi)=0$, then $\mu(\psi)=0$. Therefore, there is a constant $c\geqslant\tau_{\min}/\tau_{\max}>0$ such that for all $\psi\in C^{0}(M)$, $\mu(\psi)=c\,\mu_{\infty}(\psi)$. Since both $\mu$ and $\mu_{\infty}$ are probability measures, it must be that $c=1$, and so $\mu_{\infty}=\mu$ is invariant by $T$. Hence, $\mathcal{M}(M,T)$ is a closed set. ∎
The following proposition will only be used in the proof Proposition 2.3.
Proposition 3.2.
Let $\nu_{n}$ be a sequence in $\mathcal{M}(M)$, and define $\mu_{n}=\frac{1}{n}\sum_{i=0}^{n-1}T_{*}^{i}\nu_{n}$. Then, any accumulation point of the sequence $\mu_{n}$ belongs to $\mathcal{M}(M,T)$.
Proof.
Let $\nu_{n}\in\mathcal{M}(M)$ and define $\mu_{n}$ as in the proposition. Let $n_{k}$ be a strictly increasing sequence of integers and $\mu_{\infty}$ be such that $\mu_{n_{k}}\rightharpoonup\mu_{\infty}$. As in the proof of Proposition 3.1, we define for each $n$, $\bar{\mu}_{n}\coloneqq\left(\int\tau\,\,{\rm d}\mu_{n}\right)^{-1}\mu_{k}\otimes\lambda$. Note that $\int\tau\,\,{\rm d}\mu_{n}=\int S_{n}\tau\,{\rm d}\nu_{n}\geqslant n\tau_{\min}$.
Fix $t\in\mathbbm{R}$. From the suspension coordinates, it follows that for all $\psi\in C^{0}(\Omega,\mathbbm{R})$
$|(\Phi^{t})_{*}\bar{\mu}_{n}(\psi)-\bar{\mu}_{n}(\psi)|\leqslant\frac{2t}{n\tau_{\min}}|\psi|_{C^{0}}$
Therefore, any accumulation point of $\bar{\mu}_{n}$ is $\Phi^{t}$-invariant. Hence, up to extracting a second time, we can assume that $\bar{\mu}_{n_{k}}$ converges to some $\bar{\mu}=(\int\tau\,{\rm d}\mu)^{-1}\mu\otimes\lambda$, where $\mu\in\mathcal{M}(M,T)$ since $\bar{\mu}$ is invariant by the flow. We now prove that $\mu_{\infty}=\mu$.
Let $\psi\in C^{0}(M,\mathbbm{R})$, $\rho\in C^{0}(\mathbbm{R},\mathbbm{R})$ and $\varphi(x,t)=\psi(x)\rho(t)$ as in the proof of Proposition 3.1. Using the exact same argument, we obtain a constant $c\geqslant\tau_{\min}/\tau_{\max}>0$ such that for all $\psi\in C^{0}(M,\mathbbm{R})$, $\mu(\psi)=c\mu_{\infty}(\psi)$. Since $\mu$ and $\mu_{\infty}$ are probability measures, $c=1$ and hence $\mu_{\infty}$ is $T$-invariant. ∎
4. Typicality of Negligible Singularities condition
Assuming the genericness of the bounded complexity assumption, the genericness of the negligible singularities assumption follows from the next proposition.
Proposition 4.1.
If a Sinai billiard table has bounded complexity, then its singularities are negligible.
Proof.
Let $n_{0}\geqslant 1$, $\varphi_{0}<\pi/2$ and $\mu\in\mathcal{M}(M,T)$. Therefore
$\mu(\mathcal{S}_{0})\leqslant\mu(\{|\varphi|>\varphi_{0}\})=\frac{1}{n_{0}}\int\sum_{i=0}^{n_{0}-1}\mathbbm{1}_{\{|\varphi|>\varphi_{0}\}}\circ T^{i}\,{\rm d}\mu\leqslant\frac{1}{n_{0}}\sup_{M}\sum_{i=0}^{n_{0}-1}\mathbbm{1}_{\{|\varphi|>\varphi_{0}\}}\circ T^{i}=s_{0}(n_{0},\varphi_{0}).$
Now, Demers and Korepanov proved in [15, Proposition 5.6] that if the complexity is bounded, then $\inf_{(n_{0},\varphi_{0})}s_{0}(n_{0},\varphi_{0})=0$. Hence $\mu(\mathcal{S}_{0})=0$. By invariance of $\mu$, we get that $\mu(\mathcal{S})=0$. ∎
Since the problem of genericness of the bounded complexity assumption is still (officially) open, we now provide a complete proof of the genericness of the negligible singularities condition. First, identify the set of all Sinai billiard tables with $D$ scatterers and no corner nor cusp by ${\rm Bil}_{D}$ with the families $(f_{i})_{1\leqslant i\leqslant D}\in C^{3+\alpha}(\mathbbm{S}^{1},\mathbbm{T}^{2})$, up to permutation of the indexes, such that for all $1\leqslant i\leqslant D$
i) Each connected component of the lift of $f_{i}(\mathbbm{S}^{1})$ in $\mathbbm{R}^{2}$ is the boundary of a convex set;
ii) $|f^{\prime}_{i}|>0$ is constant, $|f^{\prime\prime}_{i}|\neq 0$, $f_{i}$ turns counter-clockwise and $f_{i}(\mathbbm{S}^{1})$ does not self intersect;
iii) For all $i\neq j$, $f_{i}(\mathbbm{S}^{1})\cap f_{j}(\mathbbm{S}^{1})=\emptyset$.
We will often simply write $Q$ for an element of ${\rm Bil}_{D}$, where $Q=\mathbbm{T}^{2}\smallsetminus\bigsqcup_{i=1}^{D}B_{i}$ and the boundary of each $B_{i}$ is parametrized by $f_{i}$. Now consider ${\rm Bil}^{\rm FH}_{D}$ the subset of ${\rm Bil}_{D}$ made of billiard tables with finite horizon, in the sense that there are no trajectory making only grazing collisions. The sets ${\rm Bil}_{D}$ and ${\rm Bil}^{\rm FH}_{D}$ are thus naturally endowed with the $C^{3}$ topology.
We first prove that ${\rm Bil}^{\rm FH}_{D}$ is an open subset of ${\rm Bil}_{D}$ with respect of the $C^{3+\alpha}$ topology.
Proposition 4.2.
${\rm Bil}^{\rm FH}_{D}$ is an open subset of ${\rm Bil}_{D}$ with respect of the $C^{3+\alpha}$ topology.
In order to prove this proposition, we need the following lemmas (the first of which [^1] The statement of Lemma 4.3 is strongly inspired from a presentation given by Tóth [31]. will be crucial for the rest of this section).
Lemma 4.3.
Let $Q_{0}\in{\rm Bil}_{D}$, $x\in\mathring{Q}_{0}\times\mathbbm{S}^{1}\subset\mathbbm{T}^{2}\times\mathbbm{S}^{1}$, $\varepsilon>0$ and $t>0$. Then, there exist $\delta_{1},\delta_{2}>0$ such that for all $Q_{1}\in{\rm Bil}_{D}\cap B_{C^{3+\alpha}}(Q_{0},\delta_{1})$, and all $y\in(\mathring{Q}_{1}\times\mathbbm{S}^{1})\cap B(x,\delta_{2})$, there exists a continuous, piecewise affine function $h_{y}:[0,t]\xrightarrow{\sim}[0,t]$, $\varepsilon$-close to the identity (in the $C^{0}$ topology), such that
$d\left(\Phi^{s}_{0}(x),\Phi^{h_{y}(s)}_{1}(y)\right)<\varepsilon,\,\forall s\in[0,t],$
where $\Phi_{0}$ and $\Phi_{1}$ are respectively the billiard flows associated to the tables $Q_{0}$ and $Q_{1}$.
Proof.
Let $Q_{0}\in{\rm Bil}_{D}$, $x\in\mathring{Q}_{0}\times\mathbbm{S}^{1}\subset\mathbbm{T}^{2}\times\mathbbm{S}^{1}$, $\varepsilon>0$ and $t>0$. Denote by $\Omega_{0}$ the phase space of $\Phi_{0}$. Since $\Phi_{0}$ is continuous on the compact set $\Omega_{0}\times[0,t]$, it is uniformly continuous. In particular, there is a $\delta_{2}>0$ such that for all $y\in B_{\Omega_{0}}(x,\delta_{2})$, $d_{\Omega_{0}}(\Phi_{0}^{s}(x),\Phi_{0}^{s}(y))<\varepsilon/2$ for all $s\in[0,t]$. Up to decreasing the value of $\delta_{2}$, we can assume that $B_{\Omega_{0}}(x,\delta_{2})\Subset\mathring{Q}_{0}\times\mathbbm{S}^{1}$. Thus, we now only need to control the distance between $\Phi^{s}_{0}(y)$ and $\Phi_{1}^{s}(y)$.
By the group property of the flows, we only need to prove Lemma 4.3 for any $t>0$ small enough. For the rest of the proof, we will assume that $30<t<\min(\tau_{0,\min},\mbox{diam}^{\prime}Q_{0})/$ 3, where $\mbox{diam}^{\prime}Q_{0}$ is the smallest diameter among the diameter of the largest disc contained in each connected component of $\mathbbm{T}^{2}\smallsetminus Q_{0}$. We now provide finitely many upper bounds on $\delta_{1}$ such that, if they are all satisfied, Lemma 4.3 holds.
First, assume that $\delta_{1}<\min(t_{0,\min},d(\pi(x),\partial Q_{0}),\mbox{diam}^{\prime}Q_{0})/3$, where $\pi:\mathbbm{T}^{2}\times\mathbbm{S}^{1}\to\mathbbm{T}^{2}$ is the canonical projection. In that case, for any $Q_{1}\in{\rm Bil}_{D}\cap B_{C^{3+\alpha}}(Q_{0},\delta_{1})$, we get that $\tau_{1,\min}\geqslant\tau_{0,\min}/3>0$, $B(x,\delta_{2})\subset\mathring{Q}_{1}\times\mathbbm{S}^{1}$ and both $\Phi^{[0,t]}_{0}(y)$ and $\Phi^{[0,t]}_{0}(y)$ each contains at most one collision, for any $y\in B(x,\delta_{2})$. We now consider the four different cases whether collisions occur or not for any fixed $y$.
Case 1: Assume that $\Phi^{[0,t]}_{0}(y)$ and $\Phi^{[0,t]}_{0}(y)$ do not contain any collision. Therefore, the two trajectories coincides. The lemma holds in this case if $h_{y}$ is the identity.
Case 2: Assume that $\Phi^{[0,t]}_{0}(y)$ makes one collision and that $\Phi^{[0,t]}_{1}(y)$ does not. In order to simplify notations (also for the subsequent cases) define $\gamma_{0}:\mathbbm{S}^{1}\to\mathbbm{S}^{1}$ by
$\gamma_{0},i(r)\coloneqq\frac{f_{0,i}^{\prime}(r)}{|f_{0,i}^{\prime}(r)|},$
where $(f_{i})=Q_{0}$. Note that $\gamma_{0,i}$ is a $C^{2}$-diffeomorphism. For $\varphi_{0}\in(0,\pi/2)$ to by specified latter, let
$F_{0,i}(r,r_{0})=\left\langle\gamma_{0,i}(r_{0,i}),\gamma_{0,i}(r)\right\rangle-\cos\left(\frac{\pi}{2}-\varphi_{0}\right).$
Note that for each $r\in\mathbbm{S}^{1}$ and $1\leqslant i\leqslant D$, there are exactly two solutions $r_{0,i}^{\pm}(r)$ to $F_{0,i}(r,r_{0,i}^{\pm}(r))=0$. We can assume that the triplets $(r_{0,i}^{-}(r),r,r_{0,i}^{+}(r))$ keep the same orientation in $\mathbbm{S}^{1}$. Since $\varphi_{0}\neq\pi/2$, $\frac{\partial F_{0,i}}{\partial r_{0}}(r,r_{0,i}^{\pm})\neq 0$, by the implicit function theorem, all $r\mapsto r_{0,i}^{\pm}(r)$ are $C^{2}$ functions. Let
$G_{0,i}^{\pm}(r)=\left\langle f_{0,i}(r_{0,i}^{\pm}(r))-f_{0,i}(r),\frac{f^{\prime\prime}_{0,i}(r)}{|f^{\prime\prime}_{0,i}(r)|}\right\rangle.$
By continuity and compactness, $\delta_{Q_{0}}\coloneqq\inf\{G_{0,i}^{*}(r)\mid r\in\mathbbm{S}^{1},*=-,+,1\leqslant i\leqslant D\}>0$. Therefore, up to decreasing the value of $\delta_{1}$, we can assume that $\delta_{1}<\delta_{Q_{0}}$. Thus, for all $Q_{1}\in{\rm Bil}_{D}\cap B_{C^{3+\alpha}}(Q_{0},\delta_{1})$, the collision in $\Phi^{[0,t]}_{0}(y)$ is $\varphi_{0}$-grazing. Now, the distance between $\pi(\Phi_{0}^{s}(y))$ and $\pi(\Phi_{1}^{s}(y))$ is at most $2t\sin(\pi/2-\varphi_{0})$. Thus, taking
$\pi/2-\varphi_{0}<\min\left(\frac{\pi}{2}\frac{\varepsilon}{8t},\frac{\varepsilon}{8}\right),$
the distances in $\mathbbm{T}^{2}$ and $\mathbbm{S}^{1}$ are both at most $\varepsilon/4$. Hence, the lemma holds if $h_{y}$ is the identity.
Case 3: Assume that $\Phi^{[0,t]}_{1}(y)$ makes one collision and that $\Phi^{[0,t]}_{0}(y)$ does not. Similarly as in Case 2, we define $\gamma_{j,i}(r,\varphi_{0})$, $F_{1,i}(r,r_{j},\varphi_{0})$, $G^{\pm}_{1,i}(r,\varphi_{0})$ and $\delta_{Q_{1},\varphi_{0}}$, $j=0,1$, for any $Q_{1}\in{\rm Bil}_{D}\cap B_{C^{3+\alpha}}(Q_{0},\delta_{1})$, where we now explicit the dependence on $\varphi_{0}$. Our goal is to find $\delta_{1}$ and $\varphi_{0}$ such that $Q_{1}\mapsto\delta_{Q_{1},\varphi_{0}}$ is bounded from below by $\delta_{Q_{0},\varphi_{0}}/2$ on a $\delta_{1}$-neighbourhood of $Q_{0}$.
We start by giving an expansion of the distance between $r^{\pm}_{j,i}(r)$ and $r$ as $\varphi\to\pi/2$. First, notice that this distance can be bounded by above by the analogous quantity in the case where the curvature of the scatterers is constant (and equals the minimum). It follows that
$d_{\mathbbm{S}^{1}}(r^{\pm}_{j,i}(r),r)\leqslant\sup_{r,i}\frac{|f^{\prime}_{j,i}|}{|f^{\prime\prime}_{j,i}(r)|}\left(\frac{\pi}{2}-\varphi_{0}\right).$
Fix some $\varphi_{0}$ to be chosen latter. Then, using that $F^{\pm}_{j,i}(r,r^{\pm}_{j,i}(r,\varphi_{0}),\varphi_{0})=0$ and expanding $\gamma_{j,i}(r^{\pm}_{j,i}(r))$ around $r$, we get
$\frac{1}{2}(r^{\pm}_{j,i}(r)-r)^{2}\left(\left\langle\gamma^{\prime\prime}_{j,i}(r),\gamma_{j,i}(r)\right\rangle+\left\langle\gamma^{\prime\prime}_{j,i}(\hat{r}^{\pm}_{j,i}(r))-\gamma^{\prime\prime}_{j,i}(r),\gamma_{j,i}(r)\right\rangle\right)=\cos\left(\frac{\pi}{2}-\varphi\right)-1,$
where $\hat{r}^{\pm}_{j,i}(r)\in(r,r^{\pm}_{j,i}(r))$. Note that $\left\langle\gamma^{\prime\prime}_{j,i}(r),\gamma_{j,i}(r)\right\rangle=-\frac{|f^{\prime\prime}_{j,i}(r)|}{|f^{\prime}_{j,i}|}$ and that
$\left|\left\langle\gamma^{\prime\prime}_{j,i}(\hat{r}^{\pm}_{j,i}(r))-\gamma^{\prime\prime}_{j,i}(r),\gamma_{j,i}(r)\right\rangle\right|\leqslant\frac{{\rm H\ddot{o}l}_{\alpha}f^{\prime\prime\prime}_{j,i}}{|f^{\prime}_{j,i}|}d_{\mathbbm{S}^{1}}(r^{\pm}_{j,i}(r),r)\leqslant\frac{{\rm H\ddot{o}l}_{\alpha}f^{\prime\prime\prime}_{j,i}}{\inf_{r,i}|f^{\prime\prime}_{j,i}(r)|}\left(\frac{\pi}{2}-\varphi_{0}\right).$
Thus,
$\left|\frac{1}{2}(r^{\pm}_{j,i}(r)-r)^{2}-\frac{1}{2}\frac{|f^{\prime}_{j,i}|}{|f^{\prime\prime}_{j,i}(r)|}\left(\frac{\pi}{2}-\varphi_{0}\right)^{2}\right|\leqslant C_{j,i}\,\left(\frac{\pi}{2}-\varphi_{0}\right)^{2+\alpha}+C_{j}\left(\frac{\pi}{2}-\varphi_{0}\right)^{4},$
where
$C_{j,i}=\frac{1}{2}\frac{{\rm H\ddot{o}l}_{\alpha}f^{\prime\prime\prime}_{j,i}}{\inf_{r,i}|f^{\prime\prime}_{j,i}(r)|}\quad\text{ and }\quad C_{j}=\frac{2}{4!}\sup_{r,i}\frac{|f^{\prime}_{j,i}|}{|f^{\prime\prime}_{j,i}(r)|}$
Now, expanding $f_{j,i}(r^{\pm}_{j,i}(r))$ at $r$, we get that
$\left|G^{\pm}_{j,i}(r,r^{\pm}_{j,i}(r),\varphi_{0})-\frac{1}{2}\frac{|f^{\prime}_{j,i}|}{|f^{\prime\prime}_{j,i}(r)|}\left(\frac{\pi}{2}-\varphi_{0}\right)^{2}\right|$
$\qquad\leqslant C_{j,i}\left(\frac{\pi}{2}-\varphi_{0}\right)^{2+\alpha}+\frac{1}{6}|f^{\prime\prime\prime}_{j,i}|_{C^{0}}\left(\frac{\pi}{2}-\varphi_{0}\right)^{3}+C\left(\frac{\pi}{2}-\varphi_{0}\right)^{4}$
as well as
$\left|G^{\pm}_{1,i}(r,\varphi_{0})-G^{\pm}_{0,i}(r,\varphi_{0})\right|$
$\qquad\leqslant\frac{1}{2}\left|\frac{|f^{\prime}_{0}|}{|f^{\prime\prime}_{0}(r)|}-\frac{|f^{\prime}_{1}|}{|f^{\prime\prime}_{1}(r)|}\right|\left(\frac{\pi}{2}-\varphi_{0}\right)^{2}+(C_{1,i}+C_{0,i})\left(\frac{\pi}{2}-\varphi_{0}\right)^{2+\alpha}+(C_{1}+C_{0})\left(\frac{\pi}{2}-\varphi_{0}\right)^{4}$
$\qquad\leqslant\bar{\varepsilon}\left(\frac{\pi}{2}-\varphi_{0}\right)^{2},$
where for the last inequality, we used that for all $\bar{\varepsilon}>0$ there exists $\bar{\delta}>0$ such that for all $Q_{1}\in B_{C^{3+\alpha}}(Q_{0},\bar{\delta})\cap{\rm Bil}_{D}$,
$\left|\frac{|f^{\prime}_{0}|}{|f^{\prime\prime}_{0}(r)|}-\frac{|f^{\prime}_{1}|}{|f^{\prime\prime}_{1}(r)|}\right|$ $<\frac{\bar{\varepsilon}}{4},$
$C_{1,i}$ $\leqslant 2C_{0,i},$
$C_{1}$ $\leqslant 2C_{0},$
and for $\varphi_{0}$ close enough to $\pi/2$ so that
$3C_{0,i}\left(\frac{\pi}{2}-\varphi_{0}\right)^{\alpha}$ $<\frac{\bar{\varepsilon}}{4},$
$3C_{0}\left(\frac{\pi}{2}-\varphi_{0}\right)^{2}$ $<\frac{\bar{\varepsilon}}{4}.$
[…]
$G^{*}_{1,i}(r,\varphi_{0})$ $\geqslant G^{*}_{0,i}(r,\varphi_{0})-\bar{\varepsilon}\left(\frac{\pi}{2}-\varphi_{0}\right)^{2}$
$\geqslant\bar{\varepsilon}\left(\frac{\pi}{2}-\varphi_{0}\right)^{2}>0.$
Up to decreasing the value of $\delta_{1}$, we can assume that $\delta_{1}<\bar{\delta}$ and $\delta_{1}<\bar{\varepsilon}\left(\frac{\pi}{2}-\varphi_{0}\right)^{2}$. In that case, the collision of $\Phi^{[0,t]}_{1}(y)$ must be $\varphi_{0}$-grazing. We can now conclude as in Case 2.
Case 4: Assume that $\Phi^{[0,t]}_{0}(y)$ and $\Phi^{[0,t]}_{1}(y)$ both make one collision. By the choice of $t$, the collisions must occur on scatterers with common indexes, say $i_{0}$. Denote $t_{0}\in(0,t)$ (resp.$t_{1}$) the time at which $\Phi^{[0,t]}_{0}(y)$ (resp. $\Phi^{[0,t]}_{1}(y)$) makes a collision. Denote also $r_{0}$ (resp. $r_{1}$) such that the collision occurs at $f_{0,i_{0}}(r_{0})$ (resp. $f_{1,i_{0}}(r_{1})$).
Since $f^{\prime\prime}_{0,i}(r)\neq 0$ for all $r\in\mathbbm{S}^{1}$ and $1\leqslant i\leqslant D$, for all $\bar{\varepsilon}>0$, there exists $\bar{\delta}>0$ such that $\mathcal{N}_{\delta}(\partial Q_{0})$ does not contain any segment of length more than $\bar{\varepsilon}$. Up to decreasing the value of $\bar{\delta}$, we can assume that $\bar{\delta}\leqslant\bar{\varepsilon}$.
For $\bar{\varepsilon}$ to be chosen latter, assume that $\delta_{1}<\bar{\delta}$. It follows that $|t_{0}-t_{1}|<\bar{\varepsilon}$, and
$d(f_{0,i_{0}}(r_{0}),f_{0,i_{0}}(r_{1}))\leqslant d(f_{0,i_{0}}(r_{0}),f_{1,i_{0}}(r_{1}))+d(f_{1,i_{0}}(r_{1}),f_{0,i_{0}}(r_{1}))\leqslant\bar{\varepsilon}+\bar{\delta}\leqslant 2\bar{\varepsilon}$
Computing the length of the arc in $\partial Q_{0}$ joining $f_{0,i_{0}}(r_{0})$ to $f_{0,i_{0}}(r_{1})$, it comes that
$d_{\mathbbm{S}^{1}}(r_{0},r_{1})\leqslant\frac{2\bar{\varepsilon}}{|f^{\prime}_{0,i_{0}}|}.$
Now, in order to estimate the distance between the outgoing trajectory angles, we compute
$\left|\gamma_{1,i_{0}}(r_{1})-\gamma_{0,i_{0}}(r_{0})\right|$ $\leqslant\frac{1}{|f^{\prime}_{0,i_{0}}(r_{0})|}\left(2{\rm Lip}(f^{\prime})d_{\mathbbm{S}^{1}}(r_{0},r_{1})+2d_{C^{3+\alpha}}(Q_{0},Q_{1})\right)$
$\leqslant 2\left(\frac{|f^{\prime\prime}_{0,i_{0}}|_{C^{0}}}{|f^{\prime}_{0,i_{0}}|_{C^{0}}^{2}}+\frac{2}{|f^{\prime}_{0,i_{0}}|_{C^{0}}}\right)\bar{\varepsilon}.$
Thus, fixing $\bar{\varepsilon}$ small enough, we can assume that the angle between the two outgoing trajectories is $2d_{\mathbbm{S}^{1}}(\gamma_{1,i_{0}}(r_{1}),\gamma_{0,i_{0}}(r_{0}))<\frac{\varepsilon}{4t}$, as well as $\bar{\varepsilon}\leqslant\varepsilon/4$. Finally, defining $h_{y}:[0,t]\to[0,t]$ by
$h_{y}(s)=\begin{cases}\frac{t_{0}}{t_{1}}s,&\text{if }0\leqslant s\leqslant t_{1},\\ \frac{t-t_{0}}{t-t_{1}}s,&\text{if }t_{1}\leqslant s\leqslant t,\end{cases}$
the collisions of $s\mapsto\Phi^{s}_{0}(y)$ and $s\mapsto\Phi^{h_{y}(s)}_{1}(y)$ both occur at $s=t_{0}$. Assuming that $\delta_{1}<\bar{\varepsilon}$ and $d_{C^{3+\alpha}}(Q_{0},Q_{1})<\delta_{1}$, we have for $s\leqslant t_{0}$,
$d(\Phi^{s}_{0}(y),\Phi_{1}^{h_{y}(s)}(y))\leqslant\bar{\varepsilon}<\varepsilon/2,$
and for $s\geqslant t_{0}$,
$d(\Phi^{s}_{0}(y),\Phi_{1}^{h_{y}(s)}(y))\leqslant\bar{\varepsilon}+2sd_{\mathbbm{S}^{1}}(\gamma_{1,i_{0}}(r_{1}),\gamma_{0,i_{0}}(r_{0}))\leqslant\varepsilon/2,$
finishing the proof of the lemma. ∎
For each $Q\in{\rm Bil}_{D}$ define $s_{0}(Q)=\inf_{(n_{0},\varphi_{0})}s_{0}(n_{0},\varphi_{0})$. The next results characterise the finiteness of the horizon in terms of the value of $s_{0}$.
Lemma 4.4.
For all $Q\in{\rm Bil}_{D}$, $Q\in{\rm Bil}^{\rm FH}_{D}$ if and only if $s_{0}(Q)<1$. Furthermore, for all $Q\in{\rm Bil}^{\rm FH}_{D}$, there exist $\varepsilon>0$ and $\delta>0$ such that $s_{0}\leqslant 1-\varepsilon$ on $B_{C^{3+\alpha}}(Q,\delta)\cap{\rm Bil}_{D}$.
Proof.
We begin with the characterisation of ${\rm Bil}^{\rm FH}_{D}$ in terms of $s_{0}$.
Let $Q\in{\rm Bil}^{\rm FH}_{D}$. By contradiction assume that $s_{0}(Q)=1$. Let $20<\varphi_{k}<\pi/$ 2 be a sequence converging to $\pi/2$ as $k\to\infty$. Therefore, for all $n$ and $k$, there exists $x_{n,k}\in M$ maximizing $s_{0}(n,\varphi_{k})=1$, that is, $x_{n,k}$ makes $n$ consecutive $\varphi_{k}$-grazing collisions. By compactness, and up to a diagonal extraction, we can assume that for all $n$, $x_{n,k}$ converges to some $x_{n}\in M$ as $k\to+\infty$. Note that the flow orbit of $x_{n}$ must be a straight line when restricted to times in $(0,n\tau_{\min})$. In particular $x_{n}$ makes at least $n\frac{\tau_{\min}}{\tau_{\max}}$ grazing collisions. By compactness, we can assume that $x_{n}$ converges to some $x\in M$ as $n\to\infty$. Therefore, for all $t>0$, $\Phi^{[0,T]}(x)$ is arbitrarily close to a straight line. Hence, it is a straight line. In particular, $x$ makes only grazing collisions, contradicting the finite horizon assumption on $Q$.
Conversely, if $Q\in{\rm Bil}_{D}\smallsetminus{\rm Bil}^{\rm FH}_{D}$, then there is a point $x\in M$ making only grazing collisions. In particular $s_{0}(n,\varphi)=1$ for all $n$ and $\varphi$. Hence $s_{0}(Q)=1$.
Let $Q_{0}\in{\rm Bil}^{\rm FH}_{D}$. By contradiction, assume that for all $\varepsilon>0$ and $k\geqslant 1$, there exists $Q_{k}^{\varepsilon}\in B_{C^{3+\alpha}}(Q_{0},1/k)\cap{\rm Bil}_{D}$ such that $s_{0}(Q_{k}^{\varepsilon})>1-\varepsilon$. Remark that for all $n_{0},\varphi_{0}$, $s_{0}(Q_{k}^{\varepsilon},n_{0},\varphi_{0})>1-\varepsilon$, therefore there exists a point $x\in M_{k,\varepsilon}$ maximizing $s_{0}(Q_{k}^{\varepsilon},n_{0},\varphi_{0})$, thus making at least $(1-\varepsilon)n_{0}$ grazing collisions among $n_{0}$ consecutive collisions. In particular, it follows that there is a point $x^{\prime}$ (belonging to $\{x,T_{k}x,\ldots T_{k}^{n_{0}}x\}$) making at least $\frac{n_{0}s_{0}(Q_{k},n_{0},\varphi_{0})}{n_{0}(1-s_{0}(Q_{k},n_{0},\varphi_{0}))+1}\geqslant\frac{1-\varepsilon}{2\varepsilon}$ consecutive grazing collisions, as long as $1/n_{0}<\varepsilon$.
Fix $t>0$ and chose some $20<\varphi<\pi/$ 2 such that
$\left|\frac{\pi}{2}-\varphi\right|<\varepsilon\left(\frac{n(n+1)}{2}\frac{\pi}{2}t\right)^{-1}.$
Therefore, any flow orbit of length at most $t$ making at most $n$ collisions, all of which being $\varphi$-grazing, deviates from a straight line by at most $\varepsilon$, when seen in $\mathbbm{T}^{2}$.
For each $k$ and $\varepsilon$, let $x_{k,\varepsilon}\in M_{k,\varepsilon}$ be a maximizer of $s_{0}(Q_{k}^{\varepsilon},n,\varphi)$. As shown above, there is a point $x^{\prime}_{k,\varepsilon}$ making at least $\frac{1-\varepsilon}{2\varepsilon}$ successive $\varphi$-grazing collisions. These points may get arbitrary close to $\partial Q_{0}$, preventing from using Lemma 4.3. For $k$ large enough so that $31/k<\tau_{\min}/$ 3, there exists $x^{\prime\prime}_{k,\varepsilon}=\Phi^{-t_{k,\varepsilon}}_{k,\varepsilon}(x^{\prime}_{k,\varepsilon})$ such that there are no collision between $x^{\prime\prime}_{k,\varepsilon}$ and $x^{\prime}_{k,\varepsilon}$, and $d(x^{\prime\prime}_{k,\varepsilon},\partial Q_{0})\geqslant\tau_{\min}/3$. Note that $t_{k,\varepsilon}>0$ can be chosen to be smaller than
$t^{\prime}\coloneqq\sqrt{\frac{\tau_{\min}}{3}\left(\frac{\tau_{\min}}{3}+\frac{2}{\inf_{r,i}|f_{0,i}^{\prime\prime}(r)|}\right)},$
whenever $21/k<\inf_{r,i}|f_{0,i}^{\prime\prime}(r)|/$ 2. By compactness, we can assume that $x^{\prime\prime}_{k,\varepsilon}$ converges to some $x^{\prime\prime}_{\varepsilon}$ as $k\to\infty$. Since, $d(x^{\prime\prime}_{\varepsilon},\partial Q_{0})>0$, we can apply Lemma 4.3 to $Q_{0}$, $x^{\prime\prime}_{\varepsilon}$, $\varepsilon$ and $t$, and obtain some positive $\delta_{1}$ and $\delta_{2}$. Denote $t_{\varepsilon}=\min\left(t,\frac{\tau_{\min}}{3}\frac{1-\varepsilon}{2\varepsilon}\right)$. Therefore, when $k$ is large enough so that $1/k<\delta_{1}$ and $d(x^{\prime\prime}_{k,\varepsilon},x^{\prime\prime}_{\varepsilon})<\delta_{2}$, $\pi(\Phi_{0}^{[0,t_{\varepsilon}]}(x^{\prime\prime}_{\varepsilon}))$ deviates from a straight line by at most $2\varepsilon$.
By compactness, let $\varepsilon_{n}\to 0$ be such that $x^{\prime\prime}_{\varepsilon_{n}}$ converges to some $x^{\prime\prime}$. By uniform continuity of $\Phi_{0}$ on $[0,t]\times\Omega_{0}$, $\pi(\Phi_{0}^{[0,t]})$ coincides with a straight line, that is, $x^{\prime\prime}$ makes only grazing collisions. In particular, the orbit of $x^{\prime\prime}$ makes at least $t/\tau_{\max}$ consecutive grazing collisions. Since this construction holds for all $t>0$, we get that $s_{0}(Q_{0})=1$, a contradiction. ∎
Proof of Proposition 4.2.
Let $Q_{0}\in{\rm Bil}^{\rm FH}_{D}$. Denote by $\tilde{\kappa}_{\min}$ the minimal value of $|f^{\prime\prime}_{i}|$, $1\leqslant i\leqslant D$, and by $V>0$ the minimal value of $|f^{\prime}_{i}|$, $1\leqslant i\leqslant D$. By compactness, $\tilde{\kappa}_{\min}>0$. Furthermore, by definition of ${\rm Bil}^{\rm FH}_{D}$, $\tau_{\min}>0$. Let $\delta_{1}=\min\{\tilde{\kappa}_{\min},V,\tau_{\min}\}/3$. In particular, for all $Q\in B_{C^{3+\alpha}}(Q_{0},\delta_{1})\cap{\rm Bil}_{D}$, the scatterers of $Q$ are pairwise disjoint, do not self intersect and are strictly convex sets (not reduced to a point). Now, let $\varepsilon>0$ and $\delta>0$ as provided by Lemma 4.4. Thus, up to decreasing the value of $\delta$ so that $\delta<\delta_{1}$,
$B_{C^{3}}(Q_{0},\delta)\cap{\rm Bil}_{D}\subset{\rm Bil}^{\rm FH}_{D},$
finishing the proof. ∎
Now, given a table $Q\in{\rm Bil}^{\rm FH}_{D}$, we characterize the measures $\mu\in\mathcal{M}(M,T)$ such that $\mu(\mathcal{S})\neq 0$. Since $\mathcal{S}$ is invariant, the conditional measures $\mu_{\mathcal{S}}$ and $\mu_{\mathcal{S}^{c}}$ are also invariant, and $\mu_{\mathcal{S}}(S)=1$. Thus, we only need to characterize measures $\mu$ such that $\mu(\mathcal{S})=1$. Denote $\mathcal{M}_{\mathcal{S}}(M,T)$ the set of those measures.
Lemma 4.5.
If $\mu\in\mathcal{M}_{\mathcal{S}}(M,T)$, then $\mu$ is supported by an at most countable family of periodic orbits, each one making at least one grazing collision.
Proof.
Let $\mu\in\mathcal{M}_{\mathcal{S}}(M,T)$. Then, by invariance, $\mu(\mathcal{S})=1$ implies $\mu(\mathcal{S}_{0})>0$. Let $(x_{k})_{1\leqslant k\leqslant k_{n}}$ be the points in $\mathcal{S}_{n}$ where multiple curves intersect. Let
[…]
Therefore $X_{n}$ is finite and $W_{n}\coloneqq\mathcal{S}_{0}\smallsetminus X_{n}$ is such that the sets $(T^{-i}W_{n})_{0\leqslant i\leqslant n}$ are pairwise disjoint. Let $W\coloneqq\bigcap_{n\geqslant 1}W_{n}$. Therefore, all the $T^{-n}W$, $n\geqslant 0$ are pairwise disjoint. In particular $\mu(W)=0$. Finally, note that $\mathcal{S}_{0}\smallsetminus W$ is at most countable.
Therefore, $\mu_{\mathcal{S}_{0}}$ is an atomic measure on $\mathcal{S}_{0}$. Since $\mu$ is a probability measure, by invariance these atoms must be supported by periodic points. ∎
From Lemma 4.5, it follows immediately that a billiard table has negligible singularities if and only if there are no periodic orbit making grazing collisions. We thus introduce the sets, for any $k\geqslant 3$,
$\mathcal{U}_{k}\coloneqq\{Q\in{\rm Bil}^{\rm FH}_{D}\mid\bigcup_{i=3}^{k}{\rm Fix}\,\,T^{i}\cap\mathcal{S}_{0}=\emptyset\}.$
Therefore, $\bigcap_{k\geqslant 3}\mathcal{U}_{k}$ is the set of billiard tables with negligible singularities. The typicality of this condition then follows from the next proposition.
Proposition 4.6.
For all $k\geqslant 3$, $\mathcal{U}_{k}$ is open and dense in ${\rm Bil}^{\rm FH}_{D}$ with respect to the $C^{3+\alpha}$ topology. In particular, the set of finite horizon Sinai billiard tables with negligible singularities is a dense $G_{\delta}$ set.
Note that billiard tables admitting grazing periodic orbits are suspected to be dense. A partial result in that direction can be found in [27].
Proof.
Let $k\geqslant 3$. We first prove that $\mathcal{U}_{k}$ is open. Let $Q_{0}\in\mathcal{U}_{k}$. By contradiction, assume that for all $n\geqslant 1$, there exists $Q_{n}\in B_{C^{3+\alpha}}(Q_{0},1/n)$ such that $Q_{n}\notin\mathcal{U}_{k}$. Denote $\Phi_{n}$ the billiard flow associated with $Q_{n}$, and $T_{n}:M_{n}\to M_{n}$ its collision map. Therefore, there is a periodic point $x_{n}\in M_{n}\subset\mathbbm{T}^{2}\times\mathbbm{S}^{1}$ of period at most $k$ with respect to $T_{n}$ making at least one grazing collision. For $n$ large enough so that $d_{C^{3+\alpha}}(Q_{0},Q_{n})<\tau_{\min}/3$, denote by $x^{\prime}_{n}$ a point along the flow orbit of $x_{n}$ such that $d(\pi(x^{\prime}_{n}),\partial Q_{0})\geqslant\tau_{\min}/3$. Let $0<p_{n}<k\tau_{\max}$ be the smallest period of $x^{\prime}_{n}$ under $\Phi_{n}$, and $t_{n}$ be such that $\Phi_{n}^{t_{n}}(x^{\prime}_{n})$ is a grazing collision. By compactness, up to extracting, we can assume that $x^{\prime}_{n}$ converges to some $x^{\prime}\in\mathbbm{T}^{2}\times\mathbbm{S}^{1}$ such that $d(\pi(x^{\prime}),\partial Q_{0})>0$, and, up to extracting further, that $p_{n}$ and $t_{n}$ converge to $p,t^{\prime}\in[0,k\tau_{\max}]$ respectively. Now, applying Lemma 4.3 for $Q_{0}$, $x^{\prime}$, $\varepsilon$ and $t=k\tau_{\max}$ we obtain some $\delta_{1},\delta_{2}>0$. For any $n$ large enough, $d_{C^{3+\alpha}}(Q_{n},Q_{0})<\delta_{1}$ and $d(x_{n},x)<\delta_{2}$. Therefore
$d(x^{\prime},\Phi_{0}^{t_{n}}(x^{\prime}))\leqslant d(x,x_{n})+d(\Phi_{n}^{t_{n}}(x_{n}),\Phi_{0}^{t_{n}}(x))\leqslant 2\varepsilon$
Thus $x$ is a periodic orbit of period $p$ for $\Phi_{0}$ and belongs to $M_{0}$. Now, since
$d(\Phi_{0}^{t_{n}}(x^{\prime}),M_{0})$ $\leqslant d(\Phi_{0}^{t_{n}}(x^{\prime}),\Phi_{n}^{t_{n}}(x^{\prime}_{n}))+d(\Phi_{n}^{t_{n}}(x^{\prime}_{n}),M_{n})+d(M_{n},M_{0})$
$\leqslant\varepsilon+1/n\leqslant 2\varepsilon,$
and by continuity of $\Phi_{0}$, $\Phi_{0}^{t^{\prime}}(x^{\prime})$ is a collision point. This is a grazing collision since the flow trajectory of $x^{\prime}$ for times near $t^{\prime}$ is arbitrarily close to segments of length at least $\tau_{\min}/3$. This contradicts $Q_{0}\in\mathcal{U}_{k}$, and thus there exists $n$ such that $B_{C^{3+\alpha}}(Q_{0},1/n)\subset\mathcal{U}_{k}$.
We now prove the density of $\mathcal{U}_{k}$ in ${\rm Bil}^{\rm FH}_{D}$. Let $Q_{0}\in{\rm Bil}^{\rm FH}_{D}$. Let $\delta$ be so small that $Q\mapsto\#\mathcal{P}_{Q}$ is uniformly bounded from above by some finite $C\geqslant 2$. Denote $\delta^{\prime}=\delta/2C^{k+1}$. Given a sequence $Q_{n}\in{\rm Bil}^{\rm FH}_{D}$, denote
$P_{n}$ $=\bigcup_{i=3}^{k}{\rm Fix}\,\,T_{n}^{i}\subset M_{n}\subset\mathbbm{T}^{2}\times\mathbbm{S}^{1},$
$R_{n}$ =$=\{x\in M_{n}\mid x\in P_{n}\text{ such that }\mathcal{O}_{n}(x)\cap\partial M_{n}=\emptyset\},$
where $\mathcal{O}_{n}(x)$ denotes the orbit of $x$ under $T_{n}$, the collision map associated with $Q_{n}$. We now proceed algorithmically by constructing a finite sequence of tables $Q_{n}$ and sets $R^{\prime}_{n}\subset R_{n}$ with $d_{C^{3+\alpha}}(Q_{n},Q_{n+1})<\delta^{\prime}$ and #$\#R^{\prime}_{n}\geqslant n$ until $R^{\prime}_{n}=P_{n}$. These constructions have to stop at some point since $d_{C^{3+\alpha}}(Q_{n},Q_{0})<\delta$ for all $n\leqslant 2C^{k+1}$, and thus $n\leqslant\#R^{\prime}_{n}\leqslant\#P_{n}\leqslant C^{k+1}$.
Initiate $R^{\prime}_{0}$ by setting $R^{\prime}_{0}=R_{0}$. Assume $Q_{n}$ and $R^{\prime}_{n}$ have been constructed. If $R^{\prime}_{n}=P_{n}$, then $Q_{n}\in\mathcal{U}_{k}$ and we stop the construction here. Else, there is a point $x\in P_{n}\smallsetminus R^{\prime}_{n}$. Since $R^{\prime}_{n}\subset R_{n}$, the point $x_{n}$ is periodic of period at most $k$ with respect to $T_{n}$ and makes at least one grazing collision. Without loss of generality, we assume that $x$ is a regular collision. Denote $\pi:\mathbbm{T}^{2}\times\mathbbm{S}^{1}\to\mathbbm{T}^{2}$ the canonical projection. Write $\mathcal{O}_{n}(x_{n})\supset\{y_{1},\ldots,y_{i_{1}},z_{1},\ldots,z_{i_{2}}\}$ the points of the orbit where the collisions are grazing, where for all $i$, $\pi(y_{i})\in\pi(R^{\prime}_{n})$ and $\pi(z_{i})\notin\pi(R^{\prime}_{n})$. Let $U_{1}$ be a neighbourhood of the $\pi(y_{i})$’s in $\partial Q_{n}\subset\mathbbm{T}^{2}$ such that $\bar{U}_{1}\cap\left(\pi(\{z_{1},\ldots z_{i_{2}}\}\cup R^{\prime}_{n})\smallsetminus\pi(\{y_{1},\ldots,y_{i_{2}}\})\right)=\emptyset$. Let $U_{2}$ be a neighbourhood of the $\pi(z_{i})$’s in $\partial Q_{n}$ such that $\bar{U}_{2}\cap\pi(R^{\prime}_{n})=\bar{U}_{1}\cap\bar{U}_{2}=\emptyset$. We construct $Q_{n+1}$ by perturbing $\partial Q_{n}$ only on $U_{1}\cup U_{2}$. On $U_{2}$ we shrink the scatterers so that the $\pi(z_{i})$’s are not part of $\partial Q_{n+1}$ and such that $d_{C^{3+\alpha}}(Q_{n+1},Q_{n})<\delta^{\prime}$. Note that the flow trajectory of $x$ is unchanged. The perturbation on $U_{1}$ needs more care, we proceed as follows.
Let $y$ be a point of $R^{\prime}_{n}\smallsetminus\{y_{1},\ldots,y_{i_{1}}\}$ such that $\pi(y)\in\pi(R^{\prime}_{n})$. Since the orbit of $y^{\prime}$ makes only regular collisions, by the hyperbolic fixed point theorem, there is a $\delta_{y}>0$ such that for any $Q\in B_{C^{3+\alpha}}(Q_{n},\delta_{y})$, there corresponds a unique periodic point $y^{\prime}$ whose orbit contains only regular collisions, with the same period as $y$ (with respect to the collision maps), and whose orbits with respect to the flows are close together. There are finitely many such points $y$. Let $\delta^{\prime}$ be the smallest $\delta_{y}$. Up to decreasing the value of $\delta_{1}$, we can assume $\delta_{1}<\delta^{\prime}$. Finally, pick $Q_{n+1}\in B_{C^{3+\alpha}}(Q_{n},\delta_{1})$ such that $\partial Q_{n+1}$ coincides with $\partial Q_{n}$ outside of $U_{1}\cup U_{2}$ and such that the orbit of $x$ in $Q_{n+1}$ does not make grazing collision. Set $R^{\prime}_{n+1}$ to be the union of the orbit of $x$ and $R^{\prime}_{n}$, where the orbits of each $y$ are replaced by the orbits of $y^{\prime}$. In particular #$\#R^{\prime}_{n+1}=\#R^{\prime}_{n}+\#\mathcal{O}_{n+1}(x)\geqslant\#R^{\prime}_{n}+1$. ∎
5. Regularity of the metric pressure and consequences
In this section we consider general finite horizon Sinai billiards (which may not have negligible singularities). We prove the bound on the defect of upper semi-continuity of the pressure map in Corollary 5.3 (generalizing Corollary 2.2) as well as some consequences, such that existence of equilibrium states as soon as there are maximizing sequences (Corollary 5.5), upper semi-continuity of the entropy map at ergodic measures (Corollary 5.6) and equidistribution of the periodic orbits (Corollary 5.8).
5.1. Bound on the defect of upper semi-continuity
Before investigating the smoothness of the pressure map, we give a preliminary result estimating the size of eventual discontinuities of each of the static pressure maps $\mu\mapsto P_{\mu}(g,\mathcal{P}_{0}^{n})\coloneqq\frac{1}{n}H_{\mu}(\mathcal{P}_{0}^{n})+\int g\,{\rm d}\mu$. Note that by invariance $P_{\mu}(g,\mathcal{P}_{0}^{n})\coloneqq\frac{1}{n}\left(H_{\mu}(\mathcal{P}_{0}^{n})+\int S_{n}g\,{\rm d}\mu\right)$.
Proposition 5.1.
For all $\mu\in\mathcal{M}(M,T)$, all potential $g:M\to\mathbbm{R}$ $\mathcal{M}_{0}^{1}$-continuous, all small enough $\varepsilon>0$ and all large enough $n$, there is a neighbourhood $V_{\varepsilon,n}\subset\mathcal{M}(M,T)$ of $\mu$ such that for all $\nu\in V_{\varepsilon,n}$,
$P_{\nu}(g,\mathcal{P}_{0}^{n})-P_{\mu}(g,\mathcal{P}_{0}^{n})$ $\leqslant\frac{\mu(\mathcal{S})}{n}\log\sum_{A\in\mathcal{P}_{0}^{n}}\sup_{A}e^{S_{n}g}-\int_{\mathcal{S}}g\,{\rm d}\mu+R_{\varepsilon,n}(g),$
where
$R_{\varepsilon,n}(g)=\varepsilon\left(4\left|\frac{1}{n}\log\sum_{A\in\mathcal{P}_{0}^{n}}\sup_{A}e^{S_{n}g}\right|+\sup|g|+2\frac{\log\#\mathcal{P}_{0}^{n}}{n}\right)+\frac{2\log 2+e^{-1}+5\varepsilon}{n}+2\varepsilon.$
To prove the above proposition, we split the contributions to $P_{\mu}(g,\mathcal{P}_{0}^{n})$ of a neighbourhood of $\partial\mathcal{P}_{0}^{n}$ and its complement. To do so, we rely on the use of conditional entropy and conditional pressure. Recall that for any two finite partitions $\mathcal{A}$ and $\mathcal{C}$, the conditional entropy of $\mathcal{A}$ with respect to $\mathcal{C}$ and $\mu$ is defined by
$H_{\mu}(\mathcal{A}\mid\mathcal{C})\coloneqq\sum_{C\in\mathcal{C}}\mu(C)H_{\mu}(\mathcal{A}\mid C)\quad\text{with }H_{\mu}(\mathcal{A}\mid C)\coloneqq H_{\mu_{C}}(\mathcal{A}),$
where $\mu_{C}$ is the conditional probability of $\mu$ on $C$. By convention, if $\mu(C)=0$, we set $H_{\mu}(\mathcal{A}\mid C)=0$. Analogously, we also denote
$P_{\mu}(g,\mathcal{A}\mid\mathcal{C})\coloneqq\sum_{C\in\mathcal{C}}\mu(C)P_{\mu}(g,\mathcal{A}\mid C)\quad\text{with }P_{\mu}(g,\mathcal{A}\mid C)\coloneqq\frac{1}{n}\left(H_{\mu_{C}}(\mathcal{A})+\int S_{n}g\,{\rm d}\mu_{C}\right).$
In order to estimate the contribution away from singularities in the static entropy part, we need the following technical lemma.
Lemma 5.2.
For all $x,y\in[0,1]$ if $|x-y|\leqslant 1/2$, then $|x\log x-y\log y|\leqslant-|x-y|\log|x-y|$.
Proof of Lemma 5.2.
Let $f_{\delta}(x)=-x\log x+(x+\delta)\log(x+\delta)$, for $20\leqslant\delta\leqslant 1/$ 2 and $x\in[0,1-\delta]$. Differentiating, we get $f^{\prime}_{\delta}(x)=\log((x+\delta)/x)>0$ on $(0,1-\delta)$. Thus
$\delta\log\delta=f_{\delta}(0)\leqslant f_{\delta}\leqslant f_{\delta}(1-\delta)=-(1-\delta)\log(1-\delta).$
Let $g(\delta)=-\delta\log\delta+(1-\delta)\log(1-\delta)$, for $20\leqslant\delta\leqslant 1/$ 2. Differentiating twice, we get $g^{\prime\prime}(\delta)=\frac{2\delta-1}{\delta(1-\delta)}<0$ on $(0,1/2)$. Thus $g$ is concave on $[0,1/2]$ and $g(\delta)\geqslant 2\delta g(1/2)+(1-2\delta)g(0)=0$. In other words, $-(1-\delta)\log(1-\delta)\leqslant-\delta\log\delta$. Taking $x\leqslant y$ and $\delta=y-x$ ends the proof. ∎
Proof of Proposition 5.1.
Let $\varepsilon>0$ and $g:M\to\mathbbm{R}$ a $\mathcal{M}_{0}^{n}$-continuous potential. First, remark that for any partition $\alpha$ of $M$ into two elements, we have on the one hand, for all $\nu\in\mathcal{M}(M,T)$ and $n\geqslant 1$,
$H_{\nu}(\mathcal{P}_{0}^{n}\vee\alpha)=H_{\nu}(\alpha)+H_{\nu}(\mathcal{P}_{0}^{n}\mid\alpha),$
where $0\leqslant H_{\nu}(\alpha)\leqslant\log 2$ by [33, Corollary 4.2.1], and on the other hand,
$H_{\nu}(\mathcal{P}_{0}^{n}\vee\alpha)=H_{\nu}(\mathcal{P}_{0}^{n})+H_{\nu}(\alpha\mid\mathcal{P}_{0}^{n}),$
where, $0\leqslant H_{\nu}(\alpha\mid\mathcal{P}_{0}^{n})\leqslant H_{\nu}(\alpha)\leqslant\log 2$ by [33, Theorem 4.3(vi)]. Thus, for any partition $\alpha$ of $M$ with #$\#\alpha=2$ and any $T$-invariant measure $\nu$,
$\left|P_{\nu}(g,\mathcal{P}_{0}^{n})-P_{\nu}(g,\mathcal{P}_{0}^{n}\mid\alpha)\right|\leqslant\frac{\log 2}{n}.$
Let $\mu\in\mathcal{M}(M,T)$. By Lemma 4.5, it follows that $\mu(\mathcal{S})=\mu(\bigcup_{n\geqslant 0}S_{n})$. Let $n$ be large enough so that $|\mu(\mathcal{S})-\mu(\mathcal{S}_{n+1}\cup\mathcal{S}_{-1})|$ is less than $\varepsilon$, and less than $\varepsilon/\sup g$ if $\sup g>0$.
Let $\delta$ be small enough so that $|\mu(\mathcal{S}_{n+1}\cup\mathcal{S}_{-1})-\mu(\mathcal{N}_{\delta}(\mathcal{S}_{n+1}\cup\mathcal{S}_{-1}))|$ is smaller than $\varepsilon$, and smaller that $\varepsilon/(-\inf g)$ if $\inf g<0$. Notice that, up to decreasing the value of $\delta$, we can assume that $\mu(\partial\mathcal{N}_{\delta}(\mathcal{S}_{n+1}\cup\mathcal{S}_{-1})\smallsetminus\mathcal{S}_{0})=0$. Now, denote $U=\mathcal{N}_{\delta}(\mathcal{S}_{n+1}\cup\mathcal{S}_{-1})$ and consider the partition $\alpha_{\varepsilon,n}=\{U,M\smallsetminus U\}$ into two elements. Similarly, let $\delta^{\prime}<\delta$ and $U^{\prime}=\mathcal{N}_{\delta^{\prime}}(\mathcal{S}_{n+1}\cup\mathcal{S}_{-1})$ such that $\mu(\partial U^{\prime}\smallsetminus\mathcal{S}_{0})=0$ and $\mu(U\smallsetminus U^{\prime})<\varepsilon/(n\sup_{M}|g|)$. Finally, let $\varphi\in C^{0}(M,\mathbbm{R})$ be such that $\varphi|_{U^{c}}\equiv 1$ and $\varphi|_{U^{\prime}}\equiv 0$.
Therefore, for any $\nu\in\mathcal{M}(M,T)$,
(5.1) $P_{\nu}(g,\mathcal{P}_{0}^{n})-P_{\mu}(g,\mathcal{P}_{0}^{n})\leqslant\frac{2\log 2}{n}+\left(P_{\nu}(g,\mathcal{P}_{0}^{n}\mid\alpha_{\varepsilon,n})-P_{\mu}(g,\mathcal{P}_{0}^{n}\mid\alpha_{\varepsilon,n})\right).$
We want to estimate the right hand side. First, notice that since for all $A\in\mathcal{P}_{0}^{n}$, $\partial(A\smallsetminus U)\subset\partial U\smallsetminus\mathcal{S}_{0}$ has $\mu$-measure zero, there is a neighbourhood $V_{A}$ of $\mu$ such that for any $\nu\in V_{A}$,
$\left|\nu(A\smallsetminus U)-\mu(A\smallsetminus U)\right|\leqslant\frac{\varepsilon}{\#\mathcal{P}_{0}^{n}}.$
Also, since $\mu(\partial U\smallsetminus\mathcal{S}_{0})=0$ and $\mathcal{S}_{0}=\partial M$, there is a neighbourhood $\tilde{V}_{1}$ of $\mu$ such that for any $\nu\in\tilde{V}_{1}$,
$\left|\nu(U)-\mu(U)\right|\leqslant\varepsilon.$
Since $\mu(\partial(M\smallsetminus U))=0$ and $\mu(\partial(U\smallsetminus U^{\prime}))=0$ there is a neighbourhood $\tilde{V}_{2}$ of $\mu$ such that for any $\nu\in\tilde{V}_{2}$,
$\left|\nu(M\smallsetminus U)-\mu(M\smallsetminus U)\right|\leqslant\varepsilon,\quad\text{ and }\left|\nu(U\smallsetminus U^{\prime})-\mu(U\smallsetminus U^{\prime})\right|\leqslant\frac{\varepsilon}{n\sup|g|}.$
Finally, since $\varphi\,S_{n}g$ is continuous, there is a neighbourhood $\tilde{V}_{3}$ of $\mu$ such that for any $\nu\in\tilde{V}_{3}$,
$\left|\int\varphi\,S_{n}g\,{\rm d}\mu-\int\varphi\,S_{n}g\,{\rm d}\nu\right|\leqslant\varepsilon.$
Define $V_{\varepsilon,n}\coloneqq\tilde{V}_{1}\cap\tilde{V}_{2}\cap\tilde{V}_{3}\cap\bigcap_{A\in\mathcal{P}_{0}^{n}}V_{A}$. Note that $V_{\varepsilon,n}$ is a neighbourhood of $\mu$.
In the case where $\mu(M\smallsetminus U)>0$, we will need to consider an even smaller neighbourhood. More precisely, notice that if the value of $\varepsilon$ decreases (as well as $\delta$, accordingly), then $\mu(M\smallsetminus U)$ increases. Thus, if $\mu(M\smallsetminus U_{\varepsilon})>0$ for some $\varepsilon$, there is an $\tilde{\varepsilon}>0$ such that for all $0<\varepsilon<\tilde{\varepsilon}$, $22\varepsilon/\mu(M\smallsetminus U)<1/$ 2. Furthermore, there is a neighbourhood $\tilde{V}_{4}$ of $\mu$ such that for all $\nu\in\tilde{V}_{4}$
$\left|\frac{\mu(M\smallsetminus U)}{\nu(M\smallsetminus U)}-1\right|\leqslant\frac{\varepsilon}{n\sup_{M}|g|}.$
If $\mu(M\smallsetminus U)>0$, replace $V_{\varepsilon,n}$ by $V_{\varepsilon,n}\cap\tilde{V}_{4}$ and take $\varepsilon<\tilde{\varepsilon}$.
Expanding the second term of the right-hand-side of (5.1) and assuming $\nu\in V_{\varepsilon,n}$, we get
(5.2) $\begin{split}&P_{\nu}(g,\mathcal{P}_{0}^{n}\mid\alpha_{\varepsilon,n})-P_{\mu}(g,\mathcal{P}_{0}^{n}\mid\alpha_{\varepsilon,n})\leqslant\left(\mu(U)P_{\mu}(g,\mathcal{P}_{0}^{n}\mid U)+\nu(U)P_{\nu}(g,\mathcal{P}_{0}^{n}\mid U)\right)\\ &\qquad\qquad\qquad\qquad\qquad+\left(\mu(M\smallsetminus U)P_{\mu}(g,\mathcal{P}_{0}^{n}\mid M\smallsetminus U)-\nu(M\smallsetminus U)P_{\nu}(g,\mathcal{P}_{0}^{n}\mid M\smallsetminus U)\right).\end{split}$
We begin with the contributions of $U$. Note that
$\mu(U)P_{\mu}(g,\mathcal{P}_{0}^{n}\mid U)$ $\geqslant\mu(U)\int\frac{1}{n}S_{n}g\,{\rm d}\mu_{U}=\int_{U}\frac{1}{n}S_{n}g\,{\rm d}\mu$
$\geqslant\int_{\mathcal{S}}\frac{1}{n}S_{n}g\,{\rm d}\mu-\int_{\mathcal{S}\smallsetminus(\mathcal{S}_{n+1}\cup\mathcal{S}_{-1})}\frac{1}{n}S_{n}g\,{\rm d}\mu+\int_{U\smallsetminus(\mathcal{S}_{n+1}\cup\mathcal{S}_{-1})}\frac{1}{n}S_{n}g\,{\rm d}\mu$
$\geqslant\int_{\mathcal{S}}g\,{\rm d}\mu-2\varepsilon,$
and,
$\nu(U)P_{\nu}(g,\mathcal{P}_{0}^{n}\mid U)\leqslant(\mu(\mathcal{S})\pm 3\varepsilon)\frac{1}{n}\log\sum_{A\in\mathcal{P}_{0}^{n}}\sup_{A}e^{S_{n}g},$
where the ±$\pm 3\varepsilon$ stands for $+3\varepsilon$ if the $\log$ term is positive and $-3\varepsilon$ otherwise. Therefore, the contributions of $U$ in (5.2) can be bounded by
$\mu(U)P_{\mu}(g,\mathcal{P}_{0}^{n}\mid U)+\nu(U)P_{\nu}(g,\mathcal{P}_{0}^{n}\mid U)\leqslant(\mu(\mathcal{S})\pm 3\varepsilon)\frac{1}{n}\log\sum_{A\in\mathcal{P}_{0}^{n}}\sup_{A}e^{S_{n}g}-\int_{\mathcal{S}}g\,{\rm d}\mu+2\varepsilon.$
We now consider the contributions of $M\smallsetminus U$, that is, the second term in (5.2). This contribution can be further split as follows
(5.3) $\begin{split}\nu(M\smallsetminus U)P_{\nu}(g,\mathcal{P}_{0}^{n}&\mid M\smallsetminus U)-\mu(M\smallsetminus U)P_{\mu}(g,\mathcal{P}_{0}^{n}\mid M\smallsetminus U)\\ &=\left(\nu(M\smallsetminus U)-\mu(M\smallsetminus U)\right)P_{\mu}(g,\mathcal{P}_{0}^{n}\mid M\smallsetminus U)\\ &\quad+\nu(M\smallsetminus U)\left(P_{\nu}(g,\mathcal{P}_{0}^{n}\mid M\smallsetminus U)-P_{\mu}(g,\mathcal{P}_{0}^{n}\mid M\smallsetminus U)\right).\end{split}$
The first term on the right-hand side can be bounded by
$\left(\nu(M\smallsetminus U)-\mu(M\smallsetminus U)\right)P_{\mu}(g,\mathcal{P}_{0}^{n}\mid M\smallsetminus U)$ $=\varepsilon\left(\left|\frac{1}{n}\log\sum_{A\in\mathcal{P}_{0}^{n}}\sup_{A}e^{S_{n}g}\right|+\left|\int\frac{1}{n}S_{n}g\,{\rm d}\mu_{U^{c}}\right|\right)$
$\leqslant\varepsilon\left(\left|\frac{1}{n}\log\sum_{A\in\mathcal{P}_{0}^{n}}\sup_{A}e^{S_{n}g}\right|+\sup|g|\right),$
whereas the second one is bounded by
$\nu(M\smallsetminus U)$ $\left(P_{\nu}(g,\mathcal{P}_{0}^{n}\mid M\smallsetminus U)-P_{\mu}(g,\mathcal{P}_{0}^{n}\mid M\smallsetminus U)\right)$
$\leqslant\frac{\nu(M\smallsetminus U)}{n}\sum_{A\in\mathcal{P}_{0}^{n}}\left|-\frac{\mu(A\smallsetminus U)}{\mu(M\smallsetminus U)}\log\frac{\mu(A\smallsetminus U)}{\mu(M\smallsetminus U)}+\frac{\nu(A\smallsetminus U)}{\nu(M\smallsetminus U)}\log\frac{\nu(A\smallsetminus U)}{\nu(M\smallsetminus U)}\right|$
$\qquad+\frac{\nu(M\smallsetminus U)}{n}\left|\int S_{n}g\,{\rm d}\mu_{U^{c}}-\int S_{n}g\,{\rm d}\nu_{U^{c}}\right|.$
Now, since
$S\coloneqq$ $\sum_{A\in\mathcal{P}_{0}^{n}}\left|\frac{\mu(A\smallsetminus U)}{\mu(M\smallsetminus U)}-\frac{\nu(A\smallsetminus U)}{\nu(M\smallsetminus U)}\right|$
$\qquad\qquad\quad\leqslant\frac{1}{\mu(M\smallsetminus U)}\sum_{A\in\mathcal{P}_{0}^{n}}\left|\left(\frac{\mu(M\smallsetminus U)}{\nu(M\smallsetminus U)}-1\right)\nu(A\smallsetminus U)\right|+\left|\mu(A\smallsetminus U)-\nu(A\smallsetminus U)\right|$
$\qquad\qquad\quad\leqslant\frac{1}{\mu(M\smallsetminus U)}\sum_{A\in\mathcal{P}_{0}^{n}}\left(\varepsilon\nu(A\smallsetminus U)+\frac{\varepsilon}{\#\mathcal{P}_{0}^{n}}\right)\leqslant\frac{2\varepsilon}{\mu(M\smallsetminus U)}<\frac{1}{2},$
we can apply Lemma 5.2 and use the Jensen inequality (for the last line) to get that
$\frac{\mu(M\smallsetminus U)}{n}\sum_{A\in\mathcal{P}_{0}^{n}}\left|-\frac{\mu(A\smallsetminus U)}{\mu(M\smallsetminus U)}\log\frac{\mu(A\smallsetminus U)}{\mu(M\smallsetminus U)}+\frac{\nu(A\smallsetminus U)}{\nu(M\smallsetminus U)}\log\frac{\nu(A\smallsetminus U)}{\nu(M\smallsetminus U)}\right|$
$\quad\leqslant\frac{\mu(M\smallsetminus U)}{n}\sum_{A\in\mathcal{P}_{0}^{n}}-\left|\frac{\mu(A\smallsetminus U)}{\mu(M\smallsetminus U)}-\frac{\nu(A\smallsetminus U)}{\nu(M\smallsetminus U)}\right|\log\left|\frac{\mu(A\smallsetminus U)}{\mu(M\smallsetminus U)}-\frac{\nu(A\smallsetminus U)}{\nu(M\smallsetminus U)}\right|$
$\quad\leqslant\frac{\mu(M\smallsetminus U)}{n}S\left(\log\#\mathcal{P}_{0}^{n}-\log S\right)\leqslant\frac{2\varepsilon}{n}\log\#\mathcal{P}_{0}^{n}+\frac{e^{-1}}{n}.$
We now turn to the difference of integrals term.
$\mu(M\smallsetminus U)$ $\left|\int S_{n}g\,{\rm d}\mu_{U^{c}}-\int S_{n}g\,{\rm d}\nu_{U^{c}}\right|$
[…]
$\leqslant n\sup|g|(\mu(U\smallsetminus U^{\prime})+\nu(U\smallsetminus U^{\prime}))+2\varepsilon\leqslant 5\varepsilon$
Thus, combining the above estimates inside (5.3), we get that the contribution of $M\smallsetminus U$ is at most
$\nu(M\smallsetminus U)$ $P_{\nu}(g,\mathcal{P}_{0}^{n}\mid M\smallsetminus U)-\mu(M\smallsetminus U)P_{\mu}(g,\mathcal{P}_{0}^{n}\mid M\smallsetminus U)$
$\leqslant\varepsilon\left(\left|\frac{1}{n}\log\sum_{A\in\mathcal{P}_{0}^{n}}\sup_{A}e^{S_{n}g}\right|+\sup|g|+2\frac{\log\#\mathcal{P}_{0}^{n}}{n}+\frac{5}{n}\right)+\frac{e^{-1}}{n}.$
Combining the estimates of the contributions of $U$ and $M\smallsetminus U$ inside (5.2), we deduce the claim from (5.1). ∎
We are now able to prove the main bound of this section.
Corollary 5.3.
For all finite horizon Sinai billiard, all $T$-invariant probability measure $\mu$, and all $\mathcal{M}_{0}^{1}$-continuous potential $g$,
$\limsup_{\nu\rightharpoonup\mu}P_{\nu}(T,g)\leqslant P_{\mu}(T,g)+\mu(\mathcal{S})\left(P_{\rm top}(T,g)-P_{\mu_{\mathcal{S}}}(T,g)\right),$
[…]
Proof.
Let $\mu\in\mathcal{M}(M,T)$ and $\varepsilon>0$ small enough as in Proposition 5.1. Since $\mathcal{P}$ is a generating partition, we can choose $n=n(\varepsilon)$ large enough so that
$h_{\mu}(T)\geqslant\frac{1}{n}H_{\mu}(\mathcal{P}_{0}^{n})-\varepsilon.$
Let $V_{\varepsilon,n}$ be the neighbourhood of $\mu$ given by Proposition 5.1 (up to increasing the value of $n$). Therefore
$P_{\mu}(T,g)$ $=h_{\mu}(T)+\mu(g)\geqslant\frac{1}{n}H_{\mu}(\mathcal{P}_{0}^{n})-\varepsilon+\mu(g)=P_{\mu}(g,\mathcal{P}_{0}^{n})-\varepsilon$
$\geqslant\sup_{\nu\in V_{\varepsilon,n}}P_{\nu}(g,\mathcal{P}_{0}^{n})-\varepsilon-R_{\varepsilon,n}(g)$
$\geqslant\sup_{\nu\in V_{\varepsilon,n}}P_{\nu}(T,g)-\varepsilon-R_{\varepsilon,n}(g),$
where we used that $P_{\nu}(T,g)=\inf_{n}P_{\nu}(g,\mathcal{P}_{0}^{n})$. We can choose $n=n(\varepsilon)$ such that, taking the limit $\varepsilon\to 0$, we get $n(\varepsilon)\to+\infty$. Therefore, taking the limit as $\varepsilon$ goes to zero of the previous expression yields
$P_{\mu}(T,g)\geqslant\limsup_{\nu\rightharpoonup\mu}P_{\nu}(T,g)-\mu(\mathcal{S})P_{\rm top}(T,g)+\int_{\mathcal{S}}g\,{\rm d}\mu\,.$
Finally, since $\mu_{\mathcal{S}}$ is atomic, it follows that $h_{\mu_{\mathcal{S}}}(T)=0$. Hence
$P_{\mu_{\mathcal{S}}}(T,g)=\int g\,{\rm d}\mu_{\mathcal{S}}=\frac{1}{\mu(\mathcal{S})}\int_{\mathcal{S}}g\,{\rm d}\mu,$
which finishes the proof. ∎
5.2. Maximizing sequences and Equilibrium States
In this part we prove that is if there exists a maximizing sequence for the pressure for a potential with small enough amplitude, then there exist associated equilibrium state. In order to give a bound on the amplitude of the potential, we need to following lemma, controlling the metric entropy on a neighbourhood of $\mathcal{M}_{\mathcal{S}}(M,T)$.
Lemma 5.4.
For all $\mu\in\mathcal{M}_{\mathcal{S}}(M,T)$ and all $\varepsilon>0$, there exists an open neighbourhood $U_{\mu}\subset\mathcal{M}(M,T)$ of $\mu$ such that $\sup\{h_{\nu}(T)\mid\nu\in U_{\mu}\}\leqslant\varepsilon$.
Note that measures arbitrarily close to $\mathcal{M}(M,T)$ of positive entropy have been constructed in [14].
Proof.
Let $\mu\in\mathcal{M}_{\mathcal{S}}(M,T)$. Then, by Lemma 4.5, $\mu=\sum_{i\in I}\omega_{i}\delta_{\mathcal{O}(x_{i})}$, where $I$ is at most countable. We first consider the case where $I$ is infinite. We identify $I$ with $\mathbbm{N}$.
Let $\varepsilon>0$ and $i_{0}$ large enough so that $\sum_{i>i_{0}}\omega_{i}<\varepsilon$. Since $\mathcal{O}_{\mu}\coloneqq\bigcup_{i=1}^{i_{0}}\mathcal{O}(x_{i})\subset M$ is a finite set and $\lim_{n\to\infty}\mbox{diam}(\mathcal{P}_{-n}^{n})=0$, there exists $n_{0}$ such that for all $n\geqslant n_{0}$ and all $A\in\mathcal{P}_{-n}^{n}$, #$\#\bar{A}\cap\mathcal{O}_{\mu}\leqslant 1$. Notice that there are at most $KnC_{0}$ elements $A$ of $\mathcal{P}_{-n}^{n}$ where the equality holds, where $C_{0}=\#\mathcal{O}_{\mu}$. Define $U_{n}$ the interior of the union of those elements of $\mathcal{P}_{-n}^{n}$. For any $x\in\mathcal{O}_{\mu}$, we write $U_{n}(x)$ the connected component of $U_{n}$ containing $x$. Up to increase the value of $n$, we can assume that the $U_{n}(x)$ are pairwise disjoint.
Since each $U_{n}(x),x\in\mathcal{O}_{\mu}$ is open, there is an open neighbourhood $V_{\varepsilon,n}$ of $\mu$ such that for all $\nu\in V_{\varepsilon,n}$, $\nu(U(x))>\mu(U(x))-\varepsilon/C_{0}$. In particular, by choice of $i_{0}$, $1-\nu(U_{n})\leqslant 2\varepsilon$. Therefore, for any $\nu\in V_{\varepsilon,n}$,
$h_{\nu}(T)$ $\leqslant\frac{1}{n}H_{\nu}(\mathcal{P}_{0}^{n})=\frac{1}{n}\nu(U_{n})H_{\nu}(\mathcal{P}_{0}^{n}|U_{n})+\frac{1}{n}(1-\nu(U_{n}))H_{\nu}(\mathcal{P}_{0}^{n}|U_{n}^{c})$
$\leqslant\frac{\log KnC_{0}}{n}+2\varepsilon\frac{\log\mathcal{P}_{0}^{n}}{n}\leqslant 3\varepsilon h_{*},$
where the last inequality holds as long as $n$ is large enough. The lemma follows by setting $U_{\mu}=V_{\varepsilon/3h_{*},n}$ for some large enough $n$. The case where $I$ is finite is analogous. ∎
Corollary 5.5.
Let $g$ be a $\mathcal{M}_{0}^{n}$-continuous potential such that $h_{*}>\sup g-\inf g$. If there exists a sequence of measure $\mu_{k}\in\mathcal{M}(M,T)$ such that
$P_{\rm top}(T,g)=\lim_{k\to\infty}P_{\mu_{k}}(T,g),$
then there exist equilibrium states associated to $g$ for the collision map. Furthermore, the variational principle holds.
Proof.
Let $g$ and $\mu_{k}$ as in the statement of the corollary. By compactness, we can assume that $\mu_{k}$ converges to some invariant measure $\mu$. Therefore, by Corollary 5.3,
$P_{\rm top}(T,g)=\lim_{k\to\infty}P_{\mu_{k}}(T,g)\leqslant P_{\mu}(T,g)+\mu(\mathcal{S})(P_{\rm top}(T,g)-P_{\mu_{\mathcal{S}}}(T,g)).$
Now, using that the Kolmogorov–Sinai entropy is harmonic, it follows that
$P_{\mu}(T,g)=(1-\mu(\mathcal{S}))P_{\mu_{\mathcal{S}^{c}}}(T,g)+\mu(\mathcal{S})P_{\mu_{\mathcal{S}}}(T,g).$
Thus,
$(1-\mu(\mathcal{S}))P_{\rm top}(T,g)\leqslant(1-\mu(\mathcal{S}))P_{\mu_{\mathcal{S}^{c}}}(T,g).$
In particular, either $\mu(\mathcal{S})=1$ (and $\mu_{\mathcal{S}^{c}}$ is not defined) or $P_{\mu_{\mathcal{S}^{c}}}=P_{\rm top}(T,g)$.
By contradiction, assume that $\mu(\mathcal{S})=1$. Let $\varepsilon>0$ be small enough so that $h_{*}>\sup g-\inf g+2\varepsilon$. By Lemma 5.4, for any $k$ large enough $h_{\mu_{k}}(T)<\varepsilon$. Hence
$P_{\mu_{k}}(T,g)\leqslant\varepsilon+\sup g<h_{*}+\inf g-\varepsilon\leqslant P_{\rm top}(T,g)-\varepsilon,$
which contradicts the assumption on the sequence $\mu_{k}$. ∎
Corollary 5.6.
The Kolmogorov–Sinai entropy is upper semi-continuous at every ergodic measure.
Proof.
Let $\mu$ be an ergodic invariant measure. Since $\mathcal{S}$ is invariant, either $\mu(\mathcal{S})=0$ or $\mu(\mathcal{S}^{c})=0$. In the first case, it follows from Corollary 5.3 that the Kolmogorov–Sinai entropy is continuous at $\mu$. In the second case, it must be that $\mu\in\mathcal{M}_{\mathcal{S}}(M,T)$. By Lemma 4.5, since $\mu$ is ergodic, it is supported by a single periodic orbit. Hence $h_{\mu}(T)=0$. The upper-semi continuity of the Kolomogorov–Sinai entropy at $\mu$ follows from Lemma 5.4. ∎
5.3. Distribution of periodic orbits
Before focusing on the distribution of periodic orbits, we need an upper-bound on #$\#{\rm Fix}\,T^{n}$. This is the content of the following lemma [^2] The estimate (5.4) is due to Mark Demers. The author is grateful to him for allowing to use his proof. [^3] Equation (5.5) was initially state for $g\equiv 0$. The author is grateful to Yuri Lima for his explanations on how to generalise it.
Lemma 5.7.
For all $n\geqslant 2$ and all $A\in\mathcal{P}_{0}^{n}$, $A$ contains at most one periodic orbit of period $n$. In particular, for any $\mathcal{M}_{0}^{1}$-piecewise continuous potentials
(5.4) $\sum_{x\in{\rm Fix}\,T^{n}}e^{S_{n}g(x)}\leqslant\sum_{A\in\mathcal{P}_{0}^{n}}\sup_{A}e^{S_{n}g}.$
Assuming that $g$ is piecewise Hölder and has sparse recurrence, that is, if $P_{\rm top}(T,g)-\sup g>s_{0}\log 2$, it exists a constant $C>0$ such that for all $n\geqslant 2$,
(5.5) $C^{-1}e^{nP_{\rm top}(T,g)}\leqslant\sum_{x\in{\rm Fix}\,T^{n}}e^{S_{n}g(x)}\leqslant Ce^{nP_{\rm top}(T,g)}$
This lemma gives a positive answer to [20, Problem 5].
Proof.
Let $n\geqslant 2$, $A\in\mathcal{P}_{0}^{n}$ and $x,y\in{\rm Fix}\,T^{n}\cap A$. Since $T^{n}$ is continuous on $A$, we get that $T^{n}x=x$ and $T^{n}y=y$ both belong to $T^{n}A\in\mathcal{P}_{-n}^{0}$. Since $A\cap T^{n}A\in\mathcal{P}_{-n}^{n}$, applying $T^{-n}$ we get that $x$ and $y$ belong to the same element of $\mathcal{P}_{0}^{2n}$. Repeating this argument, we obtain that $x$ and $y$ belong to the same element of $\mathcal{P}_{-2^{k}n}^{2^{k}n}$ for all $k\geqslant 0$. Since $\mbox{diam}\mathcal{P}_{-k}^{k}$ goes to zero as $k$ goes to infinity, it must be that $x=y$. From this, (5.4) follows.
Now, in the case $g\equiv 0$, from sparse recurrence it follows from [3, Proposition 4.6] that there is a constant $C>0$ such that #$\#\mathcal{P}_{0}^{n}\leqslant Ce^{nh_{*}}$. The lower estimate on #$\#{\rm Fix}\,T^{n}$ is the content of [10, Theorem 1.5] (itself bootstrapping from [3, Corollary 2.7] and [24, Theorem 1.3]).
For more a general potential, as in the statement of the lemma, a similar strategy can be applied. Indeed, the upper estimate is the content of [11, Proposition 3.10] (note that sparse recurrence implies the small singular pressure assumptions, as explained in [2]). For the lower estimate, we also rely on the coding $\pi:(\Sigma,\sigma)\to(M,T)$ constructed in [24, Theorem 1.3] with hyperbolicity parameter related to the unique equilibrium state $\mu_{g}$ associated to $g$.
According to [11, Corollary 6.3], $\mu_{g}$ is $T$-adapted in the sense of [24]. Since it is $\chi$-hyperbolic from the choice of $\chi$, [24, Theorem 1.3] implies the existence of $\hat{\mu}_{g}\in\mathcal{M}_{\rm erg}(\Sigma^{\#},\sigma)$ such that $\pi_{*}(\hat{\mu}_{g})=\mu_{g}$. Since $\pi_{|\Sigma^{\#}}$ is finite-to-one, $\pi_{*}$ preserves the entropy, and thus the metric pressure. Therefore, $\hat{\mu}_{g}$ is an equilibrium state for $\Sigma$.
From [30, Theorem 4.6], since the potential $\hat{g}\coloneqq g\circ\pi$ admits $\hat{\mu}_{g}$ as an equilibrium state, it is positively recurrent. Thus, by the generalised Ruelle’s Perron–Frobenius theorem (see [30, Theorem 3.4]), we get that there is a cylinder $[a]$ such that the following convergence holds,
$\lambda^{-n}L_{\hat{g}}^{n}\mathbbm{1}_{[a]}\to c>0,$
where $L_{\hat{g}}$ is the weighted transfer operator associated to $\sigma$ and $\hat{g}$, and $\lambda=\exp(P_{G}(\hat{g}))$ where $P_{G}(\hat{g})$ is the Gurevich pressure of $\hat{g}$. Note that by the existence of $\hat{\mu}_{g}$, $P_{G}(\hat{g})\geqslant P_{\rm top}(T,g)$. Now, since
$(\mathbbm{1}_{[a]}L_{\hat{g}}^{n}\mathbbm{1}_{[a]})(X)=\sum_{Y\in\sigma^{-n}(X)}e^{S_{n}\hat{g}(Y)},$
where in the sum $Y$ ranges over all the sequences of the form $(y_{0},\ldots,y_{n-1},X)$, with $y_{0}=y_{n}=a$. Since $g$ and $\pi$ are both Hölder, $\hat{g}$ is also Hölder. Thus, there is a constant $K>0$ such that $|S_{n}\hat{g}(Y)-S_{n}\hat{g}(Z_{Y})|\leqslant K$, where $Z_{Y}$ is the periodic orbit of period $n$ whose $n$-th first coordinates coincide with the ones of $Y$. Thus,
(5.6) $\sum_{Z\in[a]\cap{\rm Fix}\,\sigma^{n}}e^{S_{n}\hat{g}(Z)}\geqslant\frac{c}{2}e^{-K}e^{nP_{G}(\hat{g})}$
for all $n$ large enough. Up to decreasing the value of $c$, the claimed result holds for all $n\geqslant 2$. ∎
In this subsection, given a potential $g$, we consider the measures supported on periodic orbits, denoted
$\mu_{n}=\frac{1}{\sum_{x\in{\rm Fix}\,T^{n}}e^{S_{n}g(x)}}\sum_{x\in{\rm Fix}\,T^{n}}e^{S_{n}g(x)}\delta_{x}.$
Corollary 5.8.
Assume that sparse recurrence holds for a $\mathcal{M}_{0}^{1}$-piecewise Hölder potential $g$ and let $\mu$ be an accumulation point of the sequence $\mu_{n}$. If $\mu(\mathcal{S})<1$, then $\mu_{\mathcal{S}^{c}}=\mu_{g}$, where $\mu_{\mathcal{S}^{c}}\in\mathcal{M}(M,T)$ is the conditional measure of $\mu$ on $\mathcal{S}^{c}$
Further assuming that the singularities are negligible, then $\mu_{n}\rightharpoonup\mu_{g}$ as $n$ goes to infinity. In particular, when $g\equiv 0$, the periodic orbits equidistribute with respect to the unique measure of maximal entropy.
Proof.
From Lemma 5.7, it follows that for all $A\in\mathcal{P}_{0}^{n}$, $\mu_{n}(A)$ is equal to either to 00 or to $1/\sum_{x\in{\rm Fix}\,T^{n}}e^{S_{n}g(x)}$. Therefore, assuming sparse recurrence,
$\frac{1}{n}H_{\mu_{n}}(\mathcal{P}_{0}^{n})+\int g\,{\rm d}\mu_{n}$ $=-\frac{1}{n}\sum_{x\in{\rm Fix}\,T^{n}}\mu_{n}(\{x\})\left(\log\mu_{n}(\{x\})-S_{n}g(x)\right)$
$=\frac{1}{n}\log\sum_{x\in{\rm Fix}\,T^{n}}e^{S_{n}g(x)}\geqslant P_{\rm top}(T,g)-\frac{\log C}{n}.$
For any $q\leqslant n$, write the Euclidean division of $n$ by $q$, $n=j_{0}q+r_{0}$. Now, since $\mathcal{P}_{0}^{n}=\bigvee_{j=0}^{j_{0}-1}T^{-jq}\mathcal{P}_{0}^{q}\vee\bigvee_{i=j_{0}q+1}^{n}T^{-i}\mathcal{P}$, we get by sub-additivity and $T$-invariance of $\mu_{n}$,
$\frac{1}{n}H_{\mu_{n}}(\mathcal{P}_{0}^{n})$ $\leqslant\frac{j_{0}}{n}H_{\mu_{n}}(\mathcal{P}_{0}^{q})+\frac{r_{0}}{n}H_{\mu_{n}}(\mathcal{P})\leqslant\frac{1}{q}H_{\mu_{n}}(\mathcal{P}_{0}^{q})+\frac{q}{n}\log\#\mathcal{P}.$
Consider $n_{k}$ and $\mu$ such that $\mu_{n_{k}}\operatorname{\rightharpoonup}\mu$ as $n_{k}$ goes to infinity. Let $\varepsilon>0$. Up to extracting again, we can assume that each $\mu_{n_{k}}$ belong to the neighbourhood $V_{\varepsilon,n_{k}}$ of $\mu$ from Proposition 5.1. Therefore, combining the two above estimates, taking the limit $n_{k}\to\infty$ and applying Proposition 5.1, we get that
$P_{\rm top}(T,g)\leqslant P_{\mu}(g,\mathcal{P}_{0}^{q})+\frac{\mu(\mathcal{S})}{q}\log\sum_{A\in\mathcal{P}_{0}^{n}}\sup_{A}e^{S_{n}g}-\int_{\mathcal{S}}g\,{\rm d}\mu+R_{\varepsilon,q}(g).$
Now, taking successively the limits $q\to\infty$ and $\varepsilon\to 0$, and writing $\mu=(1-\mu(\mathcal{S}))\mu_{\mathcal{S}^{c}}+\mu(\mathcal{S})\mu_{S}$ yield
$P_{\rm top}(T,g)\leqslant(1-\mu(\mathcal{S}))P_{\mu_{\mathcal{S}^{c}}}(T,g)+\mu(\mathcal{S})P_{\rm top}(T,g),$
where we used that, since $\mu_{\mathcal{S}}\in\mathcal{M}_{\mathcal{S}}(M,T)$, $h_{\mu_{\mathcal{S}}}=0$. Therefore
$(1-\mu(\mathcal{S}))P_{\rm top}(T,g)=(1-\mu(\mathcal{S}))P_{\mu_{\mathcal{S}^{c}}}(T,g).$
Finally, if $\mu(\mathcal{S})=1$, then $\mu=\mu_{\mathcal{S}}$. Otherwise, $\mu_{\mathcal{S}^{c}}$ has maximal pressure. By [3, Theorem 2.4], the equilibrium state is unique. Therefore $\mu_{\mathcal{S}^{c}}=\mu_{g}$, where $\mu_{g}$ is the unique equilibrium state of $T$ and $g$. ∎
6. Topological tail entropy
Even though the relation between the upper semi-continuity of the entropy map and the topological tail entropy is not clear in the case of transformations with discontinuities, we give in this section an estimate on the topological tail entropy, as defined in (1.4), in Corollary 6.4. To do so, we split the contributions of points in $M\smallsetminus\mathcal{S}_{n}$ and $\mathcal{S}_{n}$ in Propositions 6.1 and 6.2 respectively.
First, recall that in the proof of [3, Lemma 3.4], Baladi and Demers proved that [^4] The definition of the Bowen metric $d_{n}$ in [3] corresponds here to $d_{n+1}$. there exists $\varepsilon_{0}>0$ with the property that for any $n\geqslant 0$, if $x,y$ lie in different elements of $\mathcal{M}_{0}^{n}$, then $d_{n+1}(x,y)\geqslant\varepsilon_{0}$.
Proposition 6.1.
There exists a constants $C>0$ such that for all $0<\varepsilon<\varepsilon_{0}$, all $\delta>0$, all $n\geqslant 2$,
$\log r(n,\delta,B(x,\varepsilon,n))\leqslant 2\frac{\log\#\mathcal{P}}{\log\Lambda}\log\delta^{-1}+C,\quad\forall x\in M\smallsetminus\mathcal{S}_{n-1}.$
In particular, taking the appropriate limits in $n$, $\delta$ and $\varepsilon$, notice that points of $M\smallsetminus\bigcup_{n\geqslant 0}\mathcal{S}_{n}$ don’t contribute to $h^{*}(T)$.
Proof.
Let $x\in M\smallsetminus\mathcal{S}_{n-1}$, $\varepsilon<\varepsilon_{0}$ and $\delta>0$. Let $\mathcal{M}_{0}^{n-1}(x)$ denote the unique element of $\mathcal{M}_{0}^{n-1}$ containing $x$. Then, $B(x,\varepsilon,n)\subset\mathcal{M}_{0}^{n-1}(x)$.
Note that if $F^{\prime}\subset B(x,\varepsilon,n)$ is $(\delta/2,n)$-separated of maximal cardinality, then $F^{\prime}$ is $(\delta,n)$-spanning $B(x,\varepsilon,n)$. In particular, $r(\delta,n,B(x,\varepsilon,n))\leqslant\#F^{\prime}$. Furthermore, if $F\subset\mathcal{M}_{0}^{n-1}(x)$ is $(\delta/2,n)$-separated with maximal cardinality, then $F$ is $(\delta,n)$-spanning and #$\#F^{\prime}\leqslant\#F$. Let $F$ be such a set.
Since $F$ is $(\delta/2,n)$-separated, according to the proof of [3, Lemma 3.4], each element of $\mathring{\mathcal{P}}_{-k_{\delta}}^{n-1+k_{\delta}}$ contains at most one point of $F$, where $k_{\delta}$ is the smallest integer such that $C_{1}\Lambda^{-k_{\delta}}<c_{1}\delta/2$, where $c_{1}$ and $C_{1}$ are constants coming respectively from [3, Lemma 3.4] and [3, eq. (3.1)]. Since $\mathcal{M}_{0}^{n-1}(x)$ contains at most #$\#\mathcal{M}_{-1}^{0}$ elements of $\mathcal{M}_{-1}^{n-1}$ and, from [3, Lemma 3.3], $\mathcal{M}_{-1}^{n-1}=\mathring{\mathcal{P}}_{0}^{n-2}$, we get that $\mathcal{M}_{0}^{n-1}(x)$ contains at most #$\#\mathcal{M}_{-1}^{0}\#\mathcal{P}_{-k_{\delta}}^{0}\#\mathcal{P}_{0}^{k_{\delta}+1}$ elements of $\mathring{\mathcal{P}}_{-k_{\delta}}^{n-1+k_{\delta}}$. Therefore
$r(n,\delta,B(x,\varepsilon,n))\leqslant\#F\leqslant\#\mathcal{M}_{-1}^{0}(\#\mathcal{P})^{2k_{\delta}+1}\leqslant\#\mathcal{M}_{-1}^{0}\#\mathcal{P}\exp\left(-2\log\#\mathcal{P}\left\lceil\frac{\log c_{1}C_{1}^{-1}\delta}{\log\Lambda}\right\rceil\right).$
Taking the $\log$ finishes the proof. ∎
Proposition 6.2.
There exists a constant $C>0$ such that for all $0<\varepsilon<\varepsilon_{0}$ small enough, all $\delta>0$ and all $n\geqslant n_{0}$,
$\log r(n,\delta,B(x,\varepsilon,n))\leqslant 2\frac{\log\#\mathcal{P}}{\log\Lambda}\log\delta^{-1}+C+\left(3+2\left\lfloor\frac{\max\tau}{\min\tau}\right\rfloor\right)(n+n_{0})s_{0}\log 2K,\quad\forall x\in\mathcal{S}_{n-1},$
where $K$ is such that $K_{n}\leqslant Kn$ for all $n$.
Actually, the estimate in Proposition 6.2 holds for all $x\in M$, but in lights of Proposition 6.1 we only state it for points of $\mathcal{S}_{n-1}$.
To prove the above, we rely on a following result which allows to count the number of times a flow orbit passes near a scatterer without making a collision.
Lemma 6.3.
For all $n_{0}$ fixed, there exists $\tilde{\varepsilon}=\tilde{\varepsilon}(n_{0})$ such that for all $\varepsilon\in(0,\tilde{\varepsilon})$ and all $x\in M$, the curve $\{\Phi_{t}(x)\mid t\in[0,\sum_{i=0}^{n_{0}-1}\tau(T^{i}x)]\}\subset\Omega$ passes at distance less than $\varepsilon$ to the scatterers without making a collision at most $\lfloor\max\tau/\min\tau\rfloor s_{0}n_{0}$ times.
Proof of Lemma 6.3.
Denote $t_{0}\coloneqq\lfloor n_{0}\max\tau/\min\tau\rfloor$. First, note that $(x,t)\mapsto\Phi_{t}(x)$ is continuous on the compact set $\Omega\times[-t_{0}\max\tau,t_{0}\max\tau]$ and, therefore, is uniformly continuous on this same set. The idea of the proof is to find for all $x$ a point $y$ in a neighbourhood such that the flow orbit of $y$ makes a grazing collision on the scatterers getting close to the trajectory of $x$. We start by considering the points making multiple grazing collisions. Let
$R_{2}\coloneqq\{x\in M\mid(T^{i}x)_{0\leqslant i\leqslant t_{0}}\text{ comprises at least two grazing collisions}\}.$
Remark that each point of $R_{2}$ must be a corner of $\bar{A}$, for some $A\in\mathcal{M}_{0}^{t_{0}}$. In particular $R_{2}$ is a finite set. Informally, define $\varepsilon_{2}$ to be the minimal distance (in $\Omega$) between the scatterers and the segment $[T^{i}x,T^{i+1}x]\subset\Omega$, $x\in R_{2}$ and $0\leqslant i<t_{0}$. More precisely, define $\pi:\Omega\to Q$ by $\pi(x)=\pi(\bar{x},v)=\bar{x}$ where $x=(\bar{x},v)\in\Omega=Q\times\mathbbm{S}^{1}/\sim$, and
$\tilde{d}([x,Tx],\partial Q)\coloneqq\inf\{\delta>0\mid\partial Q\cap(\pi([x,Tx])\times[-\delta,\delta])\text{ has $3$ or more connected components}\}$
Therefore, let
$\varepsilon_{2}\coloneqq\min_{x\in R_{2}}\min_{0\leqslant i<t_{0}}\tilde{d}([T^{i}x,T^{i+1}x],\partial Q)>0,$
and let $\delta_{2}$ be the delta associate to $\varepsilon_{2}/2$ by the uniform continuity of the flow.
We now consider the points making a single grazing collision: let
$R_{1}\coloneqq\{x\in M\mid(T^{i}x)_{0\leqslant i\leqslant t_{0}}\text{ comprises exactly one grazing collisions}\}=\mathcal{S}_{t_{0}}\smallsetminus R_{2}.$
Then $R_{1}^{\prime}\coloneqq\mathcal{S}_{t_{0}}\smallsetminus\mathcal{N}_{\delta_{2}/2}(R_{2})$ is compact. Let
$\varepsilon_{1}\coloneqq\min_{x\in R_{1}^{\prime}}\min_{0\leqslant i<t_{0}}\tilde{d}([T^{i}x,T^{i+1}x],\partial Q).$
By the uniform continuity of the flow, it follows that $\varepsilon_{1}>0$. Let $\delta_{1}>0$ be the delta associated with $\min(\varepsilon_{2}/2,\varepsilon_{1}/2,\delta_{2}/2,\mbox{diam}R_{2}/4)$ by the uniform continuity of the flow. Up to decreasing the value of $\delta_{1}$, we can assume that $\delta_{1}<\min(\varepsilon_{1}/2,\varepsilon_{2}/2)$. Finally, let $\delta_{0}$ be the delta associated to $\delta_{1}$ by the uniform continuity of the flow. We can assume that $\delta_{0}<\delta_{1}$.
We claim that any $\varepsilon>0$ such that $\varepsilon<\delta_{0}$ has the desired property. Indeed, assume that for some fixed $x\in M$ there exist two or more distinct $0\leqslant i_{j}<n_{0}$ such that $\tilde{d}([T^{i_{j}}x,T^{i_{j}+1}x],\partial Q)<\varepsilon$. Therefore, to each $i_{j}$ is associated a point $x_{j}\in M$, in a $\delta_{1}$-neighbourhood of $x$, such that the flow trajectories of $x_{j}$ and $x$ stay at distance at most $\min(\varepsilon_{2}/2,\varepsilon_{1}/2,\delta_{2}/2,\mbox{diam}R_{2}/4)$ for all times $t\in[0,n_{0}\max\tau]$. By this shadowing, we get that the orbit of $x_{j}$ passes at distance less than $\delta_{0}+\varepsilon_{1}/2<\varepsilon_{1}$ to $\partial Q$ (corresponding to the in-between collisions associated with $i_{k}$, $k\neq j$). In particular, by the definition of $\varepsilon_{1}$, we get that $x_{j}\notin R_{1}^{\prime}$. Thus $d_{M}(x_{j},R_{2})<\delta_{2}/2$ for all $j$. Since $d_{M}(x,x_{j})<\delta_{1}\leqslant\mbox{diam}R_{2}/4$, there exists a unique $y\in R_{2}$ such that $d_{M}(x_{j},y)<\delta_{2}/2$. Therefore $d(x,y)<\delta_{2}$. Thus, the flow orbits of $x$ and $y$ stay at distance less than $\varepsilon_{2}/2$ to each other for times $t\in[0,\sum_{i=0}^{n_{0}-1}\tau(T^{i}x)]$. Thus, the flow orbit of $y$ passes at distance less than $\varepsilon+\varepsilon_{2}/2<\varepsilon_{2}$ to $\partial Q$ when the orbit of $x$ passes at distance less than $\varepsilon$. Therefore, by the definition of $\varepsilon_{2}$, the flow orbit of $y$ makes a grazing collision at this instants. However, the flow orbit of $y$, for $t\in[0,n_{0}\max\tau]$ makes at most $\lfloor\max\tau/\min\tau\rfloor n_{0}$ collisions, and thus at most $\lfloor\max\tau/\min\tau\rfloor s_{0}n_{0}$ grazing collisions. Taking $\tilde{\varepsilon}=\delta_{0}$ finishes the proof. ∎
We have now all the tools in order to prove Proposition 6.2.
Proof of Proposition 6.2.
Let $0<\varepsilon<\varepsilon_{0}$ be small enough so that $\sin\varepsilon\leqslant\tilde{\varepsilon}/\max\tau$ (where $\tilde{\varepsilon}$ comes from Lemma 6.3),
$A\cap\mathcal{N}_{\varepsilon}(A\cap(\mathcal{S}_{-1}\cup\mathcal{S}_{1}))\subset\bigcup_{-1\leqslant i\leqslant 1}T^{i}\{|\varphi|>\varphi_{0}\},\quad\forall A\in\mathcal{P},$
and $B(x,\varepsilon)$ intersects at most $2K$ elements of $\mathcal{M}_{-1}^{1}$ for all $x\in M$.
Let $x\in M$. Then $B(x,\varepsilon,n)$ might intersects several elements of $\mathcal{M}_{-1}^{n}=\mathring{\mathcal{P}}_{0}^{n-1}$. We start by estimate the number of those sets.
Assume first that $n=n_{0}$. For any $A\in\mathcal{P}_{0}^{n-1}$, there exist $A_{i}\in\mathcal{P}$ such that $A=\bigcap_{i=0}^{n-1}T^{-i}A_{i}$. Thus $A\cap B(x,\varepsilon,n)\neq\emptyset$ if and only if $A_{i}\cap B(T^{-i}x,\varepsilon)\neq\emptyset$ for all $0\leqslant i<n$. In order to count the number of admissible $A_{i}$, we distinguish three mutually exclusive cases:
i) $B(y,\varepsilon)\cap(\mathcal{S}_{-1}\cup\mathcal{S}_{1})=\emptyset$. Therefore, $B(y,\varepsilon)$ is contained is a single element of $\mathcal{P}$.
ii) $y\in A\cap\mathcal{N}_{\varepsilon}(A\cap(\mathcal{S}_{-1}\cup\mathcal{S}_{1}))$ for some $A\in\mathcal{P}$. Therefore at least one of $y$, $Ty$ or $T^{-1}y$ is a $\varphi_{0}$-grazing collision, and $B(y,\varepsilon)$ intersects at most $2K$ elements of $\mathcal{P}$. This case occurs at most $3s_{0}n_{0}$ times.
iii) Otherwise, at least one of $[y,Ty]$ or $[T^{-1}y,y]$ passes $\varepsilon$ close to the scatterers without making a collision. Furthermore, $B(y,\varepsilon)$ intersects at most $2K$ elements of $\mathcal{P}$. By the Sublemma 6.3, this case occurs at most $2\lfloor\max\tau/\min\tau\rfloor s_{0}n_{0}$ times.
Combining the three above situations, we obtain
#$\#\{A\in\mathcal{P}_{0}^{n_{0}-1}\mid A\cap B(x,\varepsilon,n_{0})\neq\emptyset\}\leqslant(2K)^{\left(3+2\left\lfloor\frac{\max\tau}{\min\tau}\right\rfloor\right)s_{0}n_{0}}.$
Now, for general $n\geqslant n_{0}$, write $n=kn_{0}+\ell$, $0\leqslant\ell<n_{0}$. Then, any $A\in\mathcal{P}_{0}^{n-1}$ can be written $A=\bigcap_{i=0}^{k-1}T^{in_{0}}A_{i}\cap B$, where $A_{i}\in\mathcal{P}_{0}^{n_{0}-1}$ and $B\in\mathcal{P}_{0}^{\ell-1}$. Thus, $A\cap B(x,\varepsilon,n)\neq\emptyset$ if and only if $A_{i}\cap B(T^{in_{0}}x,\varepsilon,n_{0})\neq\emptyset$ for all $i$ and $B\cap B(T^{kn_{0}}x,\varepsilon,\ell)\neq\emptyset$. Therefore, by the computation done in the case $n=n_{0}$, and since $\mathcal{P}_{0}^{n-1}=\mathcal{M}_{-1}^{n}$, we obtain
#$\#\{A\in\mathcal{M}_{0}^{n}\mid A\cap B(x,\varepsilon,n)\neq\emptyset\}\leqslant\#\mathcal{M}_{-1}^{0}\left((2K)^{\left(3+2\left\lfloor\frac{\max\tau}{\min\tau}\right\rfloor\right)s_{0}n_{0}}\right)^{\left\lceil\frac{n}{n_{0}}\right\rceil}.$
For each $A\in\mathcal{M}_{0}^{n-1}$, let $F_{A}\subset A$ be $(\delta/2,n)$-separated with maximal cardinality if $A\cap B(x,\varepsilon,n)\neq\emptyset$, and $F_{A}=\emptyset$ otherwise. Thus, the union of the $F_{A}$ is $(\delta,n)$-spanning $B(x,\varepsilon,n)$. Therefore, using that
#$\#\{A\in\mathcal{M}_{0}^{n-1}\mid A\cap B(x,\varepsilon,n)\neq\emptyset\}\leqslant\#\{A\in\mathcal{M}_{0}^{n}\mid A\cap B(x,\varepsilon,n)\neq\emptyset\},$
and estimates on each #$\#F_{A}$ obtained in the proof of Proposition 6.1 yield
$r(n,\delta,B(x,$ $\varepsilon,n))\leqslant\sum_{A\in\mathcal{M}_{0}^{n-1}}\#F_{A}$
(#$\leqslant(\#\mathcal{M}_{-1}^{0})^{2}\#\mathcal{P}\exp\left(-2\log\#\mathcal{P}\left\lceil\frac{\log c_{1}C_{1}^{-1}\delta}{\log\Lambda}\right\rceil\right)\left((2K)^{\left(3+2\left\lfloor\frac{\max\tau}{\min\tau}\right\rfloor\right)s_{0}n_{0}}\right)^{\left\lceil\frac{n}{n_{0}}\right\rceil}.$
Taking the $\log$ of the above estimate ends the proof. ∎
Corollary 6.4.
For any finite horizon billiard table, $h^{*}(T)\leqslant\left(3+2\left\lfloor\frac{\max\tau}{\min\tau}\right\rfloor\right)s_{0}\log 2K$.
Proof.
By Propositions 6.1 and 6.2, we get that
$\limsup_{n\to\infty}\frac{1}{n}\log\sup_{x\in M}r(n,\delta,B(x,\varepsilon,n))\leqslant\left(3+2\left\lfloor\frac{\max\tau}{\min\tau}\right\rfloor\right)s_{0}\log 2K.$
Taking the appropriate limits in $\delta$ and in $\varepsilon$ yields the result. ∎
7. Examples
Denote by $Q^{(d)}$ and $Q^{(R,R^{\prime})}$ the billiard tables with parameters $d$ and $(R,R^{\prime})$ respectively, as depicted in Figure 1. These billiard tables have finite horizons and scatterers do not intersects whenever
$d\in\,$ $\mathcal{D}_{1}\coloneqq(2,4/\sqrt{3}),$
$(R,R^{\prime})\in\,$ $\mathcal{D}_{2}\coloneqq\{(r,r^{\prime})\in\mathbbm{R}^{2}\mid 0<r<r^{\prime}<\frac{1}{2},\text{ and }\frac{1}{2}<r+r^{\prime}<\frac{\sqrt{2}}{2},\text{ and }r^{\prime}>\frac{\sqrt{2}}{4}\}.$
Relying on numerical computations performed in respectively [19] and [18], Baladi and Demers proved that $Q^{(d)}$ and $Q^{(R,0.4)}$ satisfy the sparse recurrence condition for each of the finitely many values $d\in\mathcal{D}_{1}$ and $(R,0.4)\in\mathcal{D}_{2}$ for which the computations have been performed.
In this section, we prove that there is an open set of parameters $d$ (resp $R$) for which $Q^{(d)}$ (resp $Q^{(R,0.4)}$) satisfies the sparse recurrence condition. This is the content of Proposition7.1. We also prove that for most of these parameters, the associated singularities are negligible. This is the content of Proposition 7.2.
$d$$(a)$$R$$R^{\prime}$$(b)$
Figure 1. (a) The Sinai billiard on a hexagonal lattice studied in [19], scatterer of radius 1, and distance $d$ between the centers of adjacent scatterers. (b) The Sinai billiard on a square lattice with scatterers of radius $R<R^{\prime}$ studied in [18]. The boundary of a single cell is indicated by dashed lines in both tables. (Figure taken from [3, Figure 2].)
Proposition 7.1.
For any $d\in(2.15,4/\sqrt{3})$ (resp. $R\in(0.1,0.25)$), the table $Q^{(d)}$ (resp. $Q^{(R,0.4)}$) has the sparse recurrence property.
Proof.
Since $h_{*}\geqslant h_{\mu_{\tiny{\mbox{SRB}}}}(T)$, It is sufficient to provide a lower bound on $h_{\mu_{\tiny{\mbox{SRB}}}}(T)$ larger than $s_{0}\log 2$. From [12, Equations (3.49),(3.50) and (3.33)], we get that
(7.1) $h_{\mu_{\tiny{\mbox{SRB}}}}(T)\geqslant\frac{1}{2}\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\log\left(1+\tau_{\min}\frac{2\kappa_{\min}}{\cos\varphi}\right)\cos\varphi\,{\rm d}\varphi,$
where $\kappa_{\min}$ denotes the minimal curvature of the scatterers.
For the family of billiard tables studied in [19], we get that $\kappa_{\min}=1$ and $\tau_{\min}=d-2$. By differentiating, it follows that the right-hand side of (7.1) is an increasing function of the variable $d$. Now, using numerical computations, we get that for $d_{0}=2.15$,
$\frac{1}{2}\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\log\left(1+\frac{2(d_{0}-2)}{\cos\varphi}\right)\cos\varphi\,{\rm d}\varphi>0.36>\frac{1}{2}\log 2\geqslant s_{0}\log 2.$
Hence, for all $d\in(2.15,4/\sqrt{3})$, $Q^{(d)}$ has the sparse recurrence property.
For the family of billiard table studied in [18], we get that $\kappa_{\min}=\frac{1}{R^{\prime}}=2.5$ and
$\tau_{\min}=\begin{cases}1-2R^{\prime}&\quad\text{ if }0.1<R\leqslant\sqrt{2}/2-(1-2R^{\prime}),\\ \sqrt{2}/2-(R+R^{\prime})&\quad\text{ if }\sqrt{2}/2-(1-2R^{\prime})\leqslant R<\sqrt{2}/2-R^{\prime}.\end{cases}$
Fixing $R^{\prime}=0.4$, we get that the right-hand side of (7.1) is a decreasing function of the variable $R$. Now, using numerical computations, we get that for $R_{0}=1/4$
$\frac{1}{2}\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\log\left(1+\frac{5(\sqrt{2}/2-R^{\prime}-R_{0})}{\cos\varphi}\right)\cos\varphi\,{\rm d}\varphi>0.347>\frac{1}{2}\log 2\geqslant s_{0}\log 2.$
Hence, for all $R\in(0.1,1/4)$, $Q^{(R,0.4)}$ has the sparse recurrence property. ∎
Proposition 7.2.
There exists a dense $G_{\delta}$ set $G_{1}\subset\mathcal{D}_{1}$ (resp. $G_{2}\subset(0.1,\frac{\sqrt{2}}{2}-0.4)$) such that if $d\in G$ (resp. $R\in G_{2}$), then $Q^{(d)}$ (resp. $Q^{(R,0.4)}$) has negligible singularities.
Proof.
The proof is very similar to the one of Proposition 4.6. For completeness, we provide the modifications to be done. Denote
$U_{k}\coloneqq\{d\in\mathcal{D}_{1}\mid\bigcup_{i=3}^{k}{\rm Fix}\,\,T_{d}^{i}\cap\mathcal{S}_{0}=\emptyset\},\,k\geqslant 3,$
where $T_{d}$ is the collision map associated to $Q^{(d)}$. Note that for $d_{1},d_{2}\in\mathcal{D}_{1}$,
(7.2) $d_{C^{3}}(Q^{(d_{1})},Q^{(d_{2})})=|d_{1}-d_{2}|.$
To prove that $U_{k}$ is open, let $d_{0}\in U_{k}$ and $Q_{0}=Q^{(d_{0})}$. By contradiction, assume that for all $n\geqslant 1$, there exists $d_{n}\in\mathcal{D}_{1}$ such that $|d_{0}-d_{n}|<1/n$. Denote $Q_{n}=Q^{(d_{n})}$. The rest of the proof of openness is exactly as in the proof of Proposition 4.6.
To prove that $U_{k}$ is dense in $\mathcal{D}_{1}$, let $d_{0}\in\mathcal{D}_{1}$ and $\delta>0$. The proof is then as in the one of Proposition 4.6 where the successive perturbed tables $Q_{n}=Q^{(d_{n})}$ are obtained by constructing an increasing sequence $(d_{n})_{n\geqslant 1}$ bounded from above by $d_{0}+\delta$. The construction of this sequence uses the same argument as if all of the orbit of $x_{n}$ were composed of points of the type $y_{i}$ (and none of the type $z_{i}$), as well as the equivalence of metrics (7.2).
Finally, set $G_{1}=\bigcap_{k\geqslant 3}U_{k}$. The construction of $G_{2}$ is analogous (where instead of increasing values of the parameter $d$, we use decreasing values of $R$). ∎